### v2

# references.bib

Abstract
Remote sensing map products are used to obtain estimates of environmental quantities, such as deforested area or the effect of conservation zones on deforestation. However, the quality of map products varies, and — because maps are outputs of complex machine learning algorithms that take in a variety of remotely sensed variables as inputs — errors are difficult to characterize. Without capturing the biases that may be present, naive calculations of population-level estimates from such maps are statistically invalid. In this paper, we compare several uncertainty quantification methods — stratification, Olofsson area estimation method, and prediction-powered inference — that combine a small amount of randomly sampled ground truth data with large-scale remote sensing map products to generate statistically valid estimates. Applying these methods across four remote sensing use cases in area and regression coefficient estimation, we find that they result in estimates that are more reliable than naively using the map product as if it were 100% accurate and have lower uncertainty than using only the ground truth and ignoring the map product. Prediction-powered inference uses ground truth data to correct for bias in the map product estimate and (unlike stratification) does not require us to choose a map product before sampling. This is the first work to (1) apply prediction-powered inference to remote sensing estimation tasks, and (2) perform uncertainty quantification on remote sensing regression coefficients without assumptions on the structure of map product errors. To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a holdout ground truth dataset to be used for calibration in uncertainty quantification alongside their maps. Data and code are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
1 Introduction
Remote sensing maps are widely used to obtain estimates for environmental quantities [hansen2013high, baccini2012estimated, potapov2022global, wang2019great, deines2019satellites]. Map products created by training machine learning models on remotely sensed data allow researchers to cheaply access predictions for a variable of interest at a high resolution over large regions.
However, the quality of map products varies, and resulting population-level estimates of quantities of interest may be biased and statistically invalid [jain2020benefits, bastin2017extent, venter2024uncertainty]. For example, [bastin2017extent] showed that human labeling of high-resolution imagery results in an estimate of global forest extent in dryland biomes that is 40% to 47% higher than previous estimates from remotely sensed forest cover map products. Moreover, different map products may conflict with each other and result in significantly different estimates over the same region. For example, [alix2023remotely] found that two commonly used satellite-based remote sensing maps of forest cover in Mexico differ significantly from each other, implying that there is misclassification error in at least one of the algorithms used to categorize pixels as forest or non-forest. Another recent study of three satellite-based global land cover/land use map products found that the maps are biased toward different land cover classes (e.g., one map overestimates shrub cover and another overestimates grass cover relative to the other maps) [venter2022global].
Additionally, errors in map products are often difficult to characterize due to the complexity of input data and lack of interpretability in the ML models used to generate the maps. Examples of systematic errors in map products trained on satellite data include “nonrandom misclassification, saturation effects, atmospheric effects, and cloud cover,” all of which may lead to biased inference [jain2020benefits]. This has prevented environmental scientists, economists, and policymakers from using remote sensing maps in their analyses. The World Bank has cited uncertainty in remote sensing maps as a barrier to their use in water resources management [world2018mainstreaming], and the United Nations Task Team on Satellite Imagery and Geospatial Data wrote that the “shift in paradigm from traditional statistical methods (e.g., counting, measuring by humans) towards estimation from sensors (with associated aspects of validity, approximation, uncertainty)… will require convincing, statistically sound results” [united2017earth]. Correctly accounting for map errors, then, is crucial for maximizing the potential of remote sensing data to expand scientific knowledge and advance practical operations.
For those who currently reject estimation from remote sensing as too unreliable, the alternative is to use ground truth observations for the variable of interest. However, ground truth data is often sparse because it is expensive to collect, as it may require fieldwork or manual labelling of remote sensing imagery. Few ground truth points result in greater uncertainty (e.g., wider confidence intervals) in estimates. Another alternative is to use both ground truth data and remote sensing maps to generate estimates, using the ground truth points to correct for biases that may exist in the map products. This generally reduces uncertainty in the estimates compared to using ground truth alone, since map products contain a large amount of data that, while possibly biased, still give useful information about the true values we wish to measure.
Figure 1: Motivating example: map-only inference is statistically invalid. The 95% confidence interval for logistic regression coefficient relating Brazilian Amazon deforestation to slope using $N=963548$ remote sensing map points and $n=1386$ ground truth points. The imputed slope coefficient from the map alone falls outside the confidence intervals for all methods that use ground truth points.
A motivating example.
In this work, we discuss methods that correct map products with ground-truth data to give statistically-valid confidence intervals. We give a brief preview here. We seek to estimate logistic regression coefficients for the association between Brazilian Amazon deforestation in 2000–2015 and several covariates, including terrain slope. Using nearly one million remote sensing map product points and $1386$ ground truth points, we compare the $95\%$ confidence intervals for the slope coefficient with four methods (Figure 1) described shortly. See Section 4.1 for a complete description of the experimental setup.
The “classical” estimate using only ground truth points is statistically valid but has the widest confidence interval, failing to take advantage of existing map products. Meanwhile, methods that use a map product along with the ground truth points produce an estimate with lower uncertainty while still being valid.
Notably, the slope coefficient computed from the map product alone (“naive imputation”) is approximately zero, which falls outside the confidence intervals for all methods that use ground truth points. That is, if we had used only the map product, we would not have discovered the statistically significant negative relationship between slope and deforestation. Methods like PPI allow us to use ground truth points to correct for bias in the map product and thus avoid misleading conclusions.
Figure 2: Overview of uncertainty quantification using ground truth data and remote sensing map products. Olofsson area estimation, stratification, and prediction powered inference methods use a small amount of ground truth data along with a map product to produce confidence intervals for the quantity of interest. We also compare these methods against classical uncertainty quantification (using only ground truth) and naive imputation (using only the map product).
Our contribution.
In this paper, we consider uncertainty quantification methods that combine a small number of randomly sampled ground truth data ($n$ points) with large-scale remote sensing map products ($N$ points, where $N$ is much larger than $n$) in order to generate statistically valid estimates with low uncertainty (Figure 2). Stratification uses map products to determine a sampling strategy for ground truth points that reduces the variance of estimates; it has been used in the remote sensing community for decades [hay1979sampling, fitzpatrick1981comparison, stehman1998design, olofsson2014good]. Prediction-powered inference (PPI) [angelopoulos2023prediction, angelopoulos2023ppi++] is a recently developed uncertainty quantification method that uses a small set of ground truth points to calibrate and correct for bias in a much larger set of ML-generated predictions, resulting in more precise estimates than using the ground truth points alone. We compare stratification and prediction-powered inference against classical estimation (using only ground truth) and naive imputation (using only a map product).
We apply these methods to two of the most common scientific uses of remote sensing maps: area estimation and regression coefficient estimation. Area estimation includes estimating both total area occupied by a land cover/land use category and total change in land cover/land use over time. The methods applied to area estimation also apply to estimating sums or averages of continuous variables, such as total carbon stock. Regression coefficient estimation refers to estimating linear or logistic regression coefficients for causal or associative relationships between two or more environmental variables. For regressions, this paper only accounts for uncertainty in the remotely sensed dependent variable, while the covariates are assumed to have no errors. To illustrate the methods, we consider four specific remote sensing estimation tasks spanning crop area estimation, deforestation quantification, and associating deforestation and tree cover to other geophysical variables; see Table 1. Data and code for these use cases are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
Estimation task Target variable Region Year Map product(s) Ground truth Covariates Link to code
Area Maize Iowa 2022 USDA NASS Cropland Data Layers (CDL) [nass2024usda] Simulated N/A link
Area Deforestation Brazilian Amazon 2000-2015 1. NASA Global Forest Cover Change [townshend2016global]
2. Amazon ecoregion
deforestation strata [bullock2020satellite]
$n=1386$
from [bullock2020satellite]
N/A link
Logistic
regression
coefficients
Distance from roads, distance from rivers, elevation, slope link
Linear
regression
coefficients
Tree cover Contiguous USA 2021 USFS Tree Canopy Cover [forest2023usfs] $n=983$
manually labelled Google Maps high-resolution satellite images
Aridity index,
elevation, slope
link
Table 1: Overview of use cases and datasets. We estimate maize area in Iowa, deforestation area and logistic regression coefficients in the Brazilian Amazon, and tree cover linear regression coefficients in the United States. The Iowa maize area use case is a simulation: we treat a small random sample of CDL map product points as “ground truth” data and observe the effects of artificially degrading the map product relative to the ground truth. Code implementations of the use cases are linked in the rightmost column.
Combining ground truth data with map products results in statistically valid estimates that are more reliable than naive imputation and have lower uncertainty than classical estimates. As we will show, naive imputation from maps without bias correction can result in completely incorrect area and regression coefficient estimates. To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a holdout ground truth dataset (such as the usual test set) to be used for calibration in uncertainty quantification alongside their maps, or, if that is not done, for map users to generate their own holdout set for calibration, if possible (e.g., from visual inspection of remote sensing data).
Related work.
Previously, Olofsson et al. [olofsson2013making] described an area estimation method that uses the confusion matrix between land cover ground truth and map product points to correct for misclassification bias in the map product. The Olofsson method is commonly used in the remote sensing community to generate confidence intervals for area estimates. However, it does not apply to regression coefficient estimation tasks. One contribution of the present work is a comparison of this method and PPI. We find both experimentally and mathematically that these two methods produce similar outputs for area estimation.
Other methods have been developed to reduce bias in regression coefficient estimates from remote sensing data. For example, [alix2023remotely] estimates the effect of a conservation program on deforestation in Mexico. They correct for bias in the remotely sensed deforestation binary variable by modelling misclassification error as dependent on environmental covariates via a logit model. Another study [gordon2023remote] estimates logistic regression coefficients for the relationship between roads and forest cover in West Africa. They train an adversarial debiasing model with ground truth data to correct for bias arising from correlations between distance from roads and the measurement error in remotely sensed forest cover. Another study [proctor2023parameter] uses co-located ground truth and remotely sensed data to show that using remote sensing data alone leads to biased regression coefficient estimates for several effect size estimation tasks in the US. They use a linear measurement error model and apply multiple imputation: a statistical method that uses a small amount of ground truth data to estimate remote sensing measurement error and impute corrected values for remotely-sensed variables. Note that this method requires an assumption about the error structure of the remotely-sensed variables, whereas the methods investigated herein do not.
Another approach to uncertainty quantification is conformal prediction, which generates statistically valid confidence sets for machine learning predictions at individual data points. For example, [valle2023quantifying] applies conformal prediction to quantify pixel-level uncertainty for a land cover/land use map product in the Brazilian Amazon, outputting a 90% predictive set of land cover classes for each pixel (i.e., the set contains the true land cover class 90% of the time). This approach differs from our work because our methods estimate total area and regression coefficients over an entire region, rather than quantifying uncertainty at individual points or pixels.
Unlike previous works, this paper is the first work to apply prediction-powered inference to remote sensing use cases in area and regression coefficient estimation. Unlike stratified random sampling, PPI does not require that ground truth points be sampled according to a pre-determined stratification map; as a result, it allows for greater flexibility in ground truth data usage across different downstream applications. Additionally, this is the first work to combine map product and ground truth data to perform uncertainty quantification on remote sensing regression coefficients without assuming an error structure on the map product predictions.
2 Background: Uncertainty Quantification
We formally describe uncertainty quantification methods that generate confidence intervals with remote sensing variables. As two baselines, we first consider methods that use only the ground-truth data (which is valid but loses information) or only use the remote-sensing data (which is not statistically valid). We then turn to methods that use both sources of data to produce statistically-valid confidence intervals. See Table 2 for a summary.
Method Data Used Statistically Valid? Caveats
Map Ground Truth
Naive imputation
Classical
Stratification Ground truth points sampled
using stratified map
Olofsson Area estimation only
PPI
Table 2: Summary of uncertainty quantification methods. PPI is the most flexible method that incorporates both ground truth and map product data.
2.1 Baseline methods
Classical: use only the ground-truth data.
First, classical uncertainty quantification uses ground truth observations, sampled uniformly at random, to estimate the quantity of interest. For example, suppose there is a target variable $Y$ that can be observed over a given area, and we wish to generate a 95% confidence interval for its population mean $\theta^{\star}=E[Y]$. If we have $n$ ground truth observations $Y_{1},Y_{2},\dots Y_{n}$ sampled uniformly at random, we can compute the sample mean
$\hat{\theta}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$
and standard error
$SE=\sqrt{\frac{1}{n}\mathrm{Var}(Y_{i})}=\frac{1}{n}\sqrt{\sum_{i=1}^{n}(Y_{i}-\hat{\theta})^{2}}$
to produce the classical 95% confidence interval
$C=(\hat{\theta}-1.96\cdot SE,\ \hat{\theta}+1.96\cdot SE).$
This method is conservative — if the size $n$ of the ground truth dataset is small, the classical confidence interval will be large.
Naive imputation: use only the remote sensing map.
If we have a remote sensing data product with a large number $N$ of predictions $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{N}$ over the area of interest, we could treat these predictions like ground truth points and use the the formula above with $\hat{Y}$ in place of $Y$. This would produce a much smaller confidence interval, since we would use $N$ instead of $n$ in the denominator of the standard error. However, this naive imputation approach is not statistically valid, as it does not account for error or uncertainty in the model used to generate the map product predictions.
2.2 Stratification
Stratified random sampling uses remotely sensed maps to determine a sampling strategy for ground truth points that reduces the variance of estimates. Stratification is commonly used in the remote sensing community.
We first use a map product to partition the area of interest into $K$ strata (sub-areas) such that the target variable is predicted to have low variance within each stratum. Then we sample $n_{k}$ ground truth points uniformly at random from each stratum $1\leq k\leq K$. We can choose $n_{k}$ to be larger for strata predicted to have higher uncertainty to decrease the standard error within the strata and hence make the overall standard error smaller with the same total number of samples.
Suppose we again wish to estimate $\theta^{\star}=E[Y]$ using a stratified random sample. For each stratum $k$, we compute the stratum sample mean $\hat{\theta}_{k}$ and stratum standard error $SE_{k}$ using the same formulae as the classical method. Suppose that in the map product, each stratum $k$ takes up proportion $10<A_{k}<$ 1 of the total area. Then, we compute the stratified mean
$\hat{\theta}_{\text{strat}}=\sum_{k=1}^{K}A_{k}\hat{\theta}_{k}$
and stratified standard error
$SE_{\text{strat}}=\sqrt{\sum_{k=1}^{K}A_{k}^{2}\cdot SE_{k}^{2}}$
to produce the stratified 95% confidence interval
$C_{\text{strat}}=(\hat{\theta}_{\text{strat}}-1.96\cdot SE_{\text{strat}},\ \hat{\theta}_{\text{strat}}+1.96\cdot SE_{\text{strat}}).$
The advantage of stratification with an accurate map and well-chosen sampling strategy is that the variance within each stratum will be small, which results in a smaller overall confidence interval compared to the classical method. If the map predictions are inaccurate, then the ground truth points likely have higher variance within each map stratum, resulting in a larger confidence interval. Thus the uncertainty quantification incorporates the accuracy of the map product predictions.
For example, in our Brazilian Amazon deforestation area estimation use case, we use a remotely sensed stratification map created by [bullock2020satellite] using Landsat surface reflectance imagery at 30 m resolution. The map (Figure 3) partitions the Amazon into seven strata related to forest change from 1995-2017: stable forest, stable non-forest, deforestation, natural disturbance, natural disturbance and deforestation, regrowth, and buffer. The strata are designed so that points within each stratum have low variance in deforestation status (for instance, most points in the “stable forest” and “stable non-forest” strata do not experience deforestation in the time period of interest). As a result, our stratified confidence interval for deforestation area is smaller than the classical confidence interval.
Figure 3: Stratified random sample is a standard approach for estimation with remote sensing maps. This example shows a map of the Brazilian Amazon used for stratification. The strata are designed so that points within each stratum have low variance in deforestation status, which reduces the uncertainty when estimating area deforested. Adapted from [bullock2020satellite].
2.3 Prediction-Powered Inference
Prediction-powered inference (PPI) [angelopoulos2023prediction] is an uncertainty quantification method that computes confidence intervals using a small set of ground truth points and a much larger set of machine learning generated predictions. While stratified sampling requires us to commit to a map product and stratification scheme before sampling the ground truth points, the map product used in PPI can be chosen after sampling. The main idea is that PPI assesses the bias in the map product using the ground truth data and then corrects it.
PPI evaluates the accuracy of the map product by comparing the ground truth labels $Y_{1},Y_{2},\dots Y_{n}$ to the map product predictions $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{n}$ at the corresponding ground truth point locations. For a concrete example, suppose we wish to estimate $\theta^{\star}=E[Y]$ as before. The PPI point estimate is
$\hat{\theta}_{\text{PPI}}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}}_{\text{naive imputation}}-\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})}_{\text{bias correction}}.$
The PPI standard error is
$SE_{\text{PPI}}$ $=\sqrt{\frac{1}{N}\mathrm{Var}(\hat{Y}_{i}^{\prime})+\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})}$
$\approx\sqrt{\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})},$
where the approximation holds when $N$ is much larger than $n$. This yields the PPI 95% confidence interval,
$C_{\text{PPI}}=(\hat{\theta}_{\text{PPI}}-1.96\cdot SE_{\text{PPI}},\ \hat{\theta}_{\text{PPI}}+1.96\cdot SE_{\text{PPI}}).$
The first term in $SE_{\text{PPI}}$ is the variance corresponding to using only map predictions to estimate $\theta$, i.e. $\frac{1}{N}\mathrm{Var}(\hat{Y}_{i}^{\prime})$. The second term is the variance from the bias correction, $\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})$. As the number of map product points $N$ grows large, $SE_{\text{PPI}}$ approaches $\sqrt{\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})}$. This will be a realistic approximation in many cases, since remote sensing maps at a global, national, or province level can contain millions or even billions of pixels, whereas the number of ground truth points $n$ is commonly in the hundreds or thousands. The greater the accuracy of the map product, the smaller $\mathrm{Var}(\hat{Y}_{i}-Y_{i})$ will be and the smaller the PPI confidence interval will be. Note that the PPI confidence interval is smaller than the classical confidence interval if $\mathrm{Var}(\hat{Y}_{i}-Y_{i})<\mathrm{Var}(Y_{i})$, which will be true for any map whose predictions correlate positively with the ground truth.
PPI can also give confidence intervals for linear and logistic regression coefficients, but the formulae are more complex. We sketch the main idea here and refer the reader to [angelopoulos2023ppi++, angelopoulos2023prediction] for a full exposition. In brief, PPI begins like naive imputation by using a large set of map predictions $\hat{Y}^{\prime}_{1},\hat{Y}^{\prime}_{2},\dots\hat{Y}^{\prime}_{N}$ to compute a measure of fit $m_{\theta}$. Then, PPI computes a bias correction $\Delta_{\theta}$ by comparing the ground truth data $Y_{i}$ to the predictions $\hat{Y}_{i}$. The relevant formula for the measure of fit and bias correction depends on what we seek to estimate — areas, binary regression coefficients, linear regression coefficients, etc. — see Table 3. Lastly, the naive $m_{\theta}$ and the bias correction $\Delta_{\theta}$ are combined to construct the PPI 95% confidence interval
$C_{\text{PPI}}=\{\theta:|m_{\theta}+\Delta_{\theta}|\leq w_{\theta}\}$
where $w_{\theta}$ is a constant that depends on the confidence level (95% in this case).
PPI can be used with either uniform random sampling or stratified random sampling of the ground truth points [fisch2024stratified]. Stratified (weighted) PPI can be applied as follows for mean estimation. Each ground truth value $Y_{i}$ and labeled prediction $\hat{Y}_{i}$ is multiplied by weight $w_{i}=A_{k}\frac{n}{n_{k}}$ where $k$ is the stratum where point $i$ is located, $A_{k}$ is the area proportion of stratum $k$ and $n_{k}$ is the number of ground truth samples in stratum $k$. Similarly, each unlabeled prediction $\hat{Y}^{\prime}_{i}$ in stratum $k$ is multiplied by weight $w_{i}=A_{k}\frac{N}{N_{k}}$ where $N_{k}$ is the number of map product points in stratum $k$. Then the same PPI formulae as above can be used.
Estimand Measure of fit $m_{\theta}$ Bias correction $\Delta_{\theta}$
Mean $\theta-\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}$ $\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})$
Logistic regression $\frac{1}{N}\sum_{i=1}^{N}\left(\frac{X_{i}^{\prime}}{1+e^{-\theta^{T}X_{i}^{\prime}}}-\hat{Y}_{i}^{\prime}\right)$ $\frac{1}{n}\sum_{i=1}^{n}X_{i}\left(\hat{Y}_{i}-Y_{i}\right)$
Linear regression $\theta-(X^{\prime})^{\dagger}\hat{Y}^{\prime}$ $X^{\dagger}\left(\hat{Y}-Y\right)$
Table 3: Prediction-powered inference (PPI) formulae for three common estimation tasks found in remote sensing: mean estimation, logistic regression coefficients, and linear regression coefficients.
We use the ppi-python software package to implement the unweighted and weighted versions of PPI for our use cases. The software includes a tuning parameter $\lambda$ that we set using the map product points and ground-truth points. The confidence intervals are statistically valid even with this data-driven tuning [angelopoulos2023ppi++].
2.4 Olofsson et al. (2013) area estimation
Olofsson et al. [olofsson2013making] introduces an uncertainty quantification method that uses the confusion matrix between ground truth and map product land cover labels to estimate the area of a land cover class. This method is commonly used in the remote sensing community. Like PPI, the Olofsson method computes confidence intervals by using ground truth points to correct for bias in the map product estimates. In fact, we show mathematically in Appendix A that the two methods result in similar confidence intervals as the number of map product points $N$ grows large. However, unlike PPI, the Olofsson method only applies to area estimation but not regression coefficient estimation.
Turning to the method, suppose we have a map product with $K$ land cover classes, as well as $n$ randomly sampled ground truth points. The Olofsson method constructs the $K\times K$ misclassification confusion matrix between ground truth points and map product predictions at the ground truth point locations. For each $1\leq i\leq K$ and $1\leq j\leq K$, let $n_{ij}$ denote the number of points that are in map class $i$ and ground truth class $j$. For each class $1\leq i\leq K$, let $10<A_{i}<$ 1 denote the proportion of the total map area taken up by the map class.
Then, the area estimate for the area of class $j$ is
$\hat{\theta}_{j,\text{Olofsson}}=\sum_{i=1}^{K}A_{i}\frac{n_{ij}}{n_{i\cdot}}$
with corresponding standard error
$SE_{j,\text{Olofsson}}=\sqrt{\sum_{i=1}^{K}A_{i}^{2}\frac{\frac{n_{ij}}{n_{i\cdot}}(1-\frac{n_{ij}}{n_{i\cdot}})}{n_{i\cdot}}}.$
The 95% confidence interval for the area of class $j$ is
$C_{j,\text{Olofsson}}=(\hat{\theta}_{j,\text{Olofsson}}-1.96\cdot SE_{j,\text{Olofsson}},\ \hat{\theta}_{j,\text{Olofsson}}+1.96\cdot SE_{j,\text{Olofsson}}).$
Area estimation is a special case of mean estimation in which we estimate the mean of a binary variable that is $1$ for all points in class $j$ and 00 for all points outside of class $j$. Thus, we can also perform area estimation using methods such as PPI and stratification. If the $n$ ground truth points are obtained by stratified random sampling using the $K$ map land cover classes as the strata, the Olofsson confidence interval is identical to the stratified confidence interval.
2.5 Evaluating performance using effective sample size
We use effective sample size to compare the effectiveness of uncertainty quantification methods. Suppose a classical confidence interval of width $w$ is generated using $n$ ground truth points. If another uncertainty quantification method uses $n$ ground truth points (along with a map product) to generate a confidence interval of width $w^{\prime}$, then the effective sample size of the new method is
$n_{\text{effective}}=n\cdot\left(\frac{w}{w^{\prime}}\right)^{2}.$
This is the number of ground truth points that would be required to generate a classical confidence interval of width $w^{\prime}$.
3 Area estimation
In this section, we demonstrate uncertainty quantification for two area estimation use cases. First, we simulate estimating maize area in Iowa in 2022, artificially degrading the map product quality relative to simulated “ground truth” points, in order to observe the effect of map product noise and bias on the area estimate under different methods. We will see through this simulation the importance of bias correction before using maps to estimate total area. We then use real ground truth data and real map products developed by researchers to estimate change in tree cover in the Brazilian Amazon in 2000–2015. We compare the classical method, naive imputation, stratification, PPI, stratified PPI, and the method of Olofsson et al. (2013) [olofsson2013making].
In each use case, we use binary data where each point has value 1 if it belongs to the land cover class of interest (e.g., cropland or deforested land), and value 0 otherwise. Then, estimating the area fraction of the land cover class is equivalent to estimating the mean of all points in the region.
3.1 Simulation: Maize area in Iowa
Problem Formulation
We simulate estimating the fraction of maize area in Iowa in 2022. To probe the effect of map product noise and bias on different methods, we artificially degrade the map product quality relative to simulated “ground truth” points, as described next.
For this simulation, we use the USDA’s Cropland Data Layer (CDL) [nass2024usda] as the ground truth for crop types in Iowa. CDL is a crop type land cover map created using satellite imagery and agricultural ground truth data in the continental United States. It has 94.2% producer’s accuracy and 97.2% user’s accuracy for the maize land cover class in Iowa. We sample $N=100000$ locations uniformly at random from a binary maize/non-maize map derived from CDL (Figure 4). The mean of the 100000 map product points is 0.35, which we will consider the “true” fraction of maize for the simulation.
In each trial, we sample $n=100$ “ground truth” points uniformly at random from the $N=100000$ map product points. Initially, we set the map product and ground truth equal to each other at the 100 ground truth locations. We then experiment with degrading the map product quality as follows. We systematically add noise or bias to the map product points while keeping the ground truth points constant, and then compare our maize fraction estimates across naive imputation, classical estimation, and PPI. We average results over 100 trials.
Adding unbiased noise.
We define adding unbiased noise at level $p$ to the map product as resampling each of the 100000 map product points with probability $p$ from the Bernoulli(0.35) distribution. In other words, about $p$ fraction of the points are randomly labeled, while the remaining $1-p$ points are unchanged. Each resampled point is labeled as maize with probability 0.35 and non-maize with probability 0.65. We experiment with noise probabilities $p$ from 0 to 1 at increments of 0.1.
Adding bias.
We also consider adding bias to the map product. Here, we define adding bias $p$ to the map product as resampling each of the 100000 map product points with probability 0.4 from the Bernoulli($p$) distribution. In other words, about 40% of the points are randomly labeled, while the remaining 60% points are unchanged. Each resampled point is labeled as maize with probability $p$ and non-maize with probability $1-p$. We experiment with bias values $p$ from 0 to 1 at increments of 0.1. When $p=0.35$, this does not add bias, but as $p$ ranges farther from $0.35$, the amount of bias increases.
Figure 4: Iowa maize area estimation datasets. The Iowa maize binary map product is derived from the USDA’s Cropland Data Layer, and $n=100$ uniformly randomly sampled “ground truth points” are shown.
Results
As we add increasing amounts of random noise to the map, all confidence intervals for maize area remain centered at the true maize fraction of 0.35 (Figure 5). This is what we expect, since our noise-adding process resamples from the Bernoulli(0.35) distribution, causing the overall fraction of maize to remain the same. The classical confidence interval stays constant because it only uses ground truth points. The map-based imputation confidence intervals are small because $N=100000$ is large, but they do not account for error in the map product. However, the imputation confidence intervals remain unbiased because the map still has true maize fraction close to 0.35 regardless of noise level.
When the noise probability is low, the PPI confidence intervals are significantly smaller than the classical intervals. This is because the error of the map product ($\hat{Y}-Y$) is small at the ground truth locations, so the variance of this error is also small, resulting in a small PPI standard error. As noise increases, the variance of the error of the map product increases; the PPI confidence intervals grow wider, eventually converging to the classical confidence interval at noise probability $p=1$. (At $p=1$, the map is complete noise and therefore does not provide any information for estimating the area beyond the information contained in the ground truth samples.) Consequently, the PPI effective sample size decreases from $n_{\text{effective}}=578$ when $p=0.1$ to $n_{\text{effective}}=100$ when $p=1$ (which matches the number of actual ground truth points $n=100$). Thus, the effectiveness of PPI is dependent on the quality of the map product. The Olofsson confidence intervals and effective sample sizes are nearly identical to those of PPI.
Figure 5: Simulation estimating maize fraction under different levels of map noise using $N=100000$ map product points and $n=100$ “ground truth” points. Olofsson and PPI results are overlapping. Left: Maize fraction 95% confidence intervals at different noise levels. Right: Maize fraction effective sample size at different noise levels. (At noise level $p$, about $p$ fraction of the points are resampled, while the remaining $1-p$ points are unchanged. Each resampled point is labeled as maize with probability 0.35 (the true maize fraction) and non-maize with probability 0.65.)
Figure 6: Simulation estimating maize fraction 95% confidence intervals under different levels of map bias using $N=100000$ map product points and $n=100$ “ground truth” points. Olofsson and PPI confidence intervals are overlapping. (Noise is fixed at 0.4 so about 40% of the points are resampled. At bias $p$, each resampled point is labeled as maize with probability $p$ and non-maize with probability $1-p$.)
The estimated maize fraction confidence intervals after adding bias to the map product are shown in Figure 6. Again, the classical confidence interval stays constant and remains centered at the true maize fraction because it only uses the ground truth points. As bias increases, the fraction of map product points labeled as “maize” increases. As a result, the map-based imputation estimates increase with the bias value, and generally do not remain centered at the true maize fraction of 0.35. The imputation confidence intervals are small because $N=100000$ is large, but they are unreliable because they do not correct for the bias in the map product. In contrast, the PPI estimates remain centered at the true maize fraction because PPI corrects for bias in the map product. The PPI confidence intervals are smaller than the classical intervals because the variance of the map product error $\hat{Y}-Y$ is small (since only about 40% of the map product points are resampled, regardless of bias level). Thus, PPI provides unbiased estimates with lower uncertainty than using only the ground truth points. Again, the Olofsson confidence intervals and effective sample sizes are nearly identical to those of PPI, as expected.
We also plot the maize fraction confidence intervals against model accuracy for both the noise and bias experiments (Figure 7). As noise increases, model accuracy decreases, so low model accuracy is associated with wide PPI and Olofsson confidence intervals. Thus, the left panel of Figure 7 resembles an inverted version of the left panel of Figure 5.
Figure 7: Simulation estimating maize fraction 95% confidence intervals at different map product accuracy levels obtained by varying noise (left) and bias (right), using $N=100000$ map product points and $n=100$ “ground truth” points. Olofsson and PPI confidence intervals are overlapping.
As bias increases and the fraction of map product points labeled as “maize” increases, model accuracy decreases linearly. (To see this, note that if each resampled point is labeled as maize with probability $p$, then an expected $0.35(p)+0.65(1-p)=0.65-0.3p$ fraction of the resampled points will match the ground truth.) Thus, the map-based imputation estimates of maize fraction decrease linearly as model accuracy increases. In the right panel of Figure 7, a model accuracy of 0.86 corresponds to an imputed maize fraction of 0.21, significantly lower than the true fraction of 0.35. This shows that even a map with high model accuracy can yield biased estimates under naive imputation.
3.2 Deforestation area in Brazilian Amazon
Problem Formulation
Next, we move on to real data and estimate the fraction of deforestation in the Brazilian Amazon from 2000 to 2015, comparing confidence interval widths across methods.
We create a binary deforestation map product using all $N=4.4\times 10^{9}$ pixels from the NASA Global Forest Cover Change (GFCC) map [townshend2016global] (at 30 m resolution) in the study area (Figure 8). The GFCC product is derived from Landsat 5 and Landsat 7 surface reflectance imagery, and includes maps for tree cover percentage in each pixel in five-year increments from 2000 to 2015. We consider a map pixel deforested if its canopy cover percentage satisfies $\text{canopy}_{2015}-\text{canopy}_{2000}\leq-25\%$ (Appendix B).
We use a ground truth Amazon deforestation and forest disturbance dataset created by [bullock2020satellite] using time-series analysis of Landsat imagery and high-resolution imagery from Google Earth between 1995 and 2017. The dataset records the years corresponding to deforestation, degradation, and natural disturbance events at each ground truth point. We use this to construct a binary ground truth dataset restricted to Brazil, in which we consider a point deforested if it experienced a deforestation event between 2000 and 2015. The original dataset is obtained by stratified random sampling, using a stratification map (also created by [bullock2020satellite]) that contains 7 strata related to deforestation status. To generate a classical estimate baseline, we construct a “uniformly random sampled” ground truth dataset by downsampling the points in each of the strata to match its area proportion. For a fair comparison, we also downsample the original stratified ground truth dataset to match the size of the “uniformly random” dataset. This results in two ground truth datasets of $n=1386$ points each, corresponding to uniform random sampling and stratified random sampling. Both ground truth datasets are used with PPI to obtain confidence intervals.
Figure 8: Brazilian Amazon deforestation area estimation datasets. Left: Stratification map from [bullock2020satellite]. Center: Binary map of deforestation from 2000-2015 derived from NASA Global Forest Cover Change. Right: $n=1386$ ground truth points from [bullock2020satellite], colored according to the stratum from which they are sampled.
Results
Figure 9: Brazilian Amazon deforestation 95% confidence intervals using $N=4.4\times 10^{9}$ map product points and $n=1386$ ground truth points. The Olofsson confusion matrix method, PPI, and stratification all result in smaller confidence intervals for estimated deforestation than classical estimation, with Stratified PPI having the best performance.
Each method results in a point estimate that between 5% to 6% of the Brazilian Amazon was deforested between 2000 and 2015. The Olofsson confusion matrix method, PPI, and stratification all result in smaller 95% confidence intervals for estimated deforestation than classical estimation (Figure 9). The Olofsson and PPI confidence intervals are fairly similar, with respective effective sample sizes that are $1.9$ and $1.8$ times the actual number of ground truth samples. The stratified and stratified PPI confidence intervals have the best performance, with effective sample sizes that outperform the classical method by a factor of about $2.4$. (Put differently, stratified PPI has effective sample size $3382$, so using the remotely sensed deforestation map added the equivalent of $3382-1386=1996$ ground truth samples.) Thus, the methods that use both map product and ground truth reduce uncertainty in our estimates, compared to using only ground truth. In particular, stratification is effective because the strata map is designed so that points within each stratum have low variance in deforestation status.
In this case, the naive imputation estimate — using only the map to estimate deforestation without any bias correction via ground truth labels — falls within the other confidence intervals, suggesting that the NASA Global Forest Cover Change map has small bias for estimating deforestation area. However, we emphasize that this would not have been clear without the ground truth data, and ground truth data is still necessary to check the reliability of the map product for deforestation analysis.
4 Estimating regression coefficients
In this section, we apply our uncertainty quantification methods to estimating regression coefficients for associative relationships between a remotely sensed outcome and multiple covariates. Regressions are commonly used in the environmental sciences and economics to identify relationships — associative or causal — between covariates and a binary outcome (logistic regression) or between covariates and a continuous outcome (linear regression). Logistic regression coefficients give the change in the log odds ratio of a binary outcome corresponding to a one-unit increase in a covariate. Linear regression coefficients give the change in the value of a continuous outcome corresponding to a one-unit increase in a covariate. Thus, we can interpret these coefficients as effect sizes of the covariates.
In particular, we estimate (1) logistic regression coefficients for Brazilian Amazon deforestation in 2000–2015, and (2) linear regression coefficients for US tree cover in 2021. We compare the classical method, naive imputation, PPI, stratification, and stratified PPI. Our methods only account for uncertainty in the remotely sensed target variable, while ignoring uncertainty in the covariates; we leave addressing the latter to future work.
4.1 Deforestation logistic regression
Problem Formulation
We estimate the logistic regression coefficients for the association between Brazilian Amazon deforestation in 2000–2015 and several covariates: distance from major roads, distance from major rivers, elevation, and slope (Figure 10). We use logistic regression because the outcome variable (deforested or non-deforested) is binary.
Given $n$ data points, the logistic regression model is
$\log\frac{p}{1-p}=\theta_{\text{road}}\textit{x}_{\text{road}}+\theta_{\text{river}}\textit{x}_{\text{river}}+\theta_{\text{elevation}}\textit{x}_{\text{elevation}}+\theta_{\text{slope}}\textit{x}_{\text{slope}}$
where $p$ is a $n\times 1$ vector of probabilities, $\textit{x}_{j}$ is a $n\times 1$ vector of covariate $j$, and $\theta_{j}$ is the scalar logistic regression coefficient for covariate $j$. We wish to estimate the $\theta_{j}$. A one-unit increase in covariate $j$ corresponds to a $\theta_{j}$ change in the log-odds of the outcome variable.
We use the same Brazilian Amazon deforestation datasets as the previous section, with $n=1386$ ground truth points and $N=963548$ uniformly randomly sampled map product points. For the covariates, we use a map of Brazilian federal roads in 2000 [de2021transport] and the WWF HydroSHEDS Free Flowing Rivers Network map [grill2019mapping] (we include all rivers of order at most 3) to compute the distance from each data point to the nearest major road and major river. We obtain elevation and slope from the NASA Digital Elevation Model [jpl2020] at 30 m resolution. Prior work has shown that deforestation in the Brazilian Amazon is more likely to occur near major roads and navigable rivers due to human settlement and economic activities along these transportation networks [laurance2002predictors, kirby2006future, rosa2013predictive]. Elevation and slope have also been associated with forest cover change; for example, deforestation is generally less likely on sloped terrain [sandel2013human]. We do not include a comprehensive list of covariates that could influence deforestation, because the primary purpose of this regression is to illustrate uncertainty quantification rather than explain deforestation.
Figure 10: Brazilian Amazon deforestation logistic regression datasets. We regress the deforestation binary target variable (left) against four continuous covariates (right): distance from roads, distance from rivers, elevation, and slope.
Results
The 95% confidence intervals for the four logistic regression coefficients are shown in Figure 11.
The general degree of uncertainty in estimating the coefficients from our non-imputation methods are as follows. The log odds of deforestation change by about .$-.016$ to .$-.010$ for each one-kilometer increase in distance from roads, .$-.008$ to .$-.001$ for each one-kilometer increase in distance from rivers, $-3$ to $0.8$ for each one-kilometer increase in elevation, and $-0.28$ to $-0.11$ for each one-degree increase in slope. Thus, deforestation is negatively associated with all covariates except for elevation. The elevation coefficient confidence intervals intersect with zero for all non-imputation methods except for stratified PPI, suggesting that there might be no association between elevation and deforestation after accounting for the other three variables.
The PPI methods result in the smallest confidence intervals, with effective sample size outperforming the classical method by a factor of 1.1 to 2.5 depending on the covariate. The most significant improvement is for the slope covariate. Unlike in our previous area estimation examples, stratification (i.e., weighted logistic regression with weights determined by the strata) is not effective and even reduces effective sample size relative to the classical method when estimating the coefficients for elevation and distance from rivers. Similarly, stratified PPI has slightly worse performance than PPI for all covariates except distance from roads. The strata were designed to minimize variance in deforestation status, which does not necessarily correspond to better estimates of regression coefficients for the covariates.
For slope and distance from rivers, the imputed coefficient computed from the map product falls outside the confidence intervals for all methods that use ground truth points. The imputed coefficient for slope is about 0, and the imputed coefficient for distance from rivers is about $-0.010$. In particular, for the slope coefficient, the methods that use ground truth points find an association that is statistically significantly different from zero. However, if we had used only the remote sensing map, we would not have discovered the negative relationship between slope and deforestation. Thus, using only the map product could result in misleading conclusions due to biases in the deforestation map product that are correlated with covariates (e.g., the map product overestimates deforestation at locations with high slope, or underestimates deforestation at locations with low slope).
Figure 11: Brazilian Amazon deforestation logistic regression coefficient 95% confidence intervals using $N=963548$ map product points and $n=1386$ ground truth points. The PPI methods result in the smallest confidence intervals, while stratification is not generally effective. For slope and distance from rivers, the map-imputed coefficient falls outside the confidence intervals for all methods that use ground truth points.
4.2 Tree cover linear regression
Problem Formulation
Lastly, we estimate linear regression coefficients for the association between 2021 tree cover percentage in the contiguous United States and several covariates: aridity index, elevation, and slope.
Given $n$ data points, the linear regression model is
$\mu=\theta_{\text{aridity}}\textit{x}_{\text{aridity}}+\theta_{\text{elevation}}\textit{x}_{\text{elevation}}+\theta_{\text{slope}}\textit{x}_{\text{slope}}$
where $\mu$ is a $n\times 1$ vector of predictions, $\textit{x}_{j}$ is a $n\times 1$ vector of covariate $j$, and $\theta_{j}$ is the scalar linear regression coefficient for covariate $j$. We wish to estimate the $\theta_{j}$. A one-unit increase in covariate $j$ corresponds to a $\theta_{j}$ change in the outcome variable.
We use $N=983238$ map product points sampled uniformly at random from the 2021 USFS Tree Canopy Cover (TCC) product [forest2023usfs] at 30 m resolution. The TCC product is derived from Landsat and Sentinel-2 top-of-atmosphere reflectance data, USDA NASS CDL data, and terrain data. Its tree cover percentage predictions are obtained from random forest models trained on reference data from human-labeled high-resolution imagery. For ground truth, we manually label $n=983$ high-resolution satellite images from Google Maps for tree cover percentage, at locations sampled uniformly at random from the study area. Because the source of ground truth is from a uniform random sample, we do not use any stratified methods.
For the covariates, we use a map of the average Global Aridity Index from 1970-2000 (at 30 arc-seconds resolution) [trabucco2019global] and the NASA Digital Elevation Model [jpl2020] (at 30 m resolution). The aridity index is defined as a ratio between precipitation and evapo-transpiration, and is lower for more arid areas; its value ranges from 0.02 to 6.55 in our dataset. Increased aridity is associated with lower tree cover (and decreased vegetation in general) [bastin2017extent, williams2010forest, berdugo2020global]. Elevation and slope have also been associated with tree cover globally; there is generally more tree cover at lower elevations (e.g., due to higher temperatures) [mayor2017elevation] and on sloped terrain [sandel2013human]. We do not include a comprehensive list of covariates that could influence tree cover, because the primary purpose of this regression is to illustrate uncertainty quantification rather than explain tree cover.
Figure 12: US tree cover linear regression datasets. We regress the tree cover percentage target variable (left) against three covariates (right): aridity index, elevation, and slope. We use the USFS Tree Canopy Cover map product and $n=983$ uniformly randomly sampled ground truth points.
Results
Figure 13: US tree cover linear regression coefficient 95% confidence intervals using $N=983238$ map product points and $n=983$ ground truth points. PPI outperforms the classical method. For elevation, the map-imputed coefficient falls outside the PPI confidence interval.
The 95% confidence intervals for the four linear regression coefficients are shown in Figure 13.
The PPI confidence intervals for coefficients are as follows. Tree cover percentage increases by 35 to 40 percentage points for each one-unit increase in aridity index, decreases by 2.4 to 4.8 percentage points for each one-kilometer increase in elevation, and increases by 0.9 to 1.3 percentage points for each one-degree increase in slope. The PPI effective sample size outperforms the classical method by a factor of 3.3 for aridity index, 2.4 for elevation, and 3.6 for slope. As in the other use cases, PPI uses the map product along with the ground truth to reduce uncertainty in the coefficient estimates compared to only using ground truth.
For elevation, the imputed coefficient computed from the map product ($-5.3$) falls outside the PPI confidence interval. This implies that there is bias in the tree cover map product’s association with elevation (e.g., the map product overestimates tree cover at low elevation locations or underestimates tree cover at high elevation locations). Similar to the deforestation logistic regression use case, using only the map product could result in misleading conclusions. If we had used only the remotely sensed tree cover predictions, we would have overestimated the magnitude of the relationship between elevation and tree cover due to map errors.
5 Discussion
Using remote sensing-based maps to draw scientific inferences requires uncertainty quantification and propagation of uncertainty to downstream analyses — otherwise, we risk making, in the words of [mcroberts2011satellite], just pretty pictures. In this work, we applied several methods that use remote sensing map products, along with a small amount of ground truth data, to estimate area and regression coefficients while correctly taking into account the errors in the map products. We find that stratified sampling, prediction-powered inference, and (for area estimation) the Olofsson method result in lower uncertainty than classical estimates using only ground truth data sampled. Unlike naive imputation using the map product alone, these estimates are guaranteed to be statistically valid.
Bias is the main issue for naive imputation.
Our experiments illustrated that a naive imputation of area or regression coefficients from a map without bias correction can yield incorrect inferences, and we differentiated between the influence of noise versus bias. In our simulation estimating maize fraction, naive imputation deviated from the true maize fraction of 0.35 as the bias of the crop type map increased. Furthermore, accuracy is not necessarily a good proxy for bias; even at seemingly high map accuracies ($\text{accuracy}>0.85$), naive imputation estimated maize fraction at 0.21 (a 40% underestimate).
In our experiments on real data, imputation for deforestation in the Brazilian Amazon fell within stratified and PPI confidence intervals, suggesting an overall low-bias map product. However, imputation of the logistic regression coefficients on distance from rivers and slope fell outside the intervals computed through the classical, stratified, PPI, and stratified PPI methods. Naive imputation would have concluded that distance from rivers is significantly more predictive of deforestation than it actually is, and that slope has no relationship to deforestation when it actually does. Thus, even when a map product is unbiased for the average outcome, it can be biased for the purpose of estimating regression coefficients. Similarly, imputation of the elevation coefficient in our linear regression for tree cover also overestimates the effect of elevation compared to PPI.
Effect of map accuracy.
How much stratified sampling, PPI, and confusion matrix-based methods reduce uncertainty relative to classical estimates depends on the quality of the map product. In stratification, the strata map should be designed such that the variance of the quantity of interest is low within each stratum. In PPI (and the Olofsson method), the variance of the difference between the map product predictions $\hat{Y}$ and the ground truth values $Y$ should be low. In our maize fraction experiments, a highly accurate map ($\text{accuracy}=0.96$) was equivalent to having nearly 6×6$\times$6 × as much ground truth data, whereas a low-accuracy map ($\text{accuracy}=0.73$) only increased the effective sample size by $1.2$.
Stratification vs. PPI vs. Olofsson.
Comparing stratified sampling, PPI, and the Olofsson method, which estimator yields the smallest confidence intervals depends on how the strata were designed and the quality of the map product. Recall Table 2, which summarizes the conditions under which each method is most appropriate. Empirically, when estimating deforestation fraction in the Brazilian Amazon, we observed the smallest confidence intervals when stratification was used alone or in conjunction with PPI. This is a result of well-designed strata based on a remote sensing map of forest disturbance categories. PPI and the Olofsson method yielded comparable interval widths to each other. This is expected, as the Olofsson method yields similar confidence intervals as PPI when $N$ is large (Appendix A).
The picture changes, however, when we reuse the same stratified samples to estimate regression coefficients in regressions that the strata were not designed for. When we associated deforestation with covariates in a logistic regression, PPI and stratified PPI resulted in narrower confidence intervals than the stratified estimator across all 4 coefficients (distance from road, distance from river, elevation, and slope). The worse performance of the stratified estimator is because the strata were designed to minimize the variance of deforestation within each strata, which does not necessarily correspond to better regression estimation. Indeed, the stratified estimator was actually worse than the classical estimator for certain coefficients.
An advantage of PPI compared to stratification, then, is its flexibility. While stratification requires us to choose a map product before sampling, PPI can be applied to any map product and any randomly sampled ground truth dataset. Given the expensive nature of ground truth data collection, sampling without choosing a strata map and then using PPI to obtain valid confidence intervals may result in the best performance across a variety of downstream use cases. Or, if stratified samples are already available from area estimation, they can be combined with PPI to estimate regression coefficients.
Recommendations for map producers and users.
Ultimately, the best methods combine map products with ground truth to yield the narrowest intervals (highest certainty). We recommend that producers of machine learning-generated remote sensing maps use a holdout ground truth dataset for uncertainty quantification. The criterion for this holdout set is that the model, including during hyperparameter tuning, was not trained on it; a standard test set would be adequate. Map producers should release this ground truth data alongside their map product to make their map suitable for downstream scientific inferences. In the absence of a provided holdout set (currently, releasing ground truth data is not the norm), we recommend that map users generate their own ground truth data for the variable of interest if possible.
Limitations.
All of the estimation methods described in this paper (except for naive imputation) are limited to areas in which ground truth data is available. However, this would be true for any assumption-free, data-driven method that quantifies uncertainty and corrects for map bias. Furthermore, we can only ensure statistical validity of our estimates when the ground truth points are randomly sampled. Together, these attributes mean that such methods may be especially straightforward to use when ground truth data can be generated de novo, such as from human inspection of available satellite imagery, and more challenging to use when ground truth data requires bespoke surveys. Another limitation is that our estimates only quantify uncertainty in the target variable map product, and do not account for noise in either the ground truth data or (in the regression use cases) the covariates. We leave these investigations to future work.
6 Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number(s) DE-SC0023112. This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
[…]
Appendix A Mathematical comparison of PPI and the method of Olofsson
We show that when the number of map product points $N$ is large, the PPI and Olofsson area estimation confidence intervals are similar. In particular, we show (1) that they have approximately the same length when the number of samples is large and (2) they are centered on the same point. These two facts together provide a mathematical explanation for why PPI and the Olofsson method approximately agree in all of our experiments.
We begin with notation. Estimating the area proportion of a land cover class in a given region is equivalent to estimating the mean of a binary variable $Y$, where $Y=1$ for points in the class and $Y=0$ for points outside the class. Suppose we have $n$ uniformly random sampled ground truth points $Y_{1},Y_{2},\dots Y_{n}$ and $N$ map product points $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{N}$. For all $0\leq i,j\leq 1$, let $n_{ij}$ denote the number of ground truth points that are in both map class $i$ and ground truth class $j$. Let $n_{i\cdot}=n_{i0}+n_{i1}$ denote the total number of ground truth points in map class $i$, and $n_{\cdot i}=n_{0i}+n_{1i}$ denote the total number of ground truth points in ground truth class $i$. Furthermore, let $A_{i}$ denote the fraction of the map product that is in class $i$. We have $A_{0}+A_{1}=1$. The map-imputed area estimate, which we will correct for bias, is $A_{1}$. We will use the approximations $A_{0}\approx\frac{n_{0\cdot}}{n}$ and $A_{1}\approx\frac{n_{1\cdot}}{n}$. This is a reasonable approximation since the $n$ ground truth locations are sampled uniformly at random.
With this notation, we have the following,
$\sum_{i=1}^{n}Y_{i}$ $=n_{\cdot 1}$
$\sum_{i=1}^{n}\hat{Y}_{i}$ $=n_{1\cdot}$
$\mathrm{Var}_{n}(Y)$ $=\frac{n_{\cdot 0}n_{\cdot 1}}{n^{2}}$
$\mathrm{Var}_{N}(\hat{Y})$ $=A_{0}A_{1}\approx\frac{n_{0\cdot}n_{1\cdot}}{n^{2}}\text{ and}$
$\mathrm{Cov}_{n}(Y,\hat{Y})$ $=\frac{n_{00}n_{11}-n_{01}n_{10}}{n^{2}}.$
Standard errors.
We now turn to the PPI standard error. We have
$SE_{\text{PPI}}$ $\rightarrow\sqrt{\frac{1}{n}\mathrm{Var}(Y-\lambda\hat{Y})}$ since $N\gg n$
$=\sqrt{\frac{1}{n}\left(\mathrm{Var}(Y)+\lambda^{2}\mathrm{Var}(\hat{Y})-2\lambda\mathrm{Cov}(Y,\hat{Y})\right)}$ expanding out $\mathrm{Var}(Y-\lambda\hat{Y})$
$\rightarrow\sqrt{\frac{1}{n}\left(\mathrm{Var}(Y)-\frac{\mathrm{Cov}(Y,\hat{Y})^{2}}{\mathrm{Var}(\hat{Y})}\right)}$ plugging in $\lambda$
$\approx\sqrt{\frac{1}{n}\left(\frac{n_{\cdot 0}n_{\cdot 1}}{n^{2}}-\frac{(n_{00}n_{11}-n_{01}n_{10})^{2}}{n^{2}n_{0\cdot}n_{1\cdot}}\right)}$ plugging in $\mathrm{Var}(Y)$, $\mathrm{Var}(\hat{Y})$, $\mathrm{Cov}(Y,\hat{Y})$
$=\sqrt{\frac{1}{n^{2}n_{0\cdot}n_{1\cdot}}(n_{01}n_{00}n_{10}+n_{01}n_{00}n_{11}+n_{11}n_{10}n_{00}+n_{11}n_{10}n_{01})}$ using algebraic manipulation.
Above, we use the fact that as $n,N\rightarrow\infty$, the PPI tuning parameter (from Example 6.1 of [angelopoulos2023ppi++]) behaves as
$\lambda\rightarrow\frac{\mathrm{Cov}_{n}(Y,\hat{Y})}{\mathrm{Var}_{N}(\hat{Y})}\approx\frac{n_{00}n_{11}-n_{01}n_{10}}{n_{0\cdot}n_{1\cdot}}.$
We compare this with the Olofsson standard error
$SE_{\text{Olofsson}}$ $=\sqrt{A_{0}^{2}\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+A_{1}^{2}\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$
$\approx\sqrt{\left(\frac{n_{0\cdot}}{n}\right)^{2}\times\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+\left(\frac{n_{1\cdot}}{n}\right)^{2}\times\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$ plugging in $A_{0}$ and $A_{1}$
$=\sqrt{\frac{1}{n^{2}n_{0\cdot}n_{1\cdot}}(n_{01}n_{00}n_{10}+n_{01}n_{00}n_{11}+n_{11}n_{10}n_{00}+n_{11}n_{10}n_{01})}$ using algebraic manipulation.
The expressions we arrive at for PPI and the method of Olofsson agree, indicating that for large sample sizes the standard errors are approximately equal. This implies that the confidence intervals from the two methods will have the same width.
Point estimates.
Next, we show that the PPI and Olofsson confidence intervals are centered around the same point. Turning first to the PPI point estimate
$\hat{\theta}_{\text{PPI}}$ $=\lambda A_{1}-\frac{1}{n}\sum_{i=1}^{n}(\lambda\hat{Y}_{i}-Y_{i})$
$\approx\frac{\lambda n_{1\cdot}}{n}-\frac{\lambda n_{1\cdot}-n_{\cdot 1}}{n}$ plugging in $A_{1}$, $\sum_{i=1}^{n}Y_{i}$, and $\sum_{i=1}^{n}\hat{Y}_{i}$
$=\frac{n_{\cdot 1}}{n}$
The Olofsson area point estimate is
$\hat{\theta}_{\text{Olofsson}}$ $=A_{0}\frac{n_{01}}{n_{0\cdot}}+A_{1}\frac{n_{11}}{n_{1\cdot}}$
$\approx\frac{n_{0\cdot}}{n}\times\frac{n_{01}}{n_{0\cdot}}+\frac{n_{1\cdot}}{n}\times\frac{n_{11}}{n_{1\cdot}}$ plugging in $A_{0}$ and $A_{1}$
$=\frac{n_{\cdot 1}}{n}$
We see that the final expressions are equal, which means the point estimates are approximately equal.
Appendix B Binarizing the NASA Global Forest Cover Change map product for deforestation
In the Brazilian Amazon deforestation use cases, the ground truth points from [bullock2020satellite] have binary deforestation values ($Y=1$ if deforested and $Y=0$ if not deforested between 2000-2015), while the NASA GFCC map product gives canopy cover percentages over time at each pixel.
In order to binarize the continuous map product data, we classify a map pixel as “deforested” if its canopy cover percentage satisfies $\text{canopy}_{2015}-\text{canopy}_{2000}\leq-25\%$. We choose the $-25\%$ threshold by visually comparing the map-derived canopy cover change distributions for deforested and non-deforested ground truth points (Figure 14). Most of the deforested ground truth points have a map canopy cover decrease of more than $25\%$, while most of the non-deforested ground truth points have either a map canopy cover increase, no map canopy cover change, or a map canopy cover decrease of less than $25\%$.
Figure 14: NASA GFCC map canopy cover change (2000-2015) distributions for non-deforested and deforested ground truth over $n=1386$ uniformly random sampled points. The distributions are visualized using kernel density estimation. We choose a -25% deforestation threshold (shown in red) to binarize the map product.

### v3

@@
 However, the quality of map products varies, and — because maps are outputs of complex machine learning algorithms that take in a variety of remotely sensed variables as inputs — errors are difficult to characterize.
-Without capturing the biases that may be present, naive calculations of population-level estimates from such maps are statistically invalid.
-In this paper, we compare several uncertainty quantification methods — stratification, Olofsson area estimation method, and prediction-powered inference — that combine a small amount of randomly sampled ground truth data with large-scale remote sensing map products to generate statistically valid estimates.
-Applying these methods across four remote sensing use cases in area and regression coefficient estimation, we find that they result in estimates that are more reliable than naively using the map product as if it were 100% accurate and have lower uncertainty than using only the ground truth and ignoring the map product.
+Thus, population-level estimates from such maps may be biased.
+In this paper, we compare several uncertainty quantification methods — stratified estimator, post-stratified estimator, and prediction-powered inference — that combine a small amount of randomly sampled ground truth data with large-scale remote sensing map products to generate unbiased estimates.
+Applying these methods across four remote sensing use cases in area and regression coefficient estimation, we find that they result in estimates that are more reliable than using the map product as if it were 100% accurate and have lower uncertainty than using only the ground truth and ignoring the map product.
 Prediction-powered inference uses ground truth data to correct for bias in the map product estimate and (unlike stratification) does not require us to choose a map product before sampling.
 This is the first work to (1) apply prediction-powered inference to remote sensing estimation tasks, and (2) perform uncertainty quantification on remote sensing regression coefficients without assumptions on the structure of map product errors.
-To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a holdout ground truth dataset to be used for calibration in uncertainty quantification alongside their maps.
+To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a randomly sampled holdout ground truth dataset to be used for calibration in uncertainty quantification alongside their maps.
 Data and code are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
@@
 Map products created by training machine learning models on remotely sensed data allow researchers to cheaply access predictions for a variable of interest at a high resolution over large regions.
-However, the quality of map products varies, and resulting population-level estimates of quantities of interest may be biased and statistically invalid [jain2020benefits, bastin2017extent, venter2024uncertainty].
+However, the quality of map products varies, and resulting population-level estimates of quantities of interest may be biased [jain2020benefits, bastin2017extent, venter2024uncertainty].
 For example, [bastin2017extent] showed that human labeling of high-resolution imagery results in an estimate of global forest extent in dryland biomes that is 40% to 47% higher than previous estimates from remotely sensed forest cover map products.
@@
 This generally reduces uncertainty in the estimates compared to using ground truth alone, since map products contain a large amount of data that, while possibly biased, still give useful information about the true values we wish to measure.
-Figure 1: Motivating example: map-only inference is statistically invalid.
+Figure 1: Motivating example: map-only estimate is biased.
 The 95% confidence interval for logistic regression coefficient relating Brazilian Amazon deforestation to slope using $N=963548$ remote sensing map points and $n=1386$ ground truth points.
-The imputed slope coefficient from the map alone falls outside the confidence intervals for all methods that use ground truth points.
+The map-only slope coefficient estimate falls outside the confidence intervals for all methods that use ground truth points.
 A motivating example.
-In this work, we discuss methods that correct map products with ground-truth data to give statistically-valid confidence intervals.
+In this work, we discuss methods that correct map products with ground-truth data to give unbiased confidence intervals.
 We give a brief preview here.
@@
 See Section 4.1 for a complete description of the experimental setup.
-The “classical” estimate using only ground truth points is statistically valid but has the widest confidence interval, failing to take advantage of existing map products.
-Meanwhile, methods that use a map product along with the ground truth points produce an estimate with lower uncertainty while still being valid.
-Notably, the slope coefficient computed from the map product alone (“naive imputation”) is approximately zero, which falls outside the confidence intervals for all methods that use ground truth points.
+The “classical” estimate using only ground truth points is unbiased but has the widest confidence interval, failing to take advantage of existing map products.
+Meanwhile, methods that use a map product along with the ground truth points produce an estimate with lower uncertainty while still being unbiased.
+Notably, the slope coefficient computed from the map product alone is approximately zero, which falls outside the confidence intervals for all methods that use ground truth points.
 That is, if we had used only the map product, we would not have discovered the statistically significant negative relationship between slope and deforestation.
@@
 Figure 2: Overview of uncertainty quantification using ground truth data and remote sensing map products.
-Olofsson area estimation, stratification, and prediction powered inference methods use a small amount of ground truth data along with a map product to produce confidence intervals for the quantity of interest.
-We also compare these methods against classical uncertainty quantification (using only ground truth) and naive imputation (using only the map product).
+Stratified, post-stratified, and prediction powered inference estimators use a small amount of ground truth data along with a map product to produce confidence intervals for the quantity of interest.
+We also compare these methods against classical uncertainty quantification (using only ground truth) and map-only estimators.
 Our contribution.
-In this paper, we consider uncertainty quantification methods that combine a small number of randomly sampled ground truth data ($n$ points) with large-scale remote sensing map products ($N$ points, where $N$ is much larger than $n$) in order to generate statistically valid estimates with low uncertainty (Figure 2).
-Stratification uses map products to determine a sampling strategy for ground truth points that reduces the variance of estimates; it has been used in the remote sensing community for decades [hay1979sampling, fitzpatrick1981comparison, stehman1998design, olofsson2014good].
+In this paper, we consider uncertainty quantification methods that combine a small number of randomly sampled ground truth data ($n$ points) with large-scale remote sensing map products ($N$ points, where $N$ is much larger than $n$) in order to generate unbiased estimates with low uncertainty (Figure 2).
+We also compare these methods against the classical estimator (using only ground truth) and map-only estimator (using only a map product).
+We consider the following three methods:
+  • Stratified estimator uses map products to determine a sampling strategy for ground truth points that reduces the variance of estimates.
+This method has been used in the remote sensing community for decades [hay1979sampling, fitzpatrick1981comparison, stehman1998design, olofsson2014good].
+  • Post-stratified area estimator uses the confusion matrix between ground truth and map product land cover labels to correct for misclassification bias in the map product, when estimating the area of a land cover class.
+This method is also commonly used in the remote sensing community [card1982using, stehman2013estimating, olofsson2013making].
   • Prediction-powered inference (PPI) [angelopoulos2023prediction, angelopoulos2023ppi++] is a recently developed uncertainty quantification method that uses a small set of ground truth points to calibrate and correct for bias in a much larger set of ML-generated predictions, resulting in more precise estimates than using the ground truth points alone.
-We compare stratification and prediction-powered inference against classical estimation (using only ground truth) and naive imputation (using only a map product).
 We apply these methods to two of the most common scientific uses of remote sensing maps: area estimation and regression coefficient estimation.
@@
 The methods applied to area estimation also apply to estimating sums or averages of continuous variables, such as total carbon stock.
+A map-only estimator for area is known as a pixel-counting estimator.
 Regression coefficient estimation refers to estimating linear or logistic regression coefficients for causal or associative relationships between two or more environmental variables.
@@
 Code implementations of the use cases are linked in the rightmost column.
-Combining ground truth data with map products results in statistically valid estimates that are more reliable than naive imputation and have lower uncertainty than classical estimates.
-As we will show, naive imputation from maps without bias correction can result in completely incorrect area and regression coefficient estimates.
-To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a holdout ground truth dataset (such as the usual test set) to be used for calibration in uncertainty quantification alongside their maps, or, if that is not done, for map users to generate their own holdout set for calibration, if possible (e.g., from visual inspection of remote sensing data).
+Combining ground truth data with map products results in unbiased estimates that are more reliable than map-only estimates and have lower uncertainty than classical estimates.
+As we will show, map-only estimators without bias correction can result in completely incorrect area and regression coefficient estimates.
+To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a (simple or stratified) randomly sampled holdout ground truth dataset to be used for calibration in uncertainty quantification alongside their maps, or, if that is not done, for map users to generate their own randomly sampled holdout set for calibration, if possible (e.g., from visual inspection of remote sensing data).
 Related work.
-Previously, Olofsson et al.
-[olofsson2013making] described an area estimation method that uses the confusion matrix between land cover ground truth and map product points to correct for misclassification bias in the map product.
-The Olofsson method is commonly used in the remote sensing community to generate confidence intervals for area estimates.
-However, it does not apply to regression coefficient estimation tasks.
-One contribution of the present work is a comparison of this method and PPI.
+The remote sensing community commonly uses a post-stratified estimator to generate confidence intervals for area estimates.
+This estimator uses the confusion matrix between land cover ground truth and map product points to correct for misclassification bias in the map product [card1982using, stehman2013estimating, olofsson2013making].
+This method only applies to area estimation, not regression coefficient estimation tasks.
+One contribution of the present work is a comparison of the post-stratified area estimator and PPI.
 We find both experimentally and mathematically that these two methods produce similar outputs for area estimation.
-Other methods have been developed to reduce bias in regression coefficient estimates from remote sensing data.
+For other methods, [gallego2004remote] provides a survey of area estimators that combine ground truth and map product data, including regression, ratio, and calibration estimators.
+Methods have also been developed to reduce bias in regression coefficient estimates from remote sensing data.
 For example, [alix2023remotely] estimates the effect of a conservation program on deforestation in Mexico.
@@
 Note that this method requires an assumption about the error structure of the remotely-sensed variables, whereas the methods investigated herein do not.
-Another approach to uncertainty quantification is conformal prediction, which generates statistically valid confidence sets for machine learning predictions at individual data points.
+Another approach to uncertainty quantification is conformal prediction, which generates confidence sets for machine learning predictions at individual data points.
 For example, [valle2023quantifying] applies conformal prediction to quantify pixel-level uncertainty for a land cover/land use map product in the Brazilian Amazon, outputting a 90% predictive set of land cover classes for each pixel (i.e., the set contains the true land cover class 90% of the time).
@@
 We formally describe uncertainty quantification methods that generate confidence intervals with remote sensing variables.
-As two baselines, we first consider methods that use only the ground-truth data (which is valid but loses information) or only use the remote-sensing data (which is not statistically valid).
-We then turn to methods that use both sources of data to produce statistically-valid confidence intervals.
+As two baselines, we first consider methods that use only the ground-truth data (which is unbiased but loses information) or only use the remote-sensing data (which may be biased).
+We then turn to methods that use both sources of data to produce unbiased confidence intervals.
 See Table 2 for a summary.
-Method Data Used Statistically Valid? Caveats
+Method Data Used Unbiased? Caveats
 Map Ground Truth
-Naive imputation
+Map-only
 Classical
-Stratification Ground truth points sampled
+Stratified Ground truth points sampled
 using stratified map
-Olofsson Area estimation only
+Post-stratified Area estimation only
 PPI
@@
 This method is conservative — if the size $n$ of the ground truth dataset is small, the classical confidence interval will be large.
-Naive imputation: use only the remote sensing map.
+Map-only estimator: use only the remote sensing map.
 If we have a remote sensing data product with a large number $N$ of predictions $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{N}$ over the area of interest, we could treat these predictions like ground truth points and use the the formula above with $\hat{Y}$ in place of $Y$.
 This would produce a much smaller confidence interval, since we would use $N$ instead of $n$ in the denominator of the standard error.
-However, this naive imputation approach is not statistically valid, as it does not account for error or uncertainty in the model used to generate the map product predictions.
-2.2 Stratification
+However, this map-only estimator approach may result in a biased estimate, and the confidence interval is not guaranteed to have a 95% coverage rate of the true parameter of interest.
+This is because the map-only estimator does not account for error or uncertainty in the model used to generate the map product predictions.
+2.2 Stratified estimator
 Stratified random sampling uses remotely sensed maps to determine a sampling strategy for ground truth points that reduces the variance of estimates.
-Stratification is commonly used in the remote sensing community.
+The stratified estimator is commonly used in the remote sensing community.
 We first use a map product to partition the area of interest into $K$ strata (sub-areas) such that the target variable is predicted to have low variance within each stratum.
@@
 $C_{\text{strat}}=(\hat{\theta}_{\text{strat}}-1.96\cdot SE_{\text{strat}},\ \hat{\theta}_{\text{strat}}+1.96\cdot SE_{\text{strat}}).$
-The advantage of stratification with an accurate map and well-chosen sampling strategy is that the variance within each stratum will be small, which results in a smaller overall confidence interval compared to the classical method.
+The advantage of stratified estimation with an accurate map and well-chosen sampling strategy is that the variance within each stratum will be small, which results in a smaller overall confidence interval compared to the classical method.
 If the map predictions are inaccurate, then the ground truth points likely have higher variance within each map stratum, resulting in a larger confidence interval.
@@
 As a result, our stratified confidence interval for deforestation area is smaller than the classical confidence interval.
-Figure 3: Stratified random sample is a standard approach for estimation with remote sensing maps.
+Figure 3: Stratified estimation is a standard approach with remote sensing maps.
 This example shows a map of the Brazilian Amazon used for stratification.
@@
 The PPI point estimate is
-$\hat{\theta}_{\text{PPI}}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}}_{\text{naive imputation}}-\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})}_{\text{bias correction}}.$
+$\hat{\theta}_{\text{PPI}}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}}_{\text{map-only estimate}}-\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})}_{\text{bias correction}}.$
 The PPI standard error is
@@
 We sketch the main idea here and refer the reader to [angelopoulos2023ppi++, angelopoulos2023prediction] for a full exposition.
-In brief, PPI begins like naive imputation by using a large set of map predictions $\hat{Y}^{\prime}_{1},\hat{Y}^{\prime}_{2},\dots\hat{Y}^{\prime}_{N}$ to compute a measure of fit $m_{\theta}$.
+In brief, PPI begins like a map-only estimator by using a large set of map predictions $\hat{Y}^{\prime}_{1},\hat{Y}^{\prime}_{2},\dots\hat{Y}^{\prime}_{N}$ to compute a measure of fit $m_{\theta}$.
 Then, PPI computes a bias correction $\Delta_{\theta}$ by comparing the ground truth data $Y_{i}$ to the predictions $\hat{Y}_{i}$.
@@
 — see Table 3.
-Lastly, the naive $m_{\theta}$ and the bias correction $\Delta_{\theta}$ are combined to construct the PPI 95% confidence interval
+Lastly, the map-only $m_{\theta}$ and the bias correction $\Delta_{\theta}$ are combined to construct the PPI 95% confidence interval
 $C_{\text{PPI}}=\{\theta:|m_{\theta}+\Delta_{\theta}|\leq w_{\theta}\}$
@@
 The software includes a tuning parameter $\lambda$ that we set using the map product points and ground-truth points.
-The confidence intervals are statistically valid even with this data-driven tuning [angelopoulos2023ppi++].
-2.4 Olofsson et al. (2013) area estimation
-Olofsson et al.
-[olofsson2013making] introduces an uncertainty quantification method that uses the confusion matrix between ground truth and map product land cover labels to estimate the area of a land cover class.
-This method is commonly used in the remote sensing community.
-Like PPI, the Olofsson method computes confidence intervals by using ground truth points to correct for bias in the map product estimates.
-In fact, we show mathematically in Appendix A that the two methods result in similar confidence intervals as the number of map product points $N$ grows large.
-However, unlike PPI, the Olofsson method only applies to area estimation but not regression coefficient estimation.
+The confidence intervals are unbiased even with this data-driven tuning [angelopoulos2023ppi++].
+2.4 Post-stratified area estimator
+The post-stratified area estimator, commonly used in the remote sensing community, utilizes the confusion matrix between ground truth and map product land cover labels to estimate the area of a land cover class [card1982using, stehman2013estimating, olofsson2013making].
+Like PPI, the post-stratified area estimator computes confidence intervals by using ground truth points to correct for bias in the map product estimates.
+This can be viewed as a correction that uses the map errors relative to the ground truth points to modify the map-only area estimates.
+We show mathematically in Appendix A that PPI and the post-stratified area estimator result in similar confidence intervals as the number of map product points $N$ grows large.
+However, unlike PPI, the post-stratified area estimator only applies to area estimation but not regression coefficient estimation.
 Turning to the method, suppose we have a map product with $K$ land cover classes, as well as $n$ randomly sampled ground truth points.
-The Olofsson method constructs the $K\times K$ misclassification confusion matrix between ground truth points and map product predictions at the ground truth point locations.
+The post-stratified area estimator constructs the $K\times K$ misclassification confusion matrix between ground truth points and map product predictions at the ground truth point locations.
 For each $1\leq i\leq K$ and $1\leq j\leq K$, let $n_{ij}$ denote the number of points that are in map class $i$ and ground truth class $j$.
 For each class $1\leq i\leq K$, let $10<A_{i}<$ 1 denote the proportion of the total map area taken up by the map class.
-Then, the area estimate for the area of class $j$ is
-$\hat{\theta}_{j,\text{Olofsson}}=\sum_{i=1}^{K}A_{i}\frac{n_{ij}}{n_{i\cdot}}$
+Then, the post-stratified area estimate for the area of class $j$ is
+$\hat{\theta}_{j,\text{post}}=\sum_{i=1}^{K}A_{i}\frac{n_{ij}}{n_{i\cdot}}$
 with corresponding standard error
-$SE_{j,\text{Olofsson}}=\sqrt{\sum_{i=1}^{K}A_{i}^{2}\frac{\frac{n_{ij}}{n_{i\cdot}}(1-\frac{n_{ij}}{n_{i\cdot}})}{n_{i\cdot}}}.$
+$SE_{j,\text{post}}=\sqrt{\sum_{i=1}^{K}A_{i}^{2}\frac{\frac{n_{ij}}{n_{i\cdot}}(1-\frac{n_{ij}}{n_{i\cdot}})}{n_{i\cdot}}}.$
 The 95% confidence interval for the area of class $j$ is
-$C_{j,\text{Olofsson}}=(\hat{\theta}_{j,\text{Olofsson}}-1.96\cdot SE_{j,\text{Olofsson}},\ \hat{\theta}_{j,\text{Olofsson}}+1.96\cdot SE_{j,\text{Olofsson}}).$
+$C_{j,\text{post}}=(\hat{\theta}_{j,\text{post}}-1.96\cdot SE_{j,\text{post}},\ \hat{\theta}_{j,\text{post}}+1.96\cdot SE_{j,\text{post}}).$
 Area estimation is a special case of mean estimation in which we estimate the mean of a binary variable that is $1$ for all points in class $j$ and 00 for all points outside of class $j$.
-Thus, we can also perform area estimation using methods such as PPI and stratification.
-If the $n$ ground truth points are obtained by stratified random sampling using the $K$ map land cover classes as the strata, the Olofsson confidence interval is identical to the stratified confidence interval.
+Thus, we can also perform area estimation using methods such as PPI and stratified estimation.
+If the $n$ ground truth points are obtained by stratified random sampling using the $K$ map land cover classes as the strata, the post-stratified confidence interval is identical to the stratified confidence interval.
 2.5 Evaluating performance using effective sample size
@@
 We then use real ground truth data and real map products developed by researchers to estimate change in tree cover in the Brazilian Amazon in 2000–2015.
-We compare the classical method, naive imputation, stratification, PPI, stratified PPI, and the method of Olofsson et al.
-(2013) [olofsson2013making].
+We compare the classical, map-only (pixel-counting), stratified, post-stratified, PPI, and stratified PPI estimators.
 In each use case, we use binary data where each point has value 1 if it belongs to the land cover class of interest (e.g., cropland or deforested land), and value 0 otherwise.
@@
 We then experiment with degrading the map product quality as follows.
-We systematically add noise or bias to the map product points while keeping the ground truth points constant, and then compare our maize fraction estimates across naive imputation, classical estimation, and PPI.
+We systematically add noise or bias to the map product points while keeping the ground truth points constant, and then compare our maize fraction estimates across the map-only (pixel-counting), classical, and PPI estimators.
 We average results over 100 trials.
@@
 The classical confidence interval stays constant because it only uses ground truth points.
-The map-based imputation confidence intervals are small because $N=100000$ is large, but they do not account for error in the map product.
-However, the imputation confidence intervals remain unbiased because the map still has true maize fraction close to 0.35 regardless of noise level.
+The map-only confidence intervals are small because $N=100000$ is large, but they do not account for error in the map product.
+However, the map-only confidence intervals remain unbiased because the map still has true maize fraction close to 0.35 regardless of noise level.
 When the noise probability is low, the PPI confidence intervals are significantly smaller than the classical intervals.
@@
 Thus, the effectiveness of PPI is dependent on the quality of the map product.
-The Olofsson confidence intervals and effective sample sizes are nearly identical to those of PPI.
+The post-stratified confidence intervals and effective sample sizes are nearly identical to those of PPI.
 Figure 5: Simulation estimating maize fraction under different levels of map noise using $N=100000$ map product points and $n=100$ “ground truth” points.
-Olofsson and PPI results are overlapping.
+Post-stratified and PPI results are overlapping.
 Left: Maize fraction 95% confidence intervals at different noise levels.
@@
 Figure 6: Simulation estimating maize fraction 95% confidence intervals under different levels of map bias using $N=100000$ map product points and $n=100$ “ground truth” points.
-Olofsson and PPI confidence intervals are overlapping.
+Post-stratified and PPI confidence intervals are overlapping.
 (Noise is fixed at 0.4 so about 40% of the points are resampled.
@@
 As bias increases, the fraction of map product points labeled as “maize” increases.
-As a result, the map-based imputation estimates increase with the bias value, and generally do not remain centered at the true maize fraction of 0.35.
-The imputation confidence intervals are small because $N=100000$ is large, but they are unreliable because they do not correct for the bias in the map product.
+As a result, the map-only estimates increase with the bias value, and generally do not remain centered at the true maize fraction of 0.35.
+The map-only confidence intervals are small because $N=100000$ is large, but they are unreliable because they do not correct for the bias in the map product.
 In contrast, the PPI estimates remain centered at the true maize fraction because PPI corrects for bias in the map product.
@@
 Thus, PPI provides unbiased estimates with lower uncertainty than using only the ground truth points.
-Again, the Olofsson confidence intervals and effective sample sizes are nearly identical to those of PPI, as expected.
+Again, the post-stratified confidence intervals and effective sample sizes are nearly identical to those of PPI, as expected.
 We also plot the maize fraction confidence intervals against model accuracy for both the noise and bias experiments (Figure 7).
-As noise increases, model accuracy decreases, so low model accuracy is associated with wide PPI and Olofsson confidence intervals.
+As noise increases, model accuracy decreases, so low model accuracy is associated with wide PPI and post-stratified confidence intervals.
 Thus, the left panel of Figure 7 resembles an inverted version of the left panel of Figure 5.
 Figure 7: Simulation estimating maize fraction 95% confidence intervals at different map product accuracy levels obtained by varying noise (left) and bias (right), using $N=100000$ map product points and $n=100$ “ground truth” points.
-Olofsson and PPI confidence intervals are overlapping.
+Post-stratified and PPI confidence intervals are overlapping.
 As bias increases and the fraction of map product points labeled as “maize” increases, model accuracy decreases linearly.
-(To see this, note that if each resampled point is labeled as maize with probability $p$, then an expected $0.35(p)+0.65(1-p)=0.65-0.3p$ fraction of the resampled points will match the ground truth.) Thus, the map-based imputation estimates of maize fraction decrease linearly as model accuracy increases.
-In the right panel of Figure 7, a model accuracy of 0.86 corresponds to an imputed maize fraction of 0.21, significantly lower than the true fraction of 0.35.
-This shows that even a map with high model accuracy can yield biased estimates under naive imputation.
+(To see this, note that if each resampled point is labeled as maize with probability $p$, then an expected $0.35(p)+0.65(1-p)=0.65-0.3p$ fraction of the resampled points will match the ground truth.) Thus, the map-only estimates of maize fraction decrease linearly as model accuracy increases.
+In the right panel of Figure 7, a model accuracy of 0.86 corresponds to a map-only maize fraction estimate of 0.21, significantly lower than the true fraction of 0.35.
+This shows that even a map with high model accuracy can yield biased estimates under the map-only (pixel-counting) estimator.
 3.2 Deforestation area in Brazilian Amazon
@@
 Figure 9: Brazilian Amazon deforestation 95% confidence intervals using $N=4.4\times 10^{9}$ map product points and $n=1386$ ground truth points.
-The Olofsson confusion matrix method, PPI, and stratification all result in smaller confidence intervals for estimated deforestation than classical estimation, with Stratified PPI having the best performance.
+The stratified, post-stratified, and PPI estimators all result in smaller confidence intervals for estimated deforestation than classical estimation, with Stratified PPI having the best performance.
 Each method results in a point estimate that between 5% to 6% of the Brazilian Amazon was deforested between 2000 and 2015.
-The Olofsson confusion matrix method, PPI, and stratification all result in smaller 95% confidence intervals for estimated deforestation than classical estimation (Figure 9).
-The Olofsson and PPI confidence intervals are fairly similar, with respective effective sample sizes that are $1.9$ and $1.8$ times the actual number of ground truth samples.
+The stratified, post-stratified, and PPI estimators all result in smaller 95% confidence intervals for estimated deforestation than classical estimation (Figure 9).
+The post-stratified and PPI confidence intervals are fairly similar, with respective effective sample sizes that are $1.9$ and $1.8$ times the actual number of ground truth samples.
 The stratified and stratified PPI confidence intervals have the best performance, with effective sample sizes that outperform the classical method by a factor of about $2.4$.
 (Put differently, stratified PPI has effective sample size $3382$, so using the remotely sensed deforestation map added the equivalent of $3382-1386=1996$ ground truth samples.) Thus, the methods that use both map product and ground truth reduce uncertainty in our estimates, compared to using only ground truth.
-In particular, stratification is effective because the strata map is designed so that points within each stratum have low variance in deforestation status.
-In this case, the naive imputation estimate — using only the map to estimate deforestation without any bias correction via ground truth labels — falls within the other confidence intervals, suggesting that the NASA Global Forest Cover Change map has small bias for estimating deforestation area.
+In particular, the stratified estimator is effective because the strata map is designed so that points within each stratum have low variance in deforestation status.
+In this case, the map-only (pixel-counting) estimate — using only the map to estimate deforestation without any bias correction via ground truth labels — falls within the other confidence intervals, suggesting that the NASA Global Forest Cover Change map has small bias for estimating deforestation area.
 However, we emphasize that this would not have been clear without the ground truth data, and ground truth data is still necessary to check the reliability of the map product for deforestation analysis.
@@
 In particular, we estimate (1) logistic regression coefficients for Brazilian Amazon deforestation in 2000–2015, and (2) linear regression coefficients for US tree cover in 2021.
-We compare the classical method, naive imputation, PPI, stratification, and stratified PPI.
+We compare the classical method, map-only estimator, PPI, stratified estimator, and stratified PPI.
 Our methods only account for uncertainty in the remotely sensed target variable, while ignoring uncertainty in the covariates; we leave addressing the latter to future work.
@@
 The 95% confidence intervals for the four logistic regression coefficients are shown in Figure 11.
-The general degree of uncertainty in estimating the coefficients from our non-imputation methods are as follows.
+The general degree of uncertainty in estimating the coefficients from our non-map-only methods are as follows.
 The log odds of deforestation change by about .$-.016$ to .$-.010$ for each one-kilometer increase in distance from roads, .$-.008$ to .$-.001$ for each one-kilometer increase in distance from rivers, $-3$ to $0.8$ for each one-kilometer increase in elevation, and $-0.28$ to $-0.11$ for each one-degree increase in slope.
 Thus, deforestation is negatively associated with all covariates except for elevation.
-The elevation coefficient confidence intervals intersect with zero for all non-imputation methods except for stratified PPI, suggesting that there might be no association between elevation and deforestation after accounting for the other three variables.
+The elevation coefficient confidence intervals intersect with zero for all non-map-only methods except for stratified PPI, suggesting that there might be no association between elevation and deforestation after accounting for the other three variables.
 The PPI methods result in the smallest confidence intervals, with effective sample size outperforming the classical method by a factor of 1.1 to 2.5 depending on the covariate.
 The most significant improvement is for the slope covariate.
-Unlike in our previous area estimation examples, stratification (i.e., weighted logistic regression with weights determined by the strata) is not effective and even reduces effective sample size relative to the classical method when estimating the coefficients for elevation and distance from rivers.
+Unlike in our previous area estimation examples, the stratified estimator (i.e., weighted logistic regression with weights determined by the strata) is not effective and even reduces effective sample size relative to the classical method when estimating the coefficients for elevation and distance from rivers.
 Similarly, stratified PPI has slightly worse performance than PPI for all covariates except distance from roads.
 The strata were designed to minimize variance in deforestation status, which does not necessarily correspond to better estimates of regression coefficients for the covariates.
-For slope and distance from rivers, the imputed coefficient computed from the map product falls outside the confidence intervals for all methods that use ground truth points.
-The imputed coefficient for slope is about 0, and the imputed coefficient for distance from rivers is about $-0.010$.
+For slope and distance from rivers, the map-only estimated coefficient computed from the map product falls outside the confidence intervals for all methods that use ground truth points.
+The map-only coefficient for slope is about 0, and the map-only coefficient for distance from rivers is about $-0.010$.
 In particular, for the slope coefficient, the methods that use ground truth points find an association that is statistically significantly different from zero.
@@
 The PPI methods result in the smallest confidence intervals, while stratification is not generally effective.
-For slope and distance from rivers, the map-imputed coefficient falls outside the confidence intervals for all methods that use ground truth points.
+For slope and distance from rivers, the map-only coefficient falls outside the confidence intervals for all methods that use ground truth points.
 4.2 Tree cover linear regression
@@
 PPI outperforms the classical method.
-For elevation, the map-imputed coefficient falls outside the PPI confidence interval.
+For elevation, the map-only coefficient falls outside the PPI confidence interval.
 The 95% confidence intervals for the four linear regression coefficients are shown in Figure 13.
@@
 As in the other use cases, PPI uses the map product along with the ground truth to reduce uncertainty in the coefficient estimates compared to only using ground truth.
-For elevation, the imputed coefficient computed from the map product ($-5.3$) falls outside the PPI confidence interval.
+For elevation, the map-only coefficient computed from the map product ($-5.3$) falls outside the PPI confidence interval.
 This implies that there is bias in the tree cover map product’s association with elevation (e.g., the map product overestimates tree cover at low elevation locations or underestimates tree cover at high elevation locations).
@@
 In this work, we applied several methods that use remote sensing map products, along with a small amount of ground truth data, to estimate area and regression coefficients while correctly taking into account the errors in the map products.
-We find that stratified sampling, prediction-powered inference, and (for area estimation) the Olofsson method result in lower uncertainty than classical estimates using only ground truth data sampled.
-Unlike naive imputation using the map product alone, these estimates are guaranteed to be statistically valid.
-Bias is the main issue for naive imputation.
-Our experiments illustrated that a naive imputation of area or regression coefficients from a map without bias correction can yield incorrect inferences, and we differentiated between the influence of noise versus bias.
-In our simulation estimating maize fraction, naive imputation deviated from the true maize fraction of 0.35 as the bias of the crop type map increased.
-Furthermore, accuracy is not necessarily a good proxy for bias; even at seemingly high map accuracies ($\text{accuracy}>0.85$), naive imputation estimated maize fraction at 0.21 (a 40% underestimate).
-In our experiments on real data, imputation for deforestation in the Brazilian Amazon fell within stratified and PPI confidence intervals, suggesting an overall low-bias map product.
-However, imputation of the logistic regression coefficients on distance from rivers and slope fell outside the intervals computed through the classical, stratified, PPI, and stratified PPI methods.
-Naive imputation would have concluded that distance from rivers is significantly more predictive of deforestation than it actually is, and that slope has no relationship to deforestation when it actually does.
+We find that stratified, post-stratified, and prediction-powered inference estimates result in lower uncertainty than classical estimates using only ground truth data sampled.
+Unlike map-only estimates using the map product alone, these estimates are guaranteed to be unbiased.
+Bias is the main issue for map-only estimates.
+Our experiments illustrated that a map-only estimate of area or regression coefficients from a map without bias correction can yield incorrect inferences, and we differentiated between the influence of noise versus bias.
+In our simulation estimating maize fraction, map-only estimates deviated from the true maize fraction of 0.35 as the bias of the crop type map increased.
+Furthermore, accuracy is not necessarily a good proxy for bias; even at seemingly high map accuracies ($\text{accuracy}>0.85$), the map-only approach estimated maize fraction at 0.21 (a 40% underestimate).
+In our experiments on real data, the map-only estimate for deforestation in the Brazilian Amazon fell within stratified and PPI confidence intervals, suggesting an overall low-bias map product.
+However, the map-only estimates of the logistic regression coefficients on distance from rivers and slope fell outside the intervals computed through the classical, stratified, PPI, and stratified PPI methods.
+The map-only estimator would have concluded that distance from rivers is significantly more predictive of deforestation than it actually is, and that slope has no relationship to deforestation when it actually does.
 Thus, even when a map product is unbiased for the average outcome, it can be biased for the purpose of estimating regression coefficients.
-Similarly, imputation of the elevation coefficient in our linear regression for tree cover also overestimates the effect of elevation compared to PPI.
+Similarly, the map-only estimate of the elevation coefficient in our linear regression for tree cover also overestimates the effect of elevation compared to PPI.
 Effect of map accuracy.
-How much stratified sampling, PPI, and confusion matrix-based methods reduce uncertainty relative to classical estimates depends on the quality of the map product.
+How much stratified, post-stratified, and PPI estimators reduce uncertainty relative to the classical estimator depends on the quality of the map product.
 In stratification, the strata map should be designed such that the variance of the quantity of interest is low within each stratum.
-In PPI (and the Olofsson method), the variance of the difference between the map product predictions $\hat{Y}$ and the ground truth values $Y$ should be low.
-In our maize fraction experiments, a highly accurate map ($\text{accuracy}=0.96$) was equivalent to having nearly 6×6$\times$6 × as much ground truth data, whereas a low-accuracy map ($\text{accuracy}=0.73$) only increased the effective sample size by $1.2$.
-Stratification vs. PPI vs. Olofsson.
-Comparing stratified sampling, PPI, and the Olofsson method, which estimator yields the smallest confidence intervals depends on how the strata were designed and the quality of the map product.
+In PPI (and the post-stratified method), the variance of the difference between the map product predictions $\hat{Y}$ and the ground truth values $Y$ should be low.
+In our maize fraction experiments, a highly accurate map ($\text{accuracy}=0.96$) was equivalent to having nearly 6×6$\times$6 × as much ground truth data, whereas a low-accuracy map ($\text{accuracy}=0.73$) only increased the effective sample size by 1.2×1.2$\times$1.2 ×.
+Stratification vs. Post-stratification vs. PPI.
+Comparing stratified, post-stratified, and PPI estimators, which estimator yields the smallest confidence intervals depends on how the strata were designed and the quality of the map product.
 Recall Table 2, which summarizes the conditions under which each method is most appropriate.
@@
 This is a result of well-designed strata based on a remote sensing map of forest disturbance categories.
-PPI and the Olofsson method yielded comparable interval widths to each other.
-This is expected, as the Olofsson method yields similar confidence intervals as PPI when $N$ is large (Appendix A).
+PPI and the post-stratified estimator yielded comparable interval widths to each other.
+This is expected, as the post-stratified estimator yields similar confidence intervals as PPI when $N$ is large (Appendix A).
 The picture changes, however, when we reuse the same stratified samples to estimate regression coefficients in regressions that the strata were not designed for.
@@
 Indeed, the stratified estimator was actually worse than the classical estimator for certain coefficients.
-An advantage of PPI compared to stratification, then, is its flexibility.
-While stratification requires us to choose a map product before sampling, PPI can be applied to any map product and any randomly sampled ground truth dataset.
-Given the expensive nature of ground truth data collection, sampling without choosing a strata map and then using PPI to obtain valid confidence intervals may result in the best performance across a variety of downstream use cases.
+An advantage of PPI compared to the stratified estimator, then, is its flexibility.
+While the stratified estimator requires us to choose a map product before sampling, PPI can be applied to any map product and any randomly sampled ground truth dataset.
+Given the expensive nature of ground truth data collection, sampling without choosing a strata map and then using PPI to obtain confidence intervals may result in the best performance across a variety of downstream use cases.
 Or, if stratified samples are already available from area estimation, they can be combined with PPI to estimate regression coefficients.
@@
 Ultimately, the best methods combine map products with ground truth to yield the narrowest intervals (highest certainty).
-We recommend that producers of machine learning-generated remote sensing maps use a holdout ground truth dataset for uncertainty quantification.
-The criterion for this holdout set is that the model, including during hyperparameter tuning, was not trained on it; a standard test set would be adequate.
+We recommend that producers of machine learning-generated remote sensing maps use a (simple or stratified) randomly sampled holdout ground truth dataset for uncertainty quantification.
+The criterion for this holdout set is that the model, including during hyperparameter tuning, was not trained on it; a standard test set would be adequate, provided that it is randomly sampled.
 Map producers should release this ground truth data alongside their map product to make their map suitable for downstream scientific inferences.
@@
 Limitations.
-All of the estimation methods described in this paper (except for naive imputation) are limited to areas in which ground truth data is available.
+All of the estimation methods described in this paper (except for map-only estimators) are limited to areas in which ground truth data is available.
 However, this would be true for any assumption-free, data-driven method that quantifies uncertainty and corrects for map bias.
-Furthermore, we can only ensure statistical validity of our estimates when the ground truth points are randomly sampled.
+Furthermore, we can only ensure that our estimates are unbiased when the ground truth points are randomly sampled.
 Together, these attributes mean that such methods may be especially straightforward to use when ground truth data can be generated de novo, such as from human inspection of available satellite imagery, and more challenging to use when ground truth data requires bespoke surveys.
@@
 […]
-Appendix A Mathematical comparison of PPI and the method of Olofsson
-We show that when the number of map product points $N$ is large, the PPI and Olofsson area estimation confidence intervals are similar.
+Appendix A Mathematical comparison of PPI and post-stratified area estimator
+We show that when the number of map product points $N$ is large, the PPI and post-stratified area estimator confidence intervals are similar.
 In particular, we show (1) that they have approximately the same length when the number of samples is large and (2) they are centered on the same point.
-These two facts together provide a mathematical explanation for why PPI and the Olofsson method approximately agree in all of our experiments.
+These two facts together provide a mathematical explanation for why PPI and the post-stratified estimator approximately agree in all of our experiments.
 We begin with notation.
@@
 We have $A_{0}+A_{1}=1$.
-The map-imputed area estimate, which we will correct for bias, is $A_{1}$.
+The map-only area estimate, which we will correct for bias, is $A_{1}$.
 We will use the approximations $A_{0}\approx\frac{n_{0\cdot}}{n}$ and $A_{1}\approx\frac{n_{1\cdot}}{n}$.
@@
 $\lambda\rightarrow\frac{\mathrm{Cov}_{n}(Y,\hat{Y})}{\mathrm{Var}_{N}(\hat{Y})}\approx\frac{n_{00}n_{11}-n_{01}n_{10}}{n_{0\cdot}n_{1\cdot}}.$
-We compare this with the Olofsson standard error
-$SE_{\text{Olofsson}}$ $=\sqrt{A_{0}^{2}\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+A_{1}^{2}\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$
+We compare this with the post-stratified estimator standard error
+$SE_{\text{post}}$ $=\sqrt{A_{0}^{2}\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+A_{1}^{2}\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$
 $\approx\sqrt{\left(\frac{n_{0\cdot}}{n}\right)^{2}\times\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+\left(\frac{n_{1\cdot}}{n}\right)^{2}\times\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$ plugging in $A_{0}$ and $A_{1}$
 $=\sqrt{\frac{1}{n^{2}n_{0\cdot}n_{1\cdot}}(n_{01}n_{00}n_{10}+n_{01}n_{00}n_{11}+n_{11}n_{10}n_{00}+n_{11}n_{10}n_{01})}$ using algebraic manipulation.
-The expressions we arrive at for PPI and the method of Olofsson agree, indicating that for large sample sizes the standard errors are approximately equal.
+The expressions we arrive at for PPI and the post-stratified estimator agree, indicating that for large sample sizes the standard errors are approximately equal.
 This implies that the confidence intervals from the two methods will have the same width.
 Point estimates.
-Next, we show that the PPI and Olofsson confidence intervals are centered around the same point. Turning first to the PPI point estimate
+Next, we show that the PPI and post-stratified confidence intervals are centered around the same point. Turning first to the PPI point estimate
 $\hat{\theta}_{\text{PPI}}$ $=\lambda A_{1}-\frac{1}{n}\sum_{i=1}^{n}(\lambda\hat{Y}_{i}-Y_{i})$
@@
 $=\frac{n_{\cdot 1}}{n}$
-The Olofsson area point estimate is
-$\hat{\theta}_{\text{Olofsson}}$ $=A_{0}\frac{n_{01}}{n_{0\cdot}}+A_{1}\frac{n_{11}}{n_{1\cdot}}$
+The post-stratified area point estimate is
+$\hat{\theta}_{\text{post}}$ $=A_{0}\frac{n_{01}}{n_{0\cdot}}+A_{1}\frac{n_{11}}{n_{1\cdot}}$
 $\approx\frac{n_{0\cdot}}{n}\times\frac{n_{01}}{n_{0\cdot}}+\frac{n_{1\cdot}}{n}\times\frac{n_{11}}{n_{1\cdot}}$ plugging in $A_{0}$ and $A_{1}$

### v4

@@
-references.bib
+Regression coefficient estimation from remote sensing maps
 Abstract
-Remote sensing map products are used to obtain estimates of environmental quantities, such as deforested area or the effect of conservation zones on deforestation.
+Remote sensing map products are used to estimate regression coefficients relating environmental variables, such as the effect of conservation zones on deforestation.
 However, the quality of map products varies, and — because maps are outputs of complex machine learning algorithms that take in a variety of remotely sensed variables as inputs — errors are difficult to characterize.
 Thus, population-level estimates from such maps may be biased.
-In this paper, we compare several uncertainty quantification methods — stratified estimator, post-stratified estimator, and prediction-powered inference — that combine a small amount of randomly sampled ground truth data with large-scale remote sensing map products to generate unbiased estimates.
-Applying these methods across four remote sensing use cases in area and regression coefficient estimation, we find that they result in estimates that are more reliable than using the map product as if it were 100% accurate and have lower uncertainty than using only the ground truth and ignoring the map product.
-Prediction-powered inference uses ground truth data to correct for bias in the map product estimate and (unlike stratification) does not require us to choose a map product before sampling.
-This is the first work to (1) apply prediction-powered inference to remote sensing estimation tasks, and (2) perform uncertainty quantification on remote sensing regression coefficients without assumptions on the structure of map product errors.
-To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a randomly sampled holdout ground truth dataset to be used for calibration in uncertainty quantification alongside their maps.
+In this paper, we apply prediction-powered inference (PPI) to regression coefficient estimation.
+PPI generates unbiased estimates by using a small amount of randomly sampled ground truth data to correct for bias in large-scale remote sensing map products.
+Applying PPI across multiple remote sensing use cases in regression coefficient estimation, we find that it results in estimates that are (1) more reliable than using the map product as if it were 100% accurate and (2) have lower uncertainty than using only the ground truth and ignoring the map product.
+Empirically, we observe effective sample size increases of up to 17-fold using PPI compared to only using ground truth data.
+This is the first work to estimate remote sensing regression coefficients without assumptions on the structure of map product errors.
 Data and code are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
 1 Introduction
-Remote sensing maps are widely used to obtain estimates for environmental quantities [hansen2013high, baccini2012estimated, potapov2022global, wang2019great, deines2019satellites].
+Remote sensing maps are widely used to obtain estimates for environmental quantities, such as cropland expansion over time or the effect of protected areas on deforestation [hansen2013high, baccini2012estimated, potapov2022global, wang2019great, deines2019satellites].
 Map products created by training machine learning models on remotely sensed data allow researchers to cheaply access predictions for a variable of interest at a high resolution over large regions.
@@
 Another recent study of three satellite-based global land cover/land use map products found that the maps are biased toward different land cover classes (e.g., one map overestimates shrub cover and another overestimates grass cover relative to the other maps) [venter2022global].
-Additionally, errors in map products are often difficult to characterize due to the complexity of input data and lack of interpretability in the ML models used to generate the maps.
+Errors in map products are often difficult to characterize due to the complexity of input data and lack of interpretability in the ML models used to generate the maps.
 Examples of systematic errors in map products trained on satellite data include “nonrandom misclassification, saturation effects, atmospheric effects, and cloud cover,” all of which may lead to biased inference [jain2020benefits].
@@
 However, ground truth data is often sparse because it is expensive to collect, as it may require fieldwork or manual labelling of remote sensing imagery.
-Few ground truth points result in greater uncertainty (e.g., wider confidence intervals) in estimates.
+Few ground truth points result in greater uncertainty (i.e., wider confidence intervals) in estimates.
 Another alternative is to use both ground truth data and remote sensing maps to generate estimates, using the ground truth points to correct for biases that may exist in the map products.
 This generally reduces uncertainty in the estimates compared to using ground truth alone, since map products contain a large amount of data that, while possibly biased, still give useful information about the true values we wish to measure.
-Figure 1: Motivating example: map-only estimate is biased.
-The 95% confidence interval for logistic regression coefficient relating Brazilian Amazon deforestation to slope using $N=963548$ remote sensing map points and $n=1386$ ground truth points.
-The map-only slope coefficient estimate falls outside the confidence intervals for all methods that use ground truth points.
+As an example of this approach, the remote sensing community commonly uses a post-stratified estimator to generate confidence intervals for area estimation.
+A pixel-counting estimator, which uses only a map product to estimate area of a land cover class, may be biased due to misclassification error in the map product [olofsson2020mitigating].
+An unbiased area estimate can be obtained by using randomly sampled ground truth data [olofsson2014good, stehman2019key], but if ground truth observations are sparse, this may result in less precise estimates (wider confidence intervals) than desired.
+To resolve this, the post-stratified area estimator [card1982using, stehman2013estimating, olofsson2013making] combines a map product with a small amount of ground truth data, using the confusion matrix between land cover ground truth and map product points to correct for bias in the map product.
+By leveraging the map product, the post-stratified estimator also produces narrower confidence intervals compared to only using ground truth data.
+Other area estimators that combine ground truth and map product data, including regression, ratio, and calibration estimators, are described in [gallego2004remote].
+Figure 1: Overview of estimation methods using ground truth data and remote sensing map products.
+The post-stratified estimator (for area estimation) and prediction-powered inference (for regression coefficient estimation) use a small amount of ground truth data along with a map product to produce confidence intervals for the quantity of interest.
+In this paper, we focus on regression coefficient estimation using PPI.
+We also compare PPI against GT-only (using only ground truth) and map-only estimators.
+The remote sensing community also uses map products for regression coefficient estimation: estimating linear or logistic regression coefficients for causal or associative relationships between two or more environmental variables.
+Unlike area estimation with the post-stratified estimator, regression coefficient estimation from remote sensing maps still suffers from bias due to map product errors.
+An unbiased estimate can be obtained by running a regression on randomly sampled ground truth data [sarndal2003model], but similarly to the area estimation case, this may result in less precise estimates than desired.
+In this paper, we address this problem by applying prediction-powered inference (PPI) to regression coefficient estimation (Figure 1).
+PPI [angelopoulos2023prediction, angelopoulos2023ppi++, Kluger25GeneralizingPPI] is a recently developed unbiased estimation method that uses a small set of randomly sampled ground truth observations ($n$ points) to calibrate and correct for bias in a much larger set of machine learning-generated predictions ($N$ points, where $N$ is much larger than $n$).
+This method also results in more precise estimates (higher statistical power, smaller confidence intervals) than using the ground truth points alone.
+We apply PPI to regression use cases where the machine learning predictions are obtained from large-scale remote sensing map products.
+We compare results with the ground truth-only (GT-only) estimator (using only ground truth points) and map-only estimator (using only a map product).
+These estimators are described in detail in Section 2.
+Figure 2: Motivating example: PPI produces unbiased coefficient estimate with lower variance than using ground truth alone.
+We estimate the linear regression coefficient relating US forest cover to elevation (with $N=67968$ remote sensing map points and $n=500$ ground truth points).
+Left: Map-only, ground truth-only, and PPI point estimates and 95% confidence intervals.
+Right: Histogram of PPI and GT-only point estimates from repeating the experiment with 100 random sets of $n=500$ ground truth points.
 A motivating example.
-In this work, we discuss methods that correct map products with ground-truth data to give unbiased confidence intervals.
-We give a brief preview here.
-We seek to estimate logistic regression coefficients for the association between Brazilian Amazon deforestation in 2000–2015 and several covariates, including terrain slope.
-Using nearly one million remote sensing map product points and $1386$ ground truth points, we compare the $95\%$ confidence intervals for the slope coefficient with four methods (Figure 1) described shortly.
+We give a brief preview of how we use PPI to correct map products with ground-truth data to obtain an unbiased regression coefficient estimate.
+We seek to estimate linear regression coefficients for the association between US forest cover and a few covariates, including elevation.
+We use $N=67968$ remote sensing map product points and $n=500$ ground truth points.
 See Section 3.1 for a complete description of the experimental setup.
-The “classical” estimate using only ground truth points is unbiased but has the widest confidence interval, failing to take advantage of existing map products.
-Meanwhile, methods that use a map product along with the ground truth points produce an estimate with lower uncertainty while still being unbiased.
-Notably, the slope coefficient computed from the map product alone is approximately zero, which falls outside the confidence intervals for all methods that use ground truth points.
-That is, if we had used only the map product, we would not have discovered the statistically significant negative relationship between slope and deforestation.
-Methods like PPI allow us to use ground truth points to correct for bias in the map product and thus avoid misleading conclusions.
-Figure 2: Overview of uncertainty quantification using ground truth data and remote sensing map products.
-Stratified, post-stratified, and prediction powered inference estimators use a small amount of ground truth data along with a map product to produce confidence intervals for the quantity of interest.
-We also compare these methods against classical uncertainty quantification (using only ground truth) and map-only estimators.
+We compare the $95\%$ confidence intervals for the elevation coefficient estimates from PPI, GT-only, and map-only methods (Figure 2, left panel).
+The “GT-only” estimate using only ground truth points is unbiased but has the widest confidence interval, failing to take advantage of existing map products.
+Meanwhile, PPI uses the map product along with the ground truth points to produce a much smaller confidence interval while still being unbiased.
+The PPI and GT-only confidence intervals both contain the true coefficient, while the confidence interval estimated from the map product alone does not.
+That is, if we had used only the map product, we would have underestimated the magnitude of the negative relationship between forest cover and elevation.
+PPI avoids misleading conclusions by using ground truth points to correct for bias, and at the same time, it gives smaller confidence intervals than using only the ground truth points.
+We repeat the experiment using the same set of $N=67968$ map product points and 100 different randomly chosen sets of $n=500$ ground truth points (see Appendix C.4 for more detail).
+A histogram of the 100 PPI and GT-only elevation coefficient point estimates is shown in the right panel of Figure 2.
+As before, we see that the PPI and GT-only estimators are both unbiased relative to the true coefficient, but the PPI estimates are more precise as they have a lower variance than the GT-only estimates.
 Our contribution.
-In this paper, we consider uncertainty quantification methods that combine a small number of randomly sampled ground truth data ($n$ points) with large-scale remote sensing map products ($N$ points, where $N$ is much larger than $n$) in order to generate unbiased estimates with low uncertainty (Figure 2).
-We also compare these methods against the classical estimator (using only ground truth) and map-only estimator (using only a map product).
-We consider the following three methods:
-  • Stratified estimator uses map products to determine a sampling strategy for ground truth points that reduces the variance of estimates.
-This method has been used in the remote sensing community for decades [hay1979sampling, fitzpatrick1981comparison, stehman1998design, olofsson2014good].
-  • Post-stratified area estimator uses the confusion matrix between ground truth and map product land cover labels to correct for misclassification bias in the map product, when estimating the area of a land cover class.
-This method is also commonly used in the remote sensing community [card1982using, stehman2013estimating, olofsson2013making].
-  • Prediction-powered inference (PPI) [angelopoulos2023prediction, angelopoulos2023ppi++] is a recently developed uncertainty quantification method that uses a small set of ground truth points to calibrate and correct for bias in a much larger set of ML-generated predictions, resulting in more precise estimates than using the ground truth points alone.
-We apply these methods to two of the most common scientific uses of remote sensing maps: area estimation and regression coefficient estimation.
-Area estimation includes estimating both total area occupied by a land cover/land use category and total change in land cover/land use over time.
-The methods applied to area estimation also apply to estimating sums or averages of continuous variables, such as total carbon stock.
-A map-only estimator for area is known as a pixel-counting estimator.
-Regression coefficient estimation refers to estimating linear or logistic regression coefficients for causal or associative relationships between two or more environmental variables.
-For regressions, this paper only accounts for uncertainty in the remotely sensed dependent variable, while the covariates are assumed to have no errors.
-To illustrate the methods, we consider four specific remote sensing estimation tasks spanning crop area estimation, deforestation quantification, and associating deforestation and tree cover to other geophysical variables; see Table 1.
+In this paper, we illustrate the use of PPI on four specific regression coefficient estimation tasks associating deforestation, tree cover, housing price, and forest cover to other geospatial variables; see Table 1.
+We then simulate map error to understand its influence on the bias and variance of map-only, GT-only, and PPI estimators.
+We demonstrate that PPI can be applied when there is error in the remotely sensed response variable, the remotely sensed covariates, or both.
 Data and code for these use cases are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
-Estimation task Target variable Region Year Map product(s) Ground truth Covariates Link to code
-Area Maize Iowa 2022 USDA NASS Cropland Data Layers (CDL) [nass2024usda] Simulated N/A link
-Area Deforestation Brazilian Amazon 2000-2015 1. NASA Global Forest Cover Change [townshend2016global]
-2. Amazon ecoregion
-deforestation strata [bullock2020satellite]
-$n=1386$
+Regression type Response variable Covariates Region Year Map product(s) Ground truth Link to code
+Logistic Deforestation Distance from roads, distance from rivers, elevation, slope Brazilian Amazon 2000-2015 NASA Global Forest Cover Change [townshend2016global] $n=1386$
 from [bullock2020satellite]
-N/A link
-Logistic
-regression
-coefficients
-Distance from roads, distance from rivers, elevation, slope link
-Linear
-regression
-coefficients
-Tree cover Contiguous USA 2021 USFS Tree Canopy Cover [forest2023usfs] $n=983$
-manually labelled Google Maps high-resolution satellite images
-Aridity index,
+link
+Linear Tree cover Aridity index,
 elevation, slope
+Contiguous USA 2021 USFS Tree Canopy Cover [forest2023usfs] $n=983$
+manually labeled images (Sec. 3.1)
+link
+Linear Housing price Income, nightlights, road length Contiguous USA 2010-2021 MOSAIKS [rolf2021generalizable] $n=500$
+from [rolf2021generalizable]
+link
+Linear Forest cover Elevation,
+population
+Contiguous USA 2010 MOSAIKS [rolf2021generalizable] $n=500$
+from [rolf2021generalizable]
 link
 Table 1: Overview of use cases and datasets.
-We estimate maize area in Iowa, deforestation area and logistic regression coefficients in the Brazilian Amazon, and tree cover linear regression coefficients in the United States.
-The Iowa maize area use case is a simulation: we treat a small random sample of CDL map product points as “ground truth” data and observe the effects of artificially degrading the map product relative to the ground truth.
+To illustrate the mechanics and advantages of PPI, we estimate linear regression coefficients for tree cover, housing price, and forest cover in the United States.
 Code implementations of the use cases are linked in the rightmost column.
-Combining ground truth data with map products results in unbiased estimates that are more reliable than map-only estimates and have lower uncertainty than classical estimates.
-As we will show, map-only estimators without bias correction can result in completely incorrect area and regression coefficient estimates.
-To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a (simple or stratified) randomly sampled holdout ground truth dataset to be used for calibration in uncertainty quantification alongside their maps, or, if that is not done, for map users to generate their own randomly sampled holdout set for calibration, if possible (e.g., from visual inspection of remote sensing data).
+Using PPI to combine ground truth data with map products results in unbiased estimates that are more reliable than map-only estimates and have lower uncertainty than GT-only estimates.
+As we will show, map-only estimators without bias correction can result in incorrect regression coefficient estimates.
+To improve the utility of machine learning-generated remote sensing maps for downstream regression applications, we recommend that map producers provide a randomly sampled holdout ground truth dataset to be used for calibration in PPI alongside their maps, or, if that is not done, for map users to generate their own randomly sampled holdout set for calibration, if possible (e.g., from visual inspection of remote sensing data).
+This holdout set can be obtained by simple (uniform) random sampling or other sampling designs such as stratified, weighted, or clustered random sampling [Kluger25GeneralizingPPI].
 Related work.
-The remote sensing community commonly uses a post-stratified estimator to generate confidence intervals for area estimates.
-This estimator uses the confusion matrix between land cover ground truth and map product points to correct for misclassification bias in the map product [card1982using, stehman2013estimating, olofsson2013making].
-This method only applies to area estimation, not regression coefficient estimation tasks.
-One contribution of the present work is a comparison of the post-stratified area estimator and PPI.
-We find both experimentally and mathematically that these two methods produce similar outputs for area estimation.
-For other methods, [gallego2004remote] provides a survey of area estimators that combine ground truth and map product data, including regression, ratio, and calibration estimators.
-Methods have also been developed to reduce bias in regression coefficient estimates from remote sensing data.
+Previous works have developed methods to reduce bias in regression coefficient estimates from remote sensing data.
 For example, [alix2023remotely] estimates the effect of a conservation program on deforestation in Mexico.
-They correct for bias in the remotely sensed deforestation binary variable by modelling misclassification error as dependent on environmental covariates via a logit model.
+They correct for bias in the remotely sensed deforestation binary variable by modelling misclassification probabilities as a parametric function of environmental covariates.
+However, this approach imposes assumptions about the structure of the misclassification errors and moreover is only developed for settings with binary response variables.
 Another study [gordon2023remote] estimates logistic regression coefficients for the relationship between roads and forest cover in West Africa.
-They train an adversarial debiasing model with ground truth data to correct for bias arising from correlations between distance from roads and the measurement error in remotely sensed forest cover.
+They train an adversarial debiasing model with ground truth data to correct for bias arising from correlations between distance from roads and the measurement error in remotely sensed forest cover; however, they do not provide an approach for constructing confidence intervals.
 Another study [proctor2023parameter] uses co-located ground truth and remotely sensed data to show that using remote sensing data alone leads to biased regression coefficient estimates for several effect size estimation tasks in the US.
-They use a linear measurement error model and apply multiple imputation: a statistical method that uses a small amount of ground truth data to estimate remote sensing measurement error and impute corrected values for remotely-sensed variables.
-Note that this method requires an assumption about the error structure of the remotely-sensed variables, whereas the methods investigated herein do not.
-Another approach to uncertainty quantification is conformal prediction, which generates confidence sets for machine learning predictions at individual data points.
+To remedy this issue, they propose using a Bayesian linear measurement error model and applying multiple imputation: a statistical method that involves estimating the missing values in a dataset and their uncertainties and fitting the regression on multiple randomly imputed versions of the dataset.
+Note that this method requires an assumption about the error structure of the remotely-sensed variables and does not have any frequentist coverage guarantees.
+For example, in a subset of the simulation-based experiments in [proctor2023parameter] 95% confidence intervals from multiple imputation only contained the true value of the regression coefficient in fewer than 80% of the simulations as opposed to the desired $\sim$95of the simulations.
+In contrast to existing approaches in the remote sensing literature, PPI does not make any modelling assumptions about the measurement error structure, and under certain regularity conditions, its confidence intervals are guaranteed to attain the desired coverage [angelopoulos2023prediction, angelopoulos2023ppi++, Kluger25GeneralizingPPI].
+Another approach to uncertainty quantification is conformal prediction [vovk2005algorithmic, angelopoulos2023conformal], which generates confidence sets for machine learning predictions at individual data points.
 For example, [valle2023quantifying] applies conformal prediction to quantify pixel-level uncertainty for a land cover/land use map product in the Brazilian Amazon, outputting a 90% predictive set of land cover classes for each pixel (i.e., the set contains the true land cover class 90% of the time).
-This approach differs from our work because our methods estimate total area and regression coefficients over an entire region, rather than quantifying uncertainty at individual points or pixels.
-Unlike previous works, this paper is the first work to apply prediction-powered inference to remote sensing use cases in area and regression coefficient estimation.
-Unlike stratified random sampling, PPI does not require that ground truth points be sampled according to a pre-determined stratification map; as a result, it allows for greater flexibility in ground truth data usage across different downstream applications.
-Additionally, this is the first work to combine map product and ground truth data to perform uncertainty quantification on remote sensing regression coefficients without assuming an error structure on the map product predictions.
-2 Background: Uncertainty Quantification
-We formally describe uncertainty quantification methods that generate confidence intervals with remote sensing variables.
-As two baselines, we first consider methods that use only the ground-truth data (which is unbiased but loses information) or only use the remote-sensing data (which may be biased).
-We then turn to methods that use both sources of data to produce unbiased confidence intervals.
-See Table 2 for a summary.
-Method Data Used Unbiased? Caveats
-Map Ground Truth
-Map-only
-Classical
-Stratified Ground truth points sampled
-using stratified map
-Post-stratified Area estimation only
-PPI
-Table 2: Summary of uncertainty quantification methods. PPI is the most flexible method that incorporates both ground truth and map product data.
-2.1 Baseline methods
-Classical: use only the ground-truth data.
-First, classical uncertainty quantification uses ground truth observations, sampled uniformly at random, to estimate the quantity of interest.
-For example, suppose there is a target variable $Y$ that can be observed over a given area, and we wish to generate a 95% confidence interval for its population mean $\theta^{\star}=E[Y]$.
+This approach differs from our work because PPI estimates regression coefficients over an entire region, rather than quantifying uncertainty of map estimates at individual points or pixels.
+Unlike previous works, this paper is the first work to apply prediction-powered inference to remote sensing use cases in regression coefficient estimation.
+This is also the first work to combine map product and ground truth data to estimate remote sensing regression coefficients without assuming an error structure on the map product predictions.
+2 Methods: Prediction-Powered Inference for Regression
+Regressions are commonly used in the environmental sciences and economics to identify relationships — associative or causal — between covariates and a binary response (logistic regression) or between covariates and a continuous response (linear regression).
+Logistic regression coefficients give the change in the log-odds ratio of a binary response corresponding to a one-unit increase in a covariate, when all other covariates are held constant.
+Linear regression coefficients give the change in the value of a continuous response corresponding to a one-unit increase in a covariate, when all other covariates are held constant.
+Thus, we can interpret these coefficients as effect sizes of the covariates.
+Historically, scientists would run regressions using ground truth environmental samples, but regressions on data from remote sensing maps are becoming increasingly common, e.g., [taylor2023fertilizer, deines2023recent, he2022marked, taylor2022wetlands, deines2019satellites, lobell2022globally].
+When covariate or response variable values are machine learning predictions from a map product, errors in the machine learning model may result in biased regression coefficients (as we have seen in Figure 2).
+In this section, we formally describe how to use prediction-powered inference (PPI) to generate regression coefficient estimates and confidence intervals when the response variable $Y$ and/or covariates $X$ are remotely sensed map products.
+As two baselines, we consider the ground truth-only estimator that only uses ground-truth data (which is unbiased but does not leverage remote sensing maps) and the map-only estimator that only uses remote sensing maps (which may be biased).
+We then describe PPI, which uses both sources of data to produce unbiased confidence intervals.
+Figure 1 summarizes the methods.
+As a technical note, this paper focuses on inference in settings where both the ground truth data and the larger sample from a remote sensing map product are assumed to be randomly sampled from a superpopulation, as opposed to a design-based inference framework for reasons discussed in Appendix A.
+Suppose we have a remote sensing data product with a large number $N$ of predictions $(\hat{X}^{\prime}_{1},\hat{Y}^{\prime}_{1})$, $(\hat{X}^{\prime}_{2},\hat{Y}^{\prime}_{2})$, $\dots$, $(\hat{X}^{\prime}_{N},\hat{Y}^{\prime}_{N})$ where $X\in\mathbb{R}^{p}$ is a vector of covariates and $Y\in\mathbb{R}$ is the response variable.
+Suppose that for a small calibration set of $n\ll N$ points sampled uniformly at random, we have ground truth labels $(X_{1},Y_{1})$, $(X_{2},Y_{2})$$\dots$, $(X_{n},Y_{n})$ with corresponding remote sensing predictions $(\hat{X}_{1},\hat{Y}_{1}),(\hat{X}_{2},\hat{Y}_{2}),\dots,(\hat{X}_{n},\hat{Y}_{n})$.
+(Note that in some use cases, ground truth data may be widely available either for the response variable, for a subset of the covariates, or for a subset of both.
+In these cases, we would use ground truth instead of remotely sensed values for the corresponding components of the $(\hat{X},\hat{Y})$ and $(\hat{X}^{\prime},\hat{Y}^{\prime})$ vectors.)
+2.1 Linear regression
+In the case of linear regression, we wish to estimate 95% confidence intervals for the regression coefficients $\beta_{*}$ giving the best fit to the linear model
+$Y=\beta^{\mathsf{T}}X+\varepsilon,$
+where $\varepsilon$ is mean 00 and uncorrelated with $X$.
+(We assume that the first element of the vector $X$ is a constant 1, so that the first regression coefficient $\beta_{0}$ is an intercept term.)
+Ground truth-only estimator.
+The ground truth-only (GT-only) estimator uses only the $n$ ground truth calibration datapoints to compute the regression coefficient estimate
+$\hat{\beta}_{\text{calib}}=\operatorname*{arg\,min}_{\beta}\sum_{i=1}^{n}(Y_{i}-\beta^{\mathsf{T}}X_{i})^{2}.$
+However, the standard error of this ground truth coefficient estimate will be large when the number of ground truth observations $n$ is small, which is often the case in environmental applications.
+As a result, the GT-only 95% confidence interval will be large.
+Map-only estimator: Use only the remote sensing map.
+The map-only estimator uses only the $N$ map data points to compute the regression coefficient estimate
+$\hat{\gamma}_{\text{map}}=\operatorname*{arg\,min}_{\beta}\sum_{i=1}^{N}(\hat{Y}^{\prime}_{i}-\beta^{\mathsf{T}}\hat{X}^{\prime}_{i})^{2}.$
+The standard error of the map-only coefficient estimate will be small if $N$ is large, resulting in a small map-only 95% confidence interval.
+However, we employ the $\gamma$ notation to emphasize that the map-only estimator may be biased; if the remotely sensed values $\hat{X}^{\prime}_{i}$ or $\hat{Y}^{\prime}_{i}$ are inaccurate, $\hat{\gamma}_{\text{map}}$ might be biased in the sense that as the sample size goes to infinity, $\hat{\gamma}_{\text{map}}$ will converge (in probability) to some $\gamma_{*}\neq\beta_{*}$.
+Thus, the map-only confidence interval is not guaranteed to have the appropriate coverage rate (in this case, 95%) of the true parameter of interest.
+This is because the map-only estimator does not account for error or uncertainty in the model used to generate the map product predictions.
+Prediction-Powered Inference: Map + ground truth.
+Introduced in 2023, Prediction-Powered Inference (PPI) [angelopoulos2023prediction] is an uncertainty quantification approach that computes confidence intervals using a small calibration set of ground truth points and a much larger set of machine learning-generated predictions.
+The main idea is that PPI assesses the bias in the map product using the ground truth data and then corrects it.
+For regression coefficient estimation, we use variants of PPI that were developed in [Kluger25GeneralizingPPI] and have origins in [chen2000unified, zrnic2024note].
+(PPI can also be applied to mean and area estimation; see Appendices D and E.)
+PPI uses the calibration data to correct for bias in the map-based estimate.
+First, it computes the regression coefficient estimate using the remote sensing predictions at the $n$ calibration datapoints:
+$\hat{\gamma}_{\text{calib}}=\operatorname*{arg\,min}_{\beta}\sum_{i=1}^{n}(\hat{Y}_{i}-\beta^{\mathsf{T}}\hat{X}_{i})^{2}.$
+Then, the PPI regression coefficient estimate corrects the map-only estimate $\hat{\gamma}_{\text{map}}$ for bias, using the difference between the ground truth and map estimates on the calibration set $(\hat{\beta}_{\text{calib}}-\hat{\gamma}_{\text{calib}})$.
+The final PPI estimate is
+$\hat{\beta}_{\text{PPI}}=\underbrace{\hat{\gamma}_{\text{map}}}_{\text{map-only estimate}}+\underbrace{(\hat{\beta}_{\text{calib}}-\hat{\gamma}_{\text{calib}})}_{\text{bias correction}}.$
+The PPI estimate can be improved by computing a tuning matrix $\hat{\Omega}$ that minimizes the variance of each component of the estimated coefficient vector.
+The value of $\hat{\Omega}$ depends on the quality of the map product relative to the ground truth.
+The tuned PPI regression coefficient estimate is
+$\hat{\beta}_{\text{PPI}}=\underbrace{\hat{\Omega}\hat{\gamma}_{\text{map}}}_{\text{map-only estimate}}+\underbrace{(\hat{\beta}_{\text{calib}}-\hat{\Omega}\hat{\gamma}_{\text{calib}})}_{\text{bias correction}}.$ (1)
+The tuned $\hat{\beta}_{\text{PPI}}$ is still an unbiased estimate for the true coefficient $\beta_{*}$.
+The analytic formula for an optimal choice of tuning matrix is presented in Appendix B for interested readers.
+We compute a PPI confidence interval for the regression coefficients using the prediction-powered percentile bootstrap method [zrnic2024note, Kluger25GeneralizingPPI].
+In particular, we resample both the calibration set and map data points with replacement $B$ times, and for each $1\leq i\leq B$, we compute the point estimate $\hat{\beta}_{\text{PPI}}^{(i)}$ using the $i$th resampled dataset.
+For each $j=1,\dots,p$, the $(1-\alpha)$-level confidence interval for $\beta_{j}$ has endpoints that are given by the $\alpha/2$ percentile and $21-\alpha/$ 2 percentile of $j$th entries of the $B$ estimated vectors $\hat{\beta}_{\text{PPI}}^{(1)},\hat{\beta}_{\text{PPI}}^{(2)},\dots,\hat{\beta}_{\text{PPI}}^{(B)}$.
+Larger values of $B$ result in more algorithmically stable confidence intervals, at the cost of increased computational time.
+Note that an alternative approach to constructing confidence intervals for the estimator in Equation (1) would be to derive an asymptotic normal approximation and variance formula for $\hat{\beta}_{\text{PPI}}$ (e.g., such an approach was developed in [chen2000unified]; we derive it in Appendix B).
+However, in this paper we focus on bootstrap-based approaches to constructing confidence intervals because they generalize more readily to nonstandard regression models and can be modified to apply to weighted, stratified, and clustered samples [Kluger25GeneralizingPPI].
+2.2 Logistic regression
+When $Y$ is a binary variable, we can similarly use PPI to estimate the logistic regression coefficients $\beta$ that give the best fit to the model
+$\mathbb{P}(Y=1|X)=\frac{1}{1+e^{-\beta^{\mathsf{T}}X}}.$
+(We assume that the first element of the vector $X$ is a constant 1, so that the first regression coefficient $\beta_{0}$ is an intercept term.)
+The ground truth-only estimator using the $n$ ground truth calibration datapoints is
+$\hat{\beta}_{\text{calib}}=\operatorname*{arg\,min}_{\beta}\Bigl{\{}\sum_{i=1}^{n}\log(1+e^{\beta^{\mathsf{T}}X_{i}})-Y_{i}\beta^{\mathsf{T}}X_{i}\Bigr{\}}.$
+The map-only estimator using the $N$ map datapoints is
+$\hat{\gamma}_{\text{map}}=\operatorname*{arg\,min}_{\beta}\Bigl{\{}\sum_{i=1}^{N}\log(1+e^{\beta^{\mathsf{T}}\hat{X}^{\prime}_{i}})-\hat{Y}^{\prime}_{i}\beta^{\mathsf{T}}\hat{X}^{\prime}_{i}\Bigr{\}}.$
+The estimator using the remote sensing predictions at the $n$ calibration datapoints is
+$\hat{\gamma}_{\text{calib}}=\operatorname*{arg\,min}_{\beta}\Bigl{\{}\sum_{i=1}^{n}\log(1+e^{\beta^{\mathsf{T}}\hat{X}_{i}})-\hat{Y}_{i}\beta^{\mathsf{T}}\hat{X}_{i}\Bigr{\}}.$
+Similar to the linear regression case, the PPI logistic regression coefficient estimate is obtained by plugging these values into Equation (1), and the confidence interval is obtained using the percentile bootstrap.
+In the case of logistic regression, analytic formulas for an optimal choice of tuning matrix and an estimate of the covariance matrix of $\hat{\beta}_{\text{PPI}}$ are also presented in Appendix B.
+2.3 Error-in-$Y$ vs. Error-in-$X$ vs. Error-in-both
+PPI for regression coefficient estimation can be applied even when only a subset of the variables are remotely sensed, while the others have ground truth labels at all $N$ locations.
+We study PPI in three different regression regimes of interest, using the same nomenclature and categorization as [proctor2023parameter].
+In the first regime, which we call error-in-$Y$, the investigator wishes to leverage a map product for the response variable, while ground truth data for the covariates are widely available.
+In the second regime, which we call error-in-$X$, the investigator wishes to leverage a map product for some of the covariates, while ground truth data for the response variable is widely available.
+In the third regime, which we call error-in-both, ground truth labels are sparsely available for both the response variable and some of the covariates, and the investigator wishes to leverage a map product for these sparsely measured variables.
+We consider these three regimes separately for two reasons.
+First, many existing statistical methods either only apply to the error-in-$X$ regime or only apply to the error-in-$Y$ regime.
+For example, some earlier variants of PPI [angelopoulos2023prediction, angelopoulos2023ppi++] are only designed for the error-in-$Y$ regime, while many of the traditional methods in the measurement error literature [CarrolStefanski06] are only designed for error-in-$X$ regime.
+However, we emphasize that the variant of PPI presented in this paper can apply to all three regimes [Kluger25GeneralizingPPI].
+Second, the bias of the coefficient is thought to have different behavior and severity in the three regimes.
+The map-only estimator bias is hypothesized to be least severe in the error-in-$Y$ regime and most severe in the error-in-both regime.
+For example, if the map error is assumed to be additive noise, the map-only linear regression estimator will have no bias in the error-in-$Y$ regime [CarrolStefanski06] but in the error-in-$X$ regime it will often be biased towards zero (attenuation bias), and can sometimes be biased away from zero [MythsEpi5Paper, Kluger24BiasDirection].
+In the error-in-both regime, the bias issues from the error-in-$Y$ and error-in-$X$ regimes can aggregate.
+Examples
+All three of these regression settings are commonly seen in the remote sensing literature.
+We list a few examples here.
+For error-in-$Y$, studies have regressed remotely sensed algal blooms against fertilizer use in the US [taylor2023fertilizer], and remotely sensed maize and soybean yields against cover crop adoption in the US Corn Belt [deines2023recent].
+For error-in-$X$, studies have regressed crop yields against remotely sensed air pollution in China [he2022marked], and flood damages against remotely sensed wetland loss in the US [taylor2022wetlands].
+For error-in-both, studies have regressed remotely sensed maize and soybean yields against remotely sensed conservation tillage in the US Corn Belt [deines2019satellites], and remotely sensed crop greenness against remotely sensed nitrogen oxide levels in several agricultural regions [lobell2022globally].
+2.4 Evaluating performance using effective sample size
+We use effective sample size to compare the effectiveness of regression coefficient estimation methods.
+Suppose a GT-only confidence interval of width $w$ is generated using $n$ ground truth points.
+If another estimation method (such as PPI) uses $n$ ground truth points (along with a map product) to generate a confidence interval of width $w^{\prime}$, then the effective sample size of the new method is
+$n_{\text{effective}}=n\cdot\left(\frac{w}{w^{\prime}}\right)^{2}.$
+This is the number of ground truth points that would be required to generate a GT-only confidence interval of width $w^{\prime}$.
+3 Examples of regression coefficient estimation
+In this section, we apply PPI to estimate regression coefficients in four use cases associating deforestation, tree cover, housing price, and forest cover to other geospatial covariates.
+We compare point estimates and confidence intervals for the map-only estimator, GT-only estimator, and PPI estimator.
+Note that our regressions are not meant to contain a comprehensive list of covariates; our examples are primarily a means to illustrate the strengths and weaknesses of each method rather than meant to be interpreted scientifically.
+PPI can be applied when the response variable, the covariates, or both are remotely sensed, as long as each remotely sensed variable has a small amount of ground truth data for calibration.
+In the deforestation and tree cover regressions, the response variable is remotely sensed (error in $Y$), but ground truth data is widely available for all covariates.
+In the housing price regression, the response variable has ground truth widely available, but two of the covariates are remotely sensed (error in $X$).
+In the forest cover regression, both the response variable and one of the covariates are remotely sensed (error in both $X$ and $Y$).
+3.1 Experimental setup
+We first summarize the regression specifications and datasets for the four use cases. Maps for each use case are shown in Figure 3. For more detailed descriptions of the datasets, see Appendix C.
+Figure 3: Response variable and covariates for the four use cases.
+We consider use cases with remotely sensed response variable (Examples 1 and 2), remotely sensed covariates (Example 3), or both (Example 4).
+Example 1: Deforestation logistic regression
+We estimate the logistic regression coefficients for the association between Brazilian Amazon deforestation in 2000–2015 and several covariates: distance from major roads, distance from major rivers, elevation, and slope.
+We use logistic regression because the response variable (deforested or non-deforested) is binary.
+Given $n$ data points, the logistic regression model is
+$\log\frac{p}{1-p}=\beta_{0}+\beta_{\text{distRoad}}\textit{X}_{\text{distRoad}}+\beta_{\text{distRiver}}\textit{X}_{\text{distRiver}}+\beta_{\text{elevation}}\textit{X}_{\text{elevation}}+\beta_{\text{slope}}\textit{X}_{\text{slope}}$
+where $p$ is a $n\times 1$ vector of probabilities, $\textit{X}_{j}$ is a $n\times 1$ vector of covariate $j$, and $\beta_{j}$ is the scalar logistic regression coefficient for covariate $j$.
+We wish to estimate the $\beta_{j}$ giving the best fit to the above model.
+A one-unit increase in covariate $j$ corresponds to a $\beta_{j}$ change in the log-odds of the response variable.
+The scalar $\beta_{0}$ is the intercept, corresponding to the log-odds of the response when the covariates $\textit{X}_{j}$ are zero.
+The deforestation response variable is remotely sensed while the covariates are treated as ground truth.
+For remotely sensed deforestation, we create a binary deforestation map product using $N=963548$ pixels sampled uniformly at random from the NASA Global Forest Cover Change (GFCC) map [townshend2016global] (at 30 m resolution) in the study area in 2000 and 2015.
+For ground truth deforestation, we use $n=1386$ randomly sampled points from an Amazon deforestation dataset created by [bullock2020satellite] using time-series analysis of Landsat imagery and high-resolution imagery from Google Earth.
+For the covariates, we use a map of Brazilian federal roads in 2000 [de2021transport], the WWF HydroSHEDS Free Flowing Rivers Network map [grill2019mapping], and the NASA Digital Elevation Model [jpl2020].
+Example 2: Tree cover linear regression
+We estimate linear regression coefficients for the association between 2021 tree cover percentage in the contiguous United States and several covariates: aridity index, elevation, and slope.
+We use linear regression rather than logistic regression because the response variable (tree cover percentage) is continuous rather than binary.
+Given $n$ data points, the linear regression model is
+$Y=\beta_{0}+\beta_{\text{aridity}}\textit{X}_{\text{aridity}}+\beta_{\text{elevation}}\textit{X}_{\text{elevation}}+\beta_{\text{slope}}\textit{X}_{\text{slope}}+\varepsilon$
+where $Y$ is a $n\times 1$ vector of the response variable, $\textit{X}_{j}$ is a $n\times 1$ vector of covariate $j$, and $\beta_{j}$ is the scalar linear regression coefficient for covariate $j$.
+We wish to estimate the $\beta_{j}$ that give the best fit to the above model.
+A one-unit increase in covariate $j$ corresponds to a $\beta_{j}$ change in the response variable.
+The scalar $\beta_{0}$ is the intercept, corresponding to the response value when the covariates $\textit{X}_{j}$ are zero.
+The tree cover response variable is remotely sensed while the covariates are treated as ground truth.
+For remotely sensed tree cover, we use $N=983238$ map product points sampled uniformly at random from the 2021 USFS Tree Canopy Cover (TCC) product [forest2023usfs] at 30 m resolution.
+For ground truth tree cover, we manually label $n=983$ high-resolution satellite images from Google Maps (at randomly sampled locations) by visually estimating the proportion of the 30 m by 30 m image that is covered by trees.
+For the covariates, we use a map of the average Global Aridity Index from 1970-2000 [trabucco2019global] and the NASA Digital Elevation Model [jpl2020].
+Example 3: Housing price linear regression
+We estimate linear regression coefficients for the association between housing price in the United States and household income, nighttime light intensity, and road length.
+The linear regression model is
+$Y=\beta_{0}+\beta_{\text{income}}\textit{X}_{\text{income}}+\beta_{\text{nightlight}}\textit{X}_{\text{nightlight}}+\beta_{\text{roadLen}}\textit{X}_{\text{roadLen}}+\varepsilon$
+using similar notation as the previous linear regression example.
+Two of the covariates (nightlights and road length) are remotely sensed, but ground truth data is widely available for the response variable (housing price) and one covariate (income).
+For all variables, we use the MOSAIKS dataset [rolf2021generalizable] which contains a population-weighted random sample of points in the United States at roughly 1 km resolution.
+There are $N=46418$ points that have both ground truth and remotely sensed values for housing price (log dollars per square foot), household income (dollars), nighttime light intensity ($\log(1+\text{nW}/\text{cm}^{2}/\text{sr})$), and road length (km).
+For housing price and income, we use the ground truth values for all $46418$ points.
+For nightlights and road length, we take a uniform random subsample of $n=500$ points as the ground truth calibration set (and use remotely sensed values for the remaining points).
+For the GT-only method, we use only the $n=500$ ground truth points for all four variables.
+We also compare the regression coefficient point estimates from our methods against the “true coefficients” computed by running linear regression on the full set of $46418$ ground truth points for all four variables.
+Figure 4: Regression coefficient point estimates and 95% confidence intervals for the four use cases.
+PPI results in smaller 95% confidence intervals than using ground truth only.
+PPI effective sample size improvements (relative to GT-only) are shown in blue.
+For the two examples where the true coefficient is known, the bias in magnitude of each map-only estimate is shown in red.
+Example 4: Forest cover linear regression
+We estimate linear regression coefficients for the association between forest cover in the United States and two covariates: elevation and population. The linear regression model is
+$Y=\beta_{0}+\beta_{\text{elevation}}\textit{X}_{\text{elevation}}+\beta_{\text{population}}\textit{X}_{\text{population}}+\varepsilon$
+using similar notation as the previous linear regression examples.
+Both the response variable (forest cover) and one of the covariates (population) are remotely sensed, while ground truth data is widely available for the other covariate (elevation).
+Similar to the previous example, we use the MOSAIKS dataset [rolf2021generalizable].
+The dataset also provides a uniformly-weighted random sample of points in the United States at $\sim$1 km resolution.
+After removing points with missing variables, there are $N=67968$ points that have both ground truth and remotely sensed values for forest cover percentage, elevation (km), and population ($\log(1+\text{people})/\text{km}^{2}$).
+For elevation, we use the ground truth values for all $67968$ points.
+For forest cover and population, we take a uniform random subsample of $n=500$ points as the ground truth calibration set (and use remotely sensed values for the remaining points).
+For the GT-only method, we use only the $n=500$ ground truth points for all three variables.
+We also compare the regression coefficient point estimates from our methods against the “true coefficients” computed by running linear regression on the full set of $67968$ ground truth points for all four variables.
+3.2 Results
+The regression coefficient point estimates and 95% confidence intervals for all four use cases are shown in Figure 4.
+We compare results for the map-only, GT-only, and PPI estimators.
+(For all experiments in this section, we use $B=2000$ bootstrap iterations to construct PPI confidence intervals.)
+Map-only estimates are biased and their confidence intervals are too narrow.
+For all coefficients, the map-only confidence intervals are much narrower than those for GT-only and PPI.
+Unlike the GT-only and PPI intervals, the map-only intervals are not guaranteed with 95% probability to contain the true coefficients (i.e., the GT-only coefficients we would compute if we had access to ground truth data at every pixel).
+For instance, in Example 1 (deforestation logistic regression), the map-only estimate for the distance from river coefficient has a very narrow confidence interval around $-0.002$ and falls outside the PPI interval.
+Meanwhile, the GT-only and PPI confidence intervals are much wider, capturing the true uncertainty in the data, and both intersect with zero.
+Using only the maps would lead us to conclude that there is a significantly negative association between deforestation and distance from the nearest river, while the two methods that use ground truth points suggest that there may be no association.
+In Examples 3 and 4 (housing price and forest cover linear regressions), we have access to the true coefficients computed from the MOSAIKS dataset.
+We find that for all but one coefficient, the map-only confidence interval does not contain the true coefficient.
+We compute the bias (in percentages) of each map-only coefficient point estimate by subtracting the true coefficient from the map-only coefficient, then dividing the result by the true coefficient.
+Negative values of bias indicate the map-only coefficient is biased toward zero, and positive values indicate the map-only coefficient is biased away from zero.
+In our examples, the biases of the map-only coefficients range from $-18.9\%$ to $+64.3\%$.
+The single unbiased map-only estimate is for income in the housing price linear regression.
+While there are no guarantees, map-only estimates could happen to be unbiased for a variety of reasons.
+In the case of an error-in-$X$ linear regression, the estimated coefficients corresponding to a covariate for which only ground truth observations are used will be unbiased if that covariate is uncorrelated with both the map predictions and the map prediction errors of the other covariates (see Formula (5) in [Kluger24BiasDirection]).
+We suspect the map-only estimate for the income coefficient is unbiased in our experiment because (i) ground truth income data is used for all samples, (ii) income is uncorrelated with the errors in the nightlights and road length predictions, and (iii) income is only weakly correlated with the nightlights and road length predictions themselves.
+PPI uses ground truth data to correct for map bias.
+PPI gives an unbiased estimator, and the PPI confidence intervals have theoretical guarantees of containing the true coefficients with 95% probability.
+In Examples 3 and 4, the GT-only and PPI confidence intervals contain all of the true coefficients.
+PPI successfully uses the ground truth calibration set to correct for bias from the map product and account for uncertainty due to map product errors.
+PPI confidence intervals outperform the GT-only method.
+PPI uses the map product along with the ground truth to reduce uncertainty in the coefficient estimates, resulting in significantly smaller confidence intervals and larger effective sample sizes compared to the GT-only method.
+In our experiments, the PPI effective sample sizes range from 1.2×1.2$\times$1.2 × to 17.4×17.4$\times$17.4 × the ground truth calibration set size $n$.
+The largest PPI effective sample size improvement (17.4×17.4$\times$17.4 ×) is for the income coefficient in Example 3.
+We suspect that this is because the map-only income coefficient was essentially unbiased, so the PPI estimator for this coefficient involved a very small bias correction term, and therefore was similar to the low variance map-only estimator.
+The next largest PPI effective sample size improvements are for the elevation coefficient (6.2×6.2$\times$6.2 ×) in Example 4 and the aridity coefficient (6.5×6.5$\times$6.5 ×) and intercept (5.7×5.7$\times$5.7 ×) in Example 2 (tree cover linear regression).
+In contrast, adding map data did not reduce the confidence interval size significantly for the distance to river (1.2×1.2$\times$1.2 ×) and elevation (1.4×1.4$\times$1.4 ×) coefficients in Example 1.
+4 Simulating the role of map error
+To understand the effect of map product quality on PPI regression coefficient estimates, we simulate adding noise and bias to the map products in Example 4 (forest cover linear regression) from the previous section.
+We compare the resulting point estimates and 95% coefficient confidence intervals at different levels of map noise and bias using map-only, GT-only, and PPI estimators.
+4.1 Experimental Setup
+Similar to Example 4, we use the MOSAIKS dataset to regress forest cover against elevation and population, where forest cover and population are remotely sensed.
+We use the same $n=500$ randomly selected ground truth points as in the original experiment.
+However, instead of using the MOSAIKS map product datasets for forest cover and population, we take the full MOSAIKS ground truth dataset (67968 points) and use it to create simulated “map products” with different levels of noise and bias.
+We experiment with adding noise or bias to the population map (error-in-$X$), the forest cover map (error-in-$Y$), or both (error-in-both).
+No error is added to the elevation dataset.
+Simulating Noise
+For the ground truth population covariate $X_{\text{pop}}$, we define “noise at level $c$” as additive Gaussian noise with mean zero and standard deviation equal to $c$ times $\sigma_{X_{\text{pop}}}$, the standard deviation of all $X_{\text{pop}}$ values in the full ground truth dataset.
+The resulting simulated map product population value at each point is
+$X_{\text{pop}}^{\text{map}}=X_{\text{pop}}+c\cdot\mathcal{N}(0,\sigma^{2}_{X_{\text{pop}}}).$
+Similarly, for the forest cover response variable $Y$, the simulated map product value at each point is
+$Y^{\text{map}}=Y+c\cdot\mathcal{N}(0,\sigma^{2}_{Y}).$
+We experiment with noise levels $c$ from 0 to 1 at increments of 0.1. When $c=0$, the simulated map is the same as the full ground truth map.
+Simulating Bias
+[…]
+Then we create a simulated map product with “bias level $c$” by interpolating between $X_{\text{pop}}$ and $\hat{X}_{\text{pop}}^{\prime}$ to get
+$X_{\text{pop}}^{\text{map}}=c\hat{X}_{\text{pop}}^{\prime}+(1-c)X_{\text{pop}}.$
+Similarly, for the forest cover response variable $Y$, we use the MOSAIKS map product $Y^{\prime}$ to fit the square root regression $\hat{Y}^{\prime}=7.90\sqrt{Y}-3.20$.
+The simulated map product value at each point is
+$Y^{\text{map}}=c\hat{Y}^{\prime}+(1-c)Y.$
+We experiment with bias levels $c$ from 0 to 1 at increments of 0.1.
+When $c=0$, the simulated map is the same as the full ground truth map.
+When $c=1$, the simulated map is the same as the biased prediction $\hat{X}_{\text{pop}}^{\prime}$ or $\hat{Y}^{\prime}$.
+Figure 5: Simulation estimating forest cover linear regression coefficient 95% confidence intervals under different levels of map noise and map bias.
+In the bias simulations, $\hat{X}_{\text{pop}}^{\prime}$ and $\hat{Y}^{\prime}$ are nonlinearly biased proxies for true $X_{\text{pop}}$ and $Y$.
+Even with map noise or bias, PPI gives unbiased estimates with narrower confidence intervals than the GT-only method.
+Map-only estimates become increasingly biased as noise in the simulated map products for $X_{\text{pop}}$ increases or bias in the simulated map products for $X_{\text{pop}}$ or $Y$ increases.
+We also display the $R^{2}$ coefficient of determination for $X_{\text{pop}}^{\text{map}}$ at $c=1$ noise and bias levels.
+4.2 Results
+The coefficient point estimates and 95% confidence intervals at each noise and bias level are shown in Figure 5.
+We compare results for the map-only, GT-only, and PPI estimators.
+(For all experiments in this section, we use $B=200$ bootstrap iterations to construct PPI confidence intervals.)
+GT-only confidence interval is wide.
+First, the GT-only point estimate and confidence interval stay constant regardless of noise and bias, because they use a constant set of ground truth points.
+The confidence interval contains the true coefficients for both covariates ($-12$ for elevation and 1.4 for population).
+However, due to the small number of ground truth points ($n=500$), the GT-only interval for $\beta_{\text{pop}}$ also contains zero, so it does not detect a statistically significant association between forest cover and population.
+The GT-only interval is wider than the PPI and map-only intervals; exactly how much wider depends on the noise and bias levels.
+Both noise and bias lead to biased map-only estimates.
+Second, the map-only estimates suffer from bias when either noise or bias is added to the variables.
+In the noise experiments, the map-only estimates for $\beta_{\text{pop}}$ and $\beta_{\text{elevation}}$ become biased when there is noise in $X_{\text{pop}}$.
+As noise increases, the map-only estimates for $\beta_{\text{pop}}$ decrease in magnitude and become biased toward zero, because additive Gaussian noise in a covariate results in attenuation bias in the corresponding regression coefficient (see Section 2.3).
+The map-only estimates for $\beta_{\text{elevation}}$ conversely become larger in magnitude as noise in $X_{\text{pop}}$ increases.
+Furthermore, the map-only confidence intervals are narrow because the number of map product points is large ($N=67968$), but unlike the PPI intervals, they fail to capture the true estimation uncertainties due to map imperfections.
+The confidence intervals for $\beta_{\text{pop}}$ do not contain the true coefficient for $X_{\text{pop}}$ noise levels $c\geq 0.4$ (corresponding to $R^{2}\leq 0.92$ for the simulated population map).
+The confidence intervals for $\beta_{\text{elevation}}$ do not contain the true coefficient for $X_{\text{pop}}$ noise levels $c\geq 0.6$ ($R^{2}\leq 0.82$).
+By $X_{\text{pop}}$ noise levels of $c=0.8$ ($R^{2}=0.68$), the map-only estimate of $\beta_{\text{pop}}$ is roughly half of the true coefficient.
+In contrast, noise in $Y$ (forest cover) does not result in biased map-only estimates, because zero-centered noise in a response variable does not bias regression coefficients.
+Whether bias is in $X_{\text{pop}}$, $Y$, or both, the map-only estimates become biased.
+Similar to in the noise experiments, when we add bias the map-only confidence intervals remain narrow but fail to capture the actual estimation uncertainty.
+The map-only confidence intervals do not contain the true coefficient for $\beta_{\text{pop}}$ at bias levels $c\geq 0.2$ ($R^{2}\leq 0.997$ for simulated population and forest cover maps) or the true coefficient for $\beta_{\text{elevation}}$ at bias levels $c\geq 0.4$ ($R^{2}\leq 0.987$ for $X_{\text{pop}}$ and $R^{2}\leq 0.988$ for $Y$).
+When the bias level of $X_{\text{pop}}$ is at $c=1$ (in other words, $X_{\text{pop}}^{\text{map}}=\hat{X}_{\text{pop}}^{\prime}$)—corresponding to $R^{2}=0.92$ of the map compared to ground truth—treating the map as if it were 100% accurate would nearly double the estimate for $\beta_{\text{pop}}$.
+The direction of bias in coefficient estimates varies depending on the type of bias in the map and can be difficult to predict in general.
+In our bias experiments, the map-only estimates for $\beta_{\text{pop}}$ become biased away from zero while those for $\beta_{\text{elevation}}$ become biased toward zero.
+We reported the $R^{2}$ of map products above because $R^{2}$ is a common evaluation metric for remote sensing maps (MSE is another common metric, but its values are variable-specific).
+We draw attention to the fact that map $R^{2}$ can be high ($>0.9$)—in other words, the map appears very accurate—but the map-only estimator still yields significantly biased coefficients.
+Performance of PPI.
+Finally, our simulations allow us to see clearly the advantages and limitations of PPI.
+At almost all noise and bias levels across experiments, the PPI confidence intervals contain the true coefficients for both covariates.
+(The only exception is noise level $c=0.8$ for “noise in both,” corresponding to $R^{2}=0.68$ for $X_{\text{pop}}$ and $R^{2}=0.36$ for $Y$.) This is expected, as PPI uses the small set of ground truth points to correct for error and quantify uncertainty in the map product, and the PPI confidence intervals are guaranteed to contain the true coefficient with 95% probability.
+In contrast to using GT-only, PPI estimates $\beta_{\text{pop}}$ and $\beta_{\text{elevation}}$ much more precisely, allowing us to correctly conclude that there is a positive association between forest cover and population.
+The PPI estimate for $\beta_{\text{pop}}$ is only indistinguishable from zero at noise levels of $c\geq 0.7$ in $X_{\text{pop}}$ ($R^{2}\leq 0.75$), $c\geq 0.7$ in $Y$ ($R^{2}\leq 0.51$), or $c\geq 0.4$ in both ($R^{2}\leq 0.92$ for $X_{\text{pop}}$ and $R^{2}\leq 0.84$ for $Y$).
+The PPI estimates are statistically significant at all bias levels in our experiment.
+The effectiveness of PPI is dependent on the quality of the map product.
+When the noise or bias level is low, the PPI intervals are significantly smaller than the GT-only interval, because the simulated map product is close to ground truth and provides useful information for estimating the regression coefficients.
+As the noise level increases and the quality of the map deteriorates, the PPI intervals grow wider, approaching the GT-only interval.
+For example, when the noise level is $c=0.2$ in both $X_{\text{pop}}$ and $Y$ ($R^{2}=0.98$ for $X_{\text{pop}}$ and $R^{2}=0.96$ for $Y$), the PPI effective sample size outperforms the GT-only method by a factor of 18.8×18.8$\times$18.8 × for elevation and 11.2×11.2$\times$11.2 × for population.
+When the noise level increases to $c=0.9$ ($R^{2}=0.59$ for $X_{\text{pop}}$ and $R^{2}=0.19$ for $Y$), the PPI effective sample size only outperforms the GT-only method by 2.0×2.0$\times$2.0 × for elevation and $\times1$ for population.
+As the bias level increases, the PPI intervals grow wider but remain significantly smaller than the GT-only interval, because the biased map (which uses a square root transformation of the true variable values) still contains significant information about the true population and forest cover values.
+For example, when the bias level is $c=0.2$ in both $X_{\text{pop}}$ and $Y$ ($R^{2}=0.997$ for both), the PPI effective sample size outperforms the GT-only method by a factor of 133×133$\times$133 × for elevation and $\times7$ for population.
+When the bias level increases to $c=0.9$ ($R^{2}=0.94$ for both), the PPI effective sample size only outperforms the GT-only method by 20×20$\times$20 × for elevation and $\times1$ for population.
+5 Discussion
+Using remote sensing-based maps to draw scientific inferences requires uncertainty quantification and propagation of uncertainty to downstream analyses — otherwise, we risk making, in the words of [mcroberts2011satellite], just pretty pictures.
+In this work, we applied prediction-powered inference, using remote sensing map products along with a small amount of ground truth data, to estimate regression coefficients, while correctly taking into account the errors in the map products.
+We find that PPI estimates have narrower confidence intervals than the GT-only estimates.
+Unlike map-only estimates using the map product alone, the PPI estimates are guaranteed to be unbiased.
+Bias is the main issue for map-only estimates.
+Our experiments illustrate that a map-only regression coefficient estimate without bias correction can yield incorrect inferences.
+We see this clearly in our examples with MOSAIKS data in which we know the true coefficients.
+In Example 3 (housing price linear regression), the map-only estimator underestimates the intercept and overestimates the magnitude of the nightlight and road length coefficients.
+In Example 4 (forest cover linear regression), the map-only estimator underestimates the intercept, underestimates the magnitude of the elevation coefficient, and overestimates the magnitude of the population coefficient.
+In our map error simulations with forest cover linear regression (Section 4), the map-only estimates of elevation and population coefficients become increasingly biased as we add noise in $X$ or add bias in $X$ or $Y$ to the simulated maps.
+In all cases, the map-only estimator has very narrow confidence intervals that give a false sense of certainty.
+To address the bias of map-only estimators and under-coverage of their confidence intervals, we use PPI to incorporate ground truth data into our estimates.
+Performance of PPI.
+PPI uses the ground truth calibration data to correct the map product, resulting in unbiased regression coefficient estimates.
+Unlike map-only estimates, the PPI confidence intervals are guaranteed to have the appropriate coverage probability (e.g.
+95%) for the true coefficient of interest.
+This is seen in Examples 3 and 4 (housing price and forest cover linear regressions), where the PPI confidence intervals contain the true coefficients for all covariates.
+In our map error simulations, all but one of the PPI confidence intervals contain the true coefficient, even when there is a high level of noise or bias in the map.
+The width of PPI confidence intervals compared to the GT-only method depends on the quality of the map product.
+If the map product is fairly accurate (i.e., the map product predictions $(\hat{X},\hat{Y})$ are close to the ground truth values $(X,Y)$), the PPI regression coefficient confidence intervals are generally substantially smaller than the GT-only intervals.
+However, if the map product is low quality, the PPI estimator does not offer a significant advantage over the GT-only estimator.
+For instance, in our map noise simulations, the PPI confidence intervals become wider and approach the GT-only confidence intervals as noise increases and the quality of the map deteriorates.
+We also see in our experiments that the performance of PPI varies for different covariates.
+For instance, in Example 3, the PPI effective sample size outperforms the GT-only estimator by a factor of 17.4×17.4$\times$17.4 × for the income covariate, while the road length covariate has a more modest 1.5×1.5$\times$1.5 × improvement.
+Recommendations for map producers and users.
+Ultimately, the best methods combine map products with ground truth to yield the narrowest intervals (highest certainty).
+We recommend that producers of machine learning-generated remote sensing maps use a randomly sampled holdout ground truth dataset if they wish to estimate and construct confidence intervals for regression coefficients.
+The criterion for this holdout set is that the model, including during hyperparameter tuning, was not trained on it; a standard test set would be adequate, provided that it is randomly sampled and representative of the inference population.
+Map producers should release this ground truth data and details of its sampling scheme alongside their map product to make their map suitable for downstream scientific inferences.
+When there are privacy concerns for releasing ground truth data, map producers can alternatively collaborate with map users on federated approaches for PPI, such as those described in [MiaoLuNeurIPS].
+In the absence of this, we recommend that map users generate their own ground truth data for the variable of interest if possible (e.g., through manual inspection of remote sensing imagery).
+Limitations.
+PPI is limited to areas in which ground truth data is available.
+However, this would be necessary for any method that does not make further assumptions about the errors of the machine-learning model.
+Furthermore, the calibration dataset must be a sample from the ground-truth distribution of interest (this ensures that the PPI and GT-only estimators will both be unbiased).
+If the ground truth data sampling process is biased, the regression coefficient estimates using these ground truth points may not reflect the true relationships between variables.
+Together, these attributes mean that such methods may be especially straightforward to use when ground truth data can be generated de novo, such as from human inspection of available satellite imagery, and more challenging to use when ground truth data requires field data collection or survey data collection.
+The current theory for PPI focuses on the i.d.d.
+case, whereas in remote sensing we would usually have a fixed area of interest and then randomly sample a subset of the points in the area as the calibration set (design-based inference).
+In survey sampling it is known that the former often approximates the latter well when the sample is small relative to the total population, but admittedly this connection has not yet been made explicit for PPI.
+While PPI applies to settings where the ground truth sample is collected from a weighted random sample, a stratified random sample, or a clustered random sample [Kluger25GeneralizingPPI], we note that this work does not explore these common settings.
+In addition, PPI can be used for a variety of regression models beyond generalized linear models such as quantile regression (e.g., [Kluger25GeneralizingPPI, MiaoLuNeurIPS]), although due to space constraints we only consider generalized linear models in this paper.
+Acknowledgments
+This material is based upon work supported by the U.S.
+Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number(s) DE-SC0023112.
+This report was prepared as an account of work sponsored by an agency of the United States Government.
+Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights.
+Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof.
+The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
+[…]
+Appendix A Superpopulation-based versus design-based inference
+This paper focuses on settings where both the ground truth data and the data from the remote sensing map are assumed to be drawn from some superpopulation.
+This inference framework is in contrast to the design-based inference framework [sarndal2003model] that is commonly used in the remote sensing area estimation literature [Stehman2000, olofsson2014good], where the observed samples from the map data are presumed to constitute the entire population of interest.
+We focus on superpopulation-based inference approaches as opposed to design-based inference approaches for three reasons.
+First, it is common in remote sensing studies to downsample the data from a remote sensing map to save computational resources, especially when the goal is to fit a complex, computationally expensive regression model.
+For example, [ZhouEtAlCropRotationDownsamples] downsamples a crop type and crop yield map to fit a model relating crop rotation to crop yield for computational reasons (they find little change in their results across 20 different random downsamplings of the map products).
+Second, the PPI literature has thus far only developed methods that are theoretically justified in superpopulation-based inference settings.
+This is not a fundamental limitation of the PPI framework, and future work can extend PPI to design-based inference settings.
+Third, in our motivating applications $N$ is large, so the variances of the component PPI estimators $\hat{\beta}_{\text{calib}}$, $\hat{\gamma}_{\text{calib}}$, and $\hat{\gamma}_{\text{map}}$ will be similar whether or not one uses a design-based inference or a superpopulation-based inference framework (e.g.
+in design-based inference $\hat{\gamma}_{\text{map}}$ is viewed as a constant with 00 variance whereas in superpopulation-based inference $\hat{\gamma}_{\text{map}}$ has small variance due to the large sample size $N$).
+As a result, we expect that both frameworks will yield confidence intervals of similar widths, with the design-based inference confidence intervals being slightly smaller and the differences diminishing as $N\to\infty$.
+Appendix B Formulas for variance and tuning matrix of $\hat{\beta}_{\text{PPI}}$
+In this appendix, we present formulas for the covariance matrix of the proposed estimator $\hat{\beta}_{\text{PPI}}$ and its optimal tuning matrix.
+We focus on providing explicit formulas in the case of linear regression and logistic regression.
+The resulting formulas can be found in [chen2000unified], but for ease of exposition, we provide a slightly different derivation and more explicit expressions in our cases of interest.
+More formal derivations in a more general setting can be found in the appendix of [Kluger25GeneralizingPPI].
+Recall that $\hat{\gamma}_{\text{map}}$, $\hat{\beta}_{\text{calib}}$, and $\hat{\gamma}_{\text{calib}}$ are regression coefficient estimates from a generalized linear model (GLM), which could be a linear model or logistic regression model.
+Moreover, $\hat{\gamma}_{\text{map}}$ is estimated using all $N$ samples $(\hat{X}_{i}^{\prime},\hat{Y}_{i}^{\prime})_{i=1}^{N}$ of the map-based predictions.
+Meanwhile $\hat{\beta}_{\text{calib}}$ and $\hat{\gamma}_{\text{calib}}$ are estimated using a uniform random subsample of size $n$, which is a fraction $\rho=n/N$ of the $N$ samples.
+$\hat{\beta}_{\text{calib}}$ is estimated using ground truth points $(X_{i},Y_{i})_{i=1}^{n}$ and $\hat{\gamma}_{\text{calib}}$ is estimated using the corresponding proxies $(\hat{X}_{i},\hat{Y}_{i})_{i=1}^{n}$.
+For convenience we let $(X,Y)$ denote a random vector with the same distribution as $(X_{i},Y_{i})$ and $(\hat{X},\hat{Y})$ denote a random vector with the same distribution as the map-based predictions $(\hat{X}_{i},\hat{Y}_{i})$ or $(\hat{X}_{i}^{\prime},\hat{Y}_{i}^{\prime})$.
+Now let $\psi:\mathbb{R}\to\mathbb{R}$ be a function that specifies the type of GLM being run: $\psi(s)=\frac{1}{2}s^{2}$ for linear regression and $\psi(s)=\log(1+e^{s})$ for logistic regression with a canonical link function.
+Note since $\hat{\beta}_{\text{calib}}$ is a regression coefficient from a GLM, it estimates the well-defined quantity
+$\beta_{*}=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\mathbb{E}[-Y\beta^{\mathsf{T}}X+\psi(\beta^{\mathsf{T}}X)].$
+Similarly, $\hat{\gamma}_{\text{calib}}$ and $\hat{\gamma}_{\text{map}}$ estimate the well defined quantity
+$\gamma_{*}=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\mathbb{E}[-\hat{Y}\beta^{\mathsf{T}}\hat{X}+\psi(\beta^{\mathsf{T}}\hat{X})].$
+Based on asymptotic theory for M-estimators (e.g., Chapter 5 of [VanderVaartTextbook]) and since GLM estimators are M-estimators, under standard regularity conditions the 3 component estimators of $\hat{\beta}_{\text{PPI}}$ are asymptotically multivariate normal with
+$\begin{bmatrix}\hat{\beta}_{\text{calib}}\\ \hat{\gamma}_{\text{calib}}\\ \hat{\gamma}_{\text{map}}\end{bmatrix}\stackrel{{\scriptstyle\cdot}}{{\sim}}\mathcal{N}\Bigg{(}\begin{bmatrix}\beta_{*}\\ \gamma_{*}\\ \gamma_{*}\end{bmatrix},\frac{1}{n}\begin{bmatrix}D_{\beta}^{-1}C_{11}D_{\beta}^{-1}&D_{\beta}^{-1}C_{12}D_{\gamma}^{-1}&\rho D_{\beta}^{-1}C_{12}D_{\gamma}^{-1}\\ D_{\gamma}^{-1}C_{12}^{\mathsf{T}}D_{\beta}^{-1}&D_{\gamma}^{-1}C_{22}D_{\gamma}^{-1}&\rho D_{\gamma}^{-1}C_{22}D_{\gamma}^{-1}\\ \rho D_{\gamma}^{-1}C_{12}^{\mathsf{T}}D_{\beta}^{-1}&\rho D_{\gamma}^{-1}C_{22}D_{\gamma}^{-1}&\rho D_{\gamma}^{-1}C_{22}D_{\gamma}^{-1}\end{bmatrix}\Bigg{)}.$
+Above $\stackrel{{\scriptstyle\cdot}}{{\sim}}$ means approximately distributed as, $\mathcal{N}(\mu,\Sigma)$ denotes a multivariate normal distribution with mean vector $\mu$ and covariance matrix $\Sigma$, and the component matrices are given by
+$D_{\beta}=\mathbb{E}[\ddot{\psi}(\beta_{*}^{\mathsf{T}}X)XX^{\mathsf{T}}],\quad D_{\gamma}=\mathbb{E}[\ddot{\psi}(\gamma_{*}^{\mathsf{T}}\hat{X})\hat{X}\hat{X}^{\mathsf{T}}],\quad C_{11}=\mathbb{E}\big{[}\big{(}Y-\dot{\psi}(\beta_{*}^{\mathsf{T}}X)\big{)}^{2}XX^{\mathsf{T}}\big{]}$
+$C_{12}=\mathbb{E}\big{[}\big{(}Y-\dot{\psi}(\beta_{*}^{\mathsf{T}}X)\big{)}\big{(}\hat{Y}-\dot{\psi}(\gamma_{*}^{\mathsf{T}}\hat{X})\big{)}X\hat{X}^{\mathsf{T}}\big{]}\quad\text{and}\quad C_{22}=\mathbb{E}\big{[}\big{(}\hat{Y}-\dot{\psi}(\gamma_{*}^{\mathsf{T}}\hat{X})\big{)}^{2}\hat{X}\hat{X}^{\mathsf{T}}\big{]},$
+where $\dot{\psi}(\cdot)$ and $\ddot{\psi}(\cdot)$ give the first and second derivatives of the GLM class-specific function $\psi$. Now recall $\hat{\beta}_{\text{PPI}}$ takes the form
+$\hat{\beta}_{\text{PPI},\Omega}=\Omega\hat{\gamma}_{\text{map}}+(\hat{\beta}_{\text{calib}}-\Omega\hat{\gamma}_{\text{calib}}),$
+where $\Omega\in\mathbb{R}^{p\times p}$ is some tuning matrix.
+Note that by the previous asymptotic normality approximation and the formula for the distribution of a linear transformation of a multivariate Gaussian (i.e., using the fact that $AZ\sim\mathcal{N}(A\mu,A\Sigma A^{\mathsf{T}})$ when $A$ is a fixed matrix and $Z\sim\mathcal{N}(\mu,\Sigma)$),
+$\hat{\beta}_{\text{PPI},\Omega}=\begin{bmatrix}I_{p\times p}&-\Omega&\Omega\end{bmatrix}\begin{bmatrix}\hat{\beta}_{\text{calib}}\\ \hat{\gamma}_{\text{calib}}\\ \hat{\gamma}_{\text{map}}\end{bmatrix}\stackrel{{\scriptstyle\cdot}}{{\sim}}\mathcal{N}\Big{(}\beta_{*},\frac{1}{n}V(\Omega)\Big{)},$
+where
+$V(\Omega)=D_{\beta}^{-1}C_{11}D_{\beta}^{-1}-(1-\rho)\Omega[D_{\beta}^{-1}C_{12}D_{\gamma}^{-1}]^{\mathsf{T}}-(1-\rho)D_{\beta}^{-1}C_{12}D_{\gamma}^{-1}\Omega^{\mathsf{T}}+(1-\rho)\Omega D_{\gamma}^{-1}C_{22}D_{\gamma}^{-1}\Omega^{\mathsf{T}}.$
+Note that the asymptotic variance of each component of $\hat{\beta}_{\text{PPI}}$ is minimized when the diagonal entries of $V(\Omega)$ are minimized.
+Since $V(\Omega)$ is quadratic in $\Omega$ with the $j$th diagonal entry only depending on the $j$th row of $\Omega$, an exercise in linear algebra and multivariate calculus for quadratic functions shows that choosing
+$\Omega_{*}=D_{\beta}^{-1}C_{12}C_{22}^{-1}D_{\gamma},$
+simultaneously minimizes each diagonal component of $V(\Omega)$.
+Thus $\Omega_{*}$ is the optimal tuning matrix, and note that it is identical to the tuning matrix used in Equation (4) of [chen2000unified].
+Further note that
+$V(\Omega_{*})=D_{\beta}^{-1}C_{11}D_{\beta}^{-1}-(1-\rho)D_{\beta}^{-1}C_{12}C_{22}^{-1}C_{12}^{\mathsf{T}}D_{\beta}^{-1},$
+which matches the asymptotic variance found in Equation (5) of [chen2000unified].
+If $\hat{D}_{\beta}$, $\hat{D}_{\gamma}$, $\hat{C}_{11}$, $\hat{C}_{12}$ and $\hat{C}_{22}$ are data-based estimates of previously defined matrices that converge in probability to $D_{\beta}$, $D_{\gamma}$, $C_{11}$, $C_{12}$, and $C_{22}$, respectively, then choosing the tuning matrix
+$\hat{\Omega}=\hat{D}_{\beta}^{-1}\hat{C}_{12}\hat{C}_{22}^{-1}\hat{D}_{\gamma}$ (2)
+will be asymptotically optimal.
+In such settings $\hat{\Omega}$ will converge in probability to $\Omega_{*}$ and asymptotically, $\hat{\beta}_{\text{PPI}}$ will approximately follow a $\mathcal{N}\big{(}\beta_{*},\frac{1}{n}V(\Omega_{*})\big{)}$ distribution, in which case one can use an approximated covariance matrix for $\hat{\beta}_{\text{PPI}}$ given by
+$\hat{V}_{\text{PPI}}=\frac{1}{n}\Big{(}\hat{D}_{\beta}^{-1}\hat{C}_{11}\hat{D}_{\beta}^{-1}-\big{(}1-\frac{n}{N}\big{)}\hat{D}_{\beta}^{-1}\hat{C}_{12}\hat{C}_{22}^{-1}\hat{C}_{12}^{\mathsf{T}}\hat{D}_{\beta}^{-1}\Big{)}$ (3)
+and the normal approximation to calculate standard errors and confidence intervals for $\hat{\beta}_{\text{PPI}}$.
+Note that in this paper we use a bootstrap approach to constructing confidence intervals because it can generalize to other settings without having to find an explicit formula for $\hat{V}_{\text{PPI}}$.
+B.1 Explicit formulas in linear regression settings
+In linear regression settings $\psi(s)=\frac{1}{2}s^{2}$, so $\dot{\psi}(s)=s$ and $\ddot{\psi}(s)=1.$ Hence $D_{\beta}=\mathbb{E}[XX^{\mathsf{T}}]$, $D_{\gamma}=\mathbb{E}[\hat{X}\hat{X}^{\mathsf{T}}]$,
+$C_{11}=\mathbb{E}\big{[}(Y-\beta_{*}^{\mathsf{T}}X)^{2}XX^{\mathsf{T}}\big{]},C_{12}=\mathbb{E}\big{[}(Y-\beta_{*}^{\mathsf{T}}X)(\hat{Y}-\gamma_{*}^{\mathsf{T}}\hat{X})X\hat{X}^{\mathsf{T}}\big{]},\quad\text{and}\quad C_{22}=\mathbb{E}\big{[}(\hat{Y}-\gamma_{*}^{\mathsf{T}}\hat{X})^{2}\hat{X}\hat{X}^{\mathsf{T}}\big{]}.$
+Thus one can use the plug in estimators
+$\hat{D}_{\beta}=\frac{1}{n}\sum_{i=1}^{n}X_{i}X_{i}^{\mathsf{T}},\quad\hat{D}_{\gamma}=\frac{1}{n}\sum_{i=1}^{n}\hat{X}_{i}\hat{X}_{i}^{\mathsf{T}},\quad\hat{C}_{11}=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\hat{\beta}_{\text{calib}}^{\mathsf{T}}X_{i})^{2}X_{i}X_{i}^{\mathsf{T}},$
+$\hat{C}_{12}=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\hat{\beta}_{\text{calib}}^{\mathsf{T}}X_{i})(\hat{Y}_{i}-\hat{\gamma}_{\text{calib}}^{\mathsf{T}}\hat{X}_{i})X_{i}\hat{X}_{i}^{\mathsf{T}},\quad\text{and }\hat{C}_{22}=\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-\hat{\gamma}_{\text{calib}}^{\mathsf{T}}\hat{X}_{i})^{2}\hat{X}_{i}\hat{X}_{i}^{\mathsf{T}}.$
+One can then plug the above formulas into Equation (2) to obtain an asymptotically optimal tuning matrix choice.
+They can further plug these formulas into Equation (3) to estimate the variance of $\hat{\beta}_{\text{PPI}}$ with this tuning matrix choice (the estimated variance matrix of $\hat{\beta}_{\text{PPI}}$ can then be used to calculate standard errors and confidence intervals).
+B.2 Explicit formulas in logistic regression settings
+In logistic regression settings $\psi(s)=\log(1+e^{s})$, so $\dot{\psi}(s)=1/(1+e^{-s})$ and $\ddot{\psi}(s)=e^{s}/(1+e^{s})^{2}$.
+We can then similarly use the plug in estimators for $D_{\beta}$, $D_{\gamma}$, $C_{11}$, $C_{22}$, and $C_{12}$ given by
+$\hat{D}_{\beta}=\frac{1}{n}\sum_{i=1}^{n}\frac{\exp(\hat{\beta}_{\text{calib}}^{\mathsf{T}}X_{i})}{\big{(}1+\exp(\hat{\beta}_{\text{calib}}^{\mathsf{T}}X_{i})\big{)}^{2}}X_{i}X_{i}^{\mathsf{T}},\quad\hat{D}_{\gamma}=\frac{1}{n}\sum_{i=1}^{n}\frac{\exp(\hat{\gamma}_{\text{calib}}^{\mathsf{T}}\hat{X}_{i})}{\big{(}1+\exp(\hat{\gamma}_{\text{calib}}^{\mathsf{T}}\hat{X}_{i})\big{)}^{2}}\hat{X}_{i}\hat{X}_{i}^{\mathsf{T}},$
+$\hat{C}_{11}=\frac{1}{n}\sum_{i=1}^{n}\Big{(}Y_{i}-\frac{1}{1+\exp(-\hat{\beta}_{\text{calib}}^{\mathsf{T}}X_{i})}\Big{)}^{2}X_{i}X_{i}^{\mathsf{T}},\quad\hat{C}_{22}=\frac{1}{n}\sum_{i=1}^{n}\Big{(}\hat{Y}_{i}-\frac{1}{1+\exp(-\hat{\gamma}_{\text{calib}}^{\mathsf{T}}\hat{X}_{i})}\Big{)}^{2}\hat{X}_{i}\hat{X}_{i}^{\mathsf{T}}$
+$\text{and}\quad\hat{C}_{12}=\frac{1}{n}\sum_{i=1}^{n}\Big{(}Y_{i}-\frac{1}{1+\exp(-\hat{\beta}_{\text{calib}}^{\mathsf{T}}X_{i})}\Big{)}\Big{(}\hat{Y}_{i}-\frac{1}{1+\exp(-\hat{\gamma}_{\text{calib}}^{\mathsf{T}}\hat{X}_{i})}\Big{)}X_{i}\hat{X}_{i}^{\mathsf{T}}.$
+One can then plug the above formulas into Equation (2) to obtain an asymptotically optimal tuning matrix choice.
+They can further plug these formulas into Equation (3) to estimate the variance of $\hat{\beta}_{\text{PPI}}$ with this tuning matrix choice (the estimated variance matrix of $\hat{\beta}_{\text{PPI}}$ can then be used to calculate standard errors and confidence intervals).
+Appendix C Datasets
+We describe the datasets used in the four examples in the paper.
+C.1 Example 1: Deforestation logistic regression
+For the deforestation response variable, we use $N=963548$ map product pixels sampled uniformly at random from the NASA Global Forest Cover Change (GFCC) map [townshend2016global] (at 30 m resolution) in the study area.
+The GFCC product is derived from Landsat 5 and Landsat 7 surface reflectance imagery, and includes maps for tree cover percentage in each pixel in five-year increments from 2000 to 2015.
+For ground truth, we use a Amazon deforestation and forest disturbance dataset created by [bullock2020satellite] using time-series analysis of Landsat imagery and high-resolution imagery from Google Earth between 1995 and 2017.
+The dataset records the years corresponding to deforestation, degradation, and natural disturbance events at each ground truth point.
+We use this to construct a binary ground truth dataset restricted to Brazil, in which we consider a point deforested if it experienced a deforestation event between 2000 and 2015.
+The original dataset is obtained by stratified random sampling, using a stratification map (also created by [bullock2020satellite]) that contains 7 strata related to deforestation status.
+For our experiments, we construct a “uniformly random sampled” ground truth dataset of size $n=1386$ by downsampling the points in each of the strata to match its area proportion.
+Using the ground truth dataset, we find that for the “deforested” class, the remotely sensed map has 69% producer’s accuracy and 71% user’s accuracy.
+Binarizing the NASA GFCC map product for deforestation.
+The ground truth points from [bullock2020satellite] have binary deforestation values ($Y=1$ if deforested and $Y=0$ if not deforested between 2000-2015), while the NASA GFCC map product gives canopy cover percentages over time at each pixel.
+In order to binarize the continuous map product data, we classify a map pixel as “deforested” if its canopy cover percentage satisfies $\text{canopy}_{2015}-\text{canopy}_{2000}\leq-25\%$.
+We choose the $-25\%$ threshold by visually comparing the map-derived canopy cover change distributions for deforested and non-deforested ground truth points (Figure 6).
+Most of the deforested ground truth points have a map canopy cover decrease of more than $25\%$, while most of the non-deforested ground truth points have either a map canopy cover increase, no map canopy cover change, or a map canopy cover decrease of less than $25\%$.
+Figure 6: NASA GFCC map canopy cover change (2000-2015) distributions for non-deforested and deforested ground truth over $n=1386$ uniformly random sampled points.
+The distributions are visualized using kernel density estimation.
+We choose a -25% deforestation threshold (shown in red) to binarize the map product.
+For the covariates, we use a map of Brazilian federal roads in 2000 [de2021transport] and the WWF HydroSHEDS Free Flowing Rivers Network map [grill2019mapping] (we include all rivers of order at most 3) to compute the distance from each data point to the nearest major road and major river.
+We obtain elevation and slope from the NASA Digital Elevation Model [jpl2020] at 30 m resolution.
+Prior work has shown that deforestation in the Brazilian Amazon is more likely to occur near major roads and navigable rivers due to human settlement and economic activities along these transportation networks [laurance2002predictors, kirby2006future, rosa2013predictive].
+Elevation and slope have been associated with forest cover change; for example, deforestation is generally less likely on sloped terrain [sandel2013human].
+Studies have also regressed remotely sensed deforestation against other covariates such as climate, soil fertility, population, and protected area status [etter2006regional, laurance2002predictors].
+Additionally, there exist unofficial, non government built roads in Brazil that are not identified in the federal roads map but may be associated with deforestation [arima2016explaining, brandao2006mapping, perz2007unofficial, valle2024automated].
+However, in this paper, we do not use a comprehensive list of covariates that could influence deforestation, because the primary purpose of this regression is to illustrate the methods rather than explain deforestation.
+C.2 Example 2: Tree cover linear regression
+For the tree cover response variable, we use $N=983238$ map product points sampled uniformly at random from the 2021 USFS Tree Canopy Cover (TCC) product [forest2023usfs] at 30 m resolution.
+The TCC product is derived from Landsat and Sentinel-2 top-of-atmosphere reflectance data, USDA NASS CDL data, and terrain data.
+Its tree cover percentage predictions are obtained from random forest models trained on reference data from human-labeled high-resolution imagery.
+For ground truth, we manually label $n=983$ high-resolution satellite images from Google Maps for tree cover percentage, by visually estimating the proportion of the 30 m by 30 m image that is covered by trees.
+The image locations were sampled uniformly at random from the study area.
+The $R^{2}$ coefficient between the remotely sensed and ground truth tree cover percentage values is 0.68.
+For the covariates, we use a map of the average Global Aridity Index from 1970-2000 (at 30 arc-seconds resolution) [trabucco2019global] and the NASA Digital Elevation Model [jpl2020] (at 30 m resolution).
+The aridity index is defined as a ratio between precipitation and evapo-transpiration, and is lower for more arid areas; its value ranges from 0.02 to 6.55 in our dataset.
+Increased aridity is associated with lower tree cover (and decreased vegetation in general) [bastin2017extent, williams2010forest, berdugo2020global].
+Elevation and slope have also been associated with tree cover globally; there is generally more tree cover at lower elevations (e.g., due to higher temperatures) [mayor2017elevation] and on sloped terrain [sandel2013human].
+We do not include a comprehensive list of covariates that could influence tree cover, because the primary purpose of this regression is to illustrate uncertainty quantification rather than explain tree cover.
+C.3 Example 3: Housing price linear regression
+For all variables, we use the MOSAIKS dataset [rolf2021generalizable] which contains a population-weighted random sample of points in the United States at roughly 1 km resolution.
+The dataset contains ground truth values across seven variables, as well as corresponding remotely sensed predictions obtained from a machine learning model trained on a subset of the ground truth points.
+The model runs ridge regression on featurized representations of Google Static Maps satellite imagery.
+The MOSAIKS predictions generally capture the variables well, with $R^{2}$ ranging from 0.45 to 0.91 to depending on the variable.
+We note that [proctor2023parameter] also runs regressions on the MOSAIKS dataset using the ground truth and remotely sensed predictions, but unlike PPI, their coefficient estimation method relies on Bayesian modelling assumptions and does not provide frequentest statistical guarantees for e.g., 95% coverage of the true coefficients.
+The ground truth datasets are described in detail in the Supplementary Materials of [rolf2021generalizable].
+Housing price per square foot for sales since 2010 are obtained using the Zillow Transaction and Assessment Dataset.
+Median household income at the census block group level is obtained from the 2015 American Community Survey.
+Nighttime light intensity is obtained from average radiance values from the 2015 Visible Infrared Imaging Radiometer Suite (VIIRS).
+Total road length in each $\sim$1 km grid cell is computed from the United States Geological Survey National Transportation Dataset.
+After removing points with missing variables, there are $N=46418$ points that have both ground truth and remotely sensed values for housing price (log dollars per square foot), household income (dollars), nighttime light intensity ($\log(1+\text{nW}/\text{cm}^{2}/\text{sr})$), and road length (km).
+The $R^{2}$ coefficient between MOSAIKS predictions and ground truth is 0.85 for nightlights and 0.52 for road length.
+For housing price and income, we use the ground truth values for all $46418$ points.
+For nightlights and road length, we take a uniform random subsample of $n=500$ points as the ground truth calibration set (and use remotely sensed values for the remaining points).
+For the GT-only method, we use only the $n=500$ ground truth points for all four variables.
+We also compare the regression coefficient point estimates from our methods against the “true coefficients” computed by running linear regression on the full set of $46418$ ground truth points for all four variables.
+C.4 Example 4: Forest cover linear regression
+Similar to the previous example, we use the MOSAIKS dataset [rolf2021generalizable].
+The dataset also provides a uniformly-weighted random sample of points in the United States at $\sim$1 km resolution.
+After removing points with missing variables, there are $N=67968$ points that have both ground truth and remotely sensed values for forest cover percentage, elevation (km), and population ($\log(1+\text{people})/\text{km}^{2}$).
+The ground truth datasets are described in detail in the Supplementary Materials of [rolf2021generalizable].
+Forest cover percentage is obtained from the 2010 Hansen Global Forest Change maps [hansen2013high], which are derived from Landsat imagery.
+Elevation is obtained from Mapzen terrain tiles with data from NASA Jet Propulsion Laboratory’s Shuttle Radar Topography Mission.
+Population density is obtained from the Gridded Population of the World dataset, which uses the 2010 US Population and Housing Census.
+As in the previous example, the $N=67968$ remotely sensed predictions were obtained by [rolf2021generalizable] using the MOSAIKS prediction model, which runs ridge regression on featurized representations of Google Static Maps satellite imagery.
+The $R^{2}$ coefficient between the MOSAIKS predictions and ground truth is 0.91 for forest cover and 0.73 for population.
+For elevation, we use the ground truth values for all $67968$ points.
+For forest cover and population, we take a uniform random subsample of $n=500$ points as the ground truth calibration set (and use remotely sensed values for the remaining points).
+For the GT-only method, we use only the $n=500$ ground truth points for all three variables.
+We also compare the regression coefficient point estimates from our methods against the “true coefficients” computed by running linear regression on the full set of $67968$ ground truth points for all four variables.
+To generate the histogram in the right panel of Figure 2 of the main text, we compute PPI and GT-only coefficient estimates using 100 different uniform random subsamples of $n=500$ ground truth points from the full set of $67968$ ground truth points.
+We use the same set of $N=67968$ map product points in all 100 trials.
+Note that the map-only estimate remains the same in all trials.
+Appendix D PPI for mean estimation
+In this appendix, we describe how to use PPI to estimate the mean value of a remotely sensed variable.
+We describe ground truth-only, map-only, and PPI estimators.
+(Note that area estimation can be formulated as estimating the mean of a binary variable representing a land use/land cover class.
+See Appendix E for examples applying PPI for area estimation on remote sensing datasets.)
+Ground truth-only estimator.
+First, ground truth-only (GT-only) uncertainty quantification uses ground truth observations, sampled uniformly at random, to estimate the quantity of interest.
+For example, suppose there is a response variable $Y$ that can be observed over a given area, and we wish to generate a 95% confidence interval for its population mean $\theta^{\star}=E[Y]$.
 If we have $n$ ground truth observations $Y_{1},Y_{2},\dots Y_{n}$ sampled uniformly at random, we can compute the sample mean
-$\hat{\theta}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$
-and standard error
-$SE=\sqrt{\frac{1}{n}\mathrm{Var}(Y_{i})}=\frac{1}{n}\sqrt{\sum_{i=1}^{n}(Y_{i}-\hat{\theta})^{2}}$
-to produce the classical 95% confidence interval
-$C=(\hat{\theta}-1.96\cdot SE,\ \hat{\theta}+1.96\cdot SE).$
-This method is conservative — if the size $n$ of the ground truth dataset is small, the classical confidence interval will be large.
+$\hat{\theta}=\frac{1}{n}\sum_{i=1}^{n}Y_{i},\quad\text{which has standard deviation }\quad\sigma_{\hat{\theta}}=\sqrt{\frac{1}{n}\mathrm{Var}(Y_{i})},$
+and estimate the standard deviation with
+$\hat{\sigma}_{\hat{\theta}}=\frac{1}{n}\sqrt{\sum_{i=1}^{n}(Y_{i}-\hat{\theta})^{2}}$
+to produce the GT-only 95% confidence interval
+$C=(\hat{\theta}-1.96\cdot\hat{\sigma}_{\hat{\theta}},\ \hat{\theta}+1.96\cdot\hat{\sigma}_{\hat{\theta}}).$
+If the size $n$ of the ground truth dataset is small, which is often the case in environmental applications, the GT-only confidence interval will be large.
 Map-only estimator: Use only the remote sensing map.
 If we have a remote sensing data product with a large number $N$ of predictions $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{N}$ over the area of interest, we could treat these predictions like ground truth points and use the the formula above with $\hat{Y}$ in place of $Y$.
-This would produce a much smaller confidence interval, since we would use $N$ instead of $n$ in the denominator of the standard error.
+This would produce a much smaller confidence interval, since we would use $N\gg n$ in the denominator of the standard error.
 However, this map-only estimator approach may result in a biased estimate, and the confidence interval is not guaranteed to have a 95% coverage rate of the true parameter of interest.
 This is because the map-only estimator does not account for error or uncertainty in the model used to generate the map product predictions.
-2.2 Stratified estimator
+Prediction-Powered Inference: Map + ground truth.
+PPI [angelopoulos2023prediction] evaluates the accuracy of the map product by comparing the ground truth labels $Y_{1},Y_{2},\dots,Y_{n}$ to the map product predictions $\hat{Y}_{1},\hat{Y}_{2},\dots,\hat{Y}_{n}$ at the corresponding ground truth point locations.
+This calibration set of $n$ points is a simple or stratified random sample over the region of interest, and it must be separate from the training dataset used to train the machine learning model.
+PPI uses the calibration set to correct for bias in the map-only estimate computed from the $N$ unlabeled map points $\hat{Y}^{\prime}_{1},\hat{Y}^{\prime}_{2},\dots\hat{Y}^{\prime}_{N}$.
+For a concrete example, suppose we wish to estimate $\theta^{\star}=E[Y]$ as before. In the special case of estimating $\theta^{\star}=E[Y]$, the PPI point estimate is
+$\hat{\theta}_{\text{PPI}}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}}_{\text{map-only estimate}}-\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})}_{\text{bias correction}},$
+which coincides precisely with an estimator previously studied in the remote sensing literature (see Equation 4c of [MCROBERTS2022113168]).
+Assuming the map sample and calibration sample are independent, the standard deviation of this estimator is
+$\sigma_{\text{PPI}}$ $=\sqrt{\frac{1}{N}\mathrm{Var}(\hat{Y}_{i}^{\prime})+\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})}$
+$\approx\sqrt{\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})},$
+where the approximation holds when $N$ is much larger than $n$.
+Letting $\hat{\sigma}_{\text{PPI}}$ be an estimate of $\sigma_{\text{PPI}}$ based on plugging sample variance estimates into the formula for $\sigma_{\text{PPI}}$, the PPI 95% confidence interval is given by
+$C_{\text{PPI}}=(\hat{\theta}_{\text{PPI}}-1.96\cdot\hat{\sigma}_{\text{PPI}},\ \hat{\theta}_{\text{PPI}}+1.96\cdot\hat{\sigma}_{\text{PPI}}).$
+The first term in $\sigma_{\text{PPI}}$ is the variance corresponding to using only map predictions to estimate $\theta$, i.e.
+$\frac{1}{N}\mathrm{Var}(\hat{Y}_{i}^{\prime})$.
+The second term is the variance from the bias correction, $\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})$.
+As the number of map product points $N$ grows large, $\sigma_{\text{PPI}}$ approaches $\sqrt{\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})}$.
+This will be a realistic approximation in many cases, since remote sensing maps at a global, national, or province level can contain millions or even billions of pixels, whereas the number of ground truth points $n$ is commonly in the hundreds or thousands.
+The greater the accuracy of the map product, the smaller $\mathrm{Var}(\hat{Y}_{i}-Y_{i})$ will be and the smaller the PPI confidence interval will be.
+Note that the PPI confidence interval is smaller than the GT-only confidence interval if $\mathrm{Var}(\hat{Y}_{i}-Y_{i})<\mathrm{Var}(Y_{i})$.
+Like PPI, the post-stratified area estimator combines a map product with a small amount of ground truth data to make a precise and unbiased estimate.
+The post-stratified estimator and PPI produce similar outputs for area estimation; for more detailed comparisons of the two methods, see Appendices E and F.
+However, the post-stratified estimator only applies to area estimation, not regression coefficient estimation, while PPI can be applied to both (Figure 1 in the main text).
+Similarly, an estimator from [MCROBERTS2022113168] coincides precisely with the PPI estimator in the context of mean estimation, but does not apply to the setting of regression coefficient estimation.
+Appendix E Area estimation experiments
+In this appendix, we apply PPI and other methods to area estimation tasks with both ground truth and remotely sensed data.
+We first describe two well-known methods for area estimation (stratified estimator and post-stratified estimator).
+These methods were introduced in [cochran1977sampling] and are commonly used in the remote sensing community [olofsson2013making, olofsson2014good].
+PPI for area estimation (described in Appendix D) produces similar confidence intervals as the post-stratified estimator; for a mathematical comparison, see Appendix F.
+Then, we apply PPI, the stratified estimator, and the post-stratified estimator to two area estimation use cases (Iowa maize crop area and Brazilian Amazon deforestation area).
+We also compare results with the GT-only and map-only mean estimation methods described in Appendix D.
+E.1 Stratified estimator
 Stratified random sampling uses remotely sensed maps to determine a sampling strategy for ground truth points that reduces the variance of estimates.
 The stratified estimator is commonly used in the remote sensing community.
-We first use a map product to partition the area of interest into $K$ strata (sub-areas) such that the target variable is predicted to have low variance within each stratum.
+We first use a map product to partition the area of interest into $K$ strata (sub-areas) such that the response variable is predicted to have low variance within each stratum.
 Then we sample $n_{k}$ ground truth points uniformly at random from each stratum $1\leq k\leq K$.
-We can choose $n_{k}$ to be larger for strata predicted to have higher uncertainty to decrease the standard error within the strata and hence make the overall standard error smaller with the same total number of samples.
-Suppose we again wish to estimate $\theta^{\star}=E[Y]$ using a stratified random sample.
-For each stratum $k$, we compute the stratum sample mean $\hat{\theta}_{k}$ and stratum standard error $SE_{k}$ using the same formulae as the classical method.
-Suppose that in the map product, each stratum $k$ takes up proportion $10<A_{k}<$ 1 of the total area.
+We can choose $n_{k}$ to be larger for strata predicted to have higher uncertainty to decrease the standard error within the strata and hence make the overall standard error smaller with the same total number of sample units.
+Suppose we observe the response variable $Y$ over a given area, and we wish to estimate $\theta^{\star}=E[Y]$ using a stratified random sample.
+For each stratum $k$, we compute the stratum sample mean $\hat{\theta}_{k}$ and stratum standard error $\hat{\sigma}_{k}$ using the same formulae as the GT-only method from Appendix D.
+Suppose that in the map product, each stratum $k$ takes up proportion $10<A_{k}<$ 1 of the total area (with $\sum_{k=1}^{K}A_{k}=1$).
 Then, we compute the stratified mean
@@
 and stratified standard error
-$SE_{\text{strat}}=\sqrt{\sum_{k=1}^{K}A_{k}^{2}\cdot SE_{k}^{2}}$
+$\hat{\sigma}_{\text{strat}}=\sqrt{\sum_{k=1}^{K}A_{k}^{2}\cdot\hat{\sigma}_{k}^{2}}$
 to produce the stratified 95% confidence interval
-$C_{\text{strat}}=(\hat{\theta}_{\text{strat}}-1.96\cdot SE_{\text{strat}},\ \hat{\theta}_{\text{strat}}+1.96\cdot SE_{\text{strat}}).$
-The advantage of stratified estimation with an accurate map and well-chosen sampling strategy is that the variance within each stratum will be small, which results in a smaller overall confidence interval compared to the classical method.
+$C_{\text{strat}}=(\hat{\theta}_{\text{strat}}-1.96\cdot\hat{\sigma}_{\text{strat}},\ \hat{\theta}_{\text{strat}}+1.96\cdot\hat{\sigma}_{\text{strat}}).$
+The advantage of stratified estimation with an accurate map and well-chosen sampling strategy is that the variance within each stratum will be small, which results in a smaller overall confidence interval compared to the GT-only method.
 If the map predictions are inaccurate, then the ground truth points likely have higher variance within each map stratum, resulting in a larger confidence interval.
@@
 The strata are designed so that points within each stratum have low variance in deforestation status (for instance, most points in the “stable forest” and “stable non-forest” strata do not experience deforestation in the time period of interest).
-As a result, our stratified confidence interval for deforestation area is smaller than the classical confidence interval.
-Figure 3: Stratified estimation is a standard approach with remote sensing maps.
+As a result, our stratified confidence interval for deforestation area is smaller than the GT-only confidence interval.
+Figure 7: Stratified estimation is a commonly used approach with remote sensing maps.
 This example shows a map of the Brazilian Amazon used for stratification.
@@
 Adapted from [bullock2020satellite].
-2.3 Prediction-Powered Inference
-Prediction-powered inference (PPI) [angelopoulos2023prediction] is an uncertainty quantification method that computes confidence intervals using a small set of ground truth points and a much larger set of machine learning generated predictions.
-While stratified sampling requires us to commit to a map product and stratification scheme before sampling the ground truth points, the map product used in PPI can be chosen after sampling.
-The main idea is that PPI assesses the bias in the map product using the ground truth data and then corrects it.
-PPI evaluates the accuracy of the map product by comparing the ground truth labels $Y_{1},Y_{2},\dots Y_{n}$ to the map product predictions $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{n}$ at the corresponding ground truth point locations.
-For a concrete example, suppose we wish to estimate $\theta^{\star}=E[Y]$ as before.
-The PPI point estimate is
-$\hat{\theta}_{\text{PPI}}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}}_{\text{map-only estimate}}-\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})}_{\text{bias correction}}.$
-The PPI standard error is
-$SE_{\text{PPI}}$ $=\sqrt{\frac{1}{N}\mathrm{Var}(\hat{Y}_{i}^{\prime})+\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})}$
-$\approx\sqrt{\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})},$
-where the approximation holds when $N$ is much larger than $n$. This yields the PPI 95% confidence interval,
-$C_{\text{PPI}}=(\hat{\theta}_{\text{PPI}}-1.96\cdot SE_{\text{PPI}},\ \hat{\theta}_{\text{PPI}}+1.96\cdot SE_{\text{PPI}}).$
-The first term in $SE_{\text{PPI}}$ is the variance corresponding to using only map predictions to estimate $\theta$, i.e.
-$\frac{1}{N}\mathrm{Var}(\hat{Y}_{i}^{\prime})$.
-The second term is the variance from the bias correction, $\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})$.
-As the number of map product points $N$ grows large, $SE_{\text{PPI}}$ approaches $\sqrt{\frac{1}{n}\mathrm{Var}(\hat{Y}_{i}-Y_{i})}$.
-This will be a realistic approximation in many cases, since remote sensing maps at a global, national, or province level can contain millions or even billions of pixels, whereas the number of ground truth points $n$ is commonly in the hundreds or thousands.
-The greater the accuracy of the map product, the smaller $\mathrm{Var}(\hat{Y}_{i}-Y_{i})$ will be and the smaller the PPI confidence interval will be.
-Note that the PPI confidence interval is smaller than the classical confidence interval if $\mathrm{Var}(\hat{Y}_{i}-Y_{i})<\mathrm{Var}(Y_{i})$, which will be true for any map whose predictions correlate positively with the ground truth.
-PPI can also give confidence intervals for linear and logistic regression coefficients, but the formulae are more complex.
-We sketch the main idea here and refer the reader to [angelopoulos2023ppi++, angelopoulos2023prediction] for a full exposition.
-In brief, PPI begins like a map-only estimator by using a large set of map predictions $\hat{Y}^{\prime}_{1},\hat{Y}^{\prime}_{2},\dots\hat{Y}^{\prime}_{N}$ to compute a measure of fit $m_{\theta}$.
-Then, PPI computes a bias correction $\Delta_{\theta}$ by comparing the ground truth data $Y_{i}$ to the predictions $\hat{Y}_{i}$.
-The relevant formula for the measure of fit and bias correction depends on what we seek to estimate — areas, binary regression coefficients, linear regression coefficients, etc.
-— see Table 3.
-Lastly, the map-only $m_{\theta}$ and the bias correction $\Delta_{\theta}$ are combined to construct the PPI 95% confidence interval
-$C_{\text{PPI}}=\{\theta:|m_{\theta}+\Delta_{\theta}|\leq w_{\theta}\}$
-where $w_{\theta}$ is a constant that depends on the confidence level (95% in this case).
 PPI can be used with either uniform random sampling or stratified random sampling of the ground truth points [fisch2024stratified].
 Stratified (weighted) PPI can be applied as follows for mean estimation.
-Each ground truth value $Y_{i}$ and labeled prediction $\hat{Y}_{i}$ is multiplied by weight $w_{i}=A_{k}\frac{n}{n_{k}}$ where $k$ is the stratum where point $i$ is located, $A_{k}$ is the area proportion of stratum $k$ and $n_{k}$ is the number of ground truth samples in stratum $k$.
-Similarly, each unlabeled prediction $\hat{Y}^{\prime}_{i}$ in stratum $k$ is multiplied by weight $w_{i}=A_{k}\frac{N}{N_{k}}$ where $N_{k}$ is the number of map product points in stratum $k$.
-Then the same PPI formulae as above can be used.
-Estimand Measure of fit $m_{\theta}$ Bias correction $\Delta_{\theta}$
-Mean $\theta-\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}$ $\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})$
-Logistic regression $\frac{1}{N}\sum_{i=1}^{N}\left(\frac{X_{i}^{\prime}}{1+e^{-\theta^{T}X_{i}^{\prime}}}-\hat{Y}_{i}^{\prime}\right)$ $\frac{1}{n}\sum_{i=1}^{n}X_{i}\left(\hat{Y}_{i}-Y_{i}\right)$
-Linear regression $\theta-(X^{\prime})^{\dagger}\hat{Y}^{\prime}$ $X^{\dagger}\left(\hat{Y}-Y\right)$
-Table 3: Prediction-powered inference (PPI) formulae for three common estimation tasks found in remote sensing: mean estimation, logistic regression coefficients, and linear regression coefficients.
-We use the ppi-python software package to implement the unweighted and weighted versions of PPI for our use cases.
-The software includes a tuning parameter $\lambda$ that we set using the map product points and ground-truth points.
-The confidence intervals are unbiased even with this data-driven tuning [angelopoulos2023ppi++].
-2.4 Post-stratified area estimator
+Each ground truth value $Y_{i}$ and labeled prediction $\hat{Y}_{i}$ is multiplied by weight $w_{i}=A_{k}\frac{n}{n_{k}}$ where $k$ is the stratum where point $i$ is located, $A_{k}$ is the area proportion of stratum $k$, $n_{k}$ is the number of ground truth sample units in stratum $k$, and $n=\sum_{i=1}^{K}n_{k}$.
+(Note that the weights are designed such that the sum of weights of all $n_{k}$ ground truth points in stratum $k$ is $A_{k}n$, and the sum of weights of all $n$ ground truth points is $\sum_{i=1}^{K}A_{k}n=n$.
+If the points were uniformly randomly sampled, this would be equivalent to only having a single “stratum,” and each point would have weight 1.) Similarly, each unlabeled prediction $\hat{Y}^{\prime}_{i}$ in stratum $k$ is multiplied by weight $w_{i}=A_{k}\frac{N}{N_{k}}$ where $N_{k}$ is the number of map product points in stratum $k$ and $N=\sum_{i=1}^{K}N_{k}$.
+Then the same PPI formulae as in Appendix D can be used.
+E.2 Post-stratified area estimator
 The post-stratified area estimator, commonly used in the remote sensing community, utilizes the confusion matrix between ground truth and map product land cover labels to estimate the area of a land cover class [card1982using, stehman2013estimating, olofsson2013making].
@@
 This can be viewed as a correction that uses the map errors relative to the ground truth points to modify the map-only area estimates.
-We show mathematically in Appendix A that PPI and the post-stratified area estimator result in similar confidence intervals as the number of map product points $N$ grows large.
+We show mathematically in Appendix F that PPI and the post-stratified area estimator result in similar confidence intervals as the number of map product points $N$ grows large.
 However, unlike PPI, the post-stratified area estimator only applies to area estimation but not regression coefficient estimation.
@@
 For each $1\leq i\leq K$ and $1\leq j\leq K$, let $n_{ij}$ denote the number of points that are in map class $i$ and ground truth class $j$.
+Also let $n_{i\cdot}=\sum_{j=1}^{K}n_{ij}$ denote the total number of ground truth samples that have map class $i$.
 For each class $1\leq i\leq K$, let $10<A_{i}<$ 1 denote the proportion of the total map area taken up by the map class.
-Then, the post-stratified area estimate for the area of class $j$ is
+Then, motivated by the formula $\mathbb{P}(Y=j)=\sum_{i=1}^{K}\mathbb{P}(\hat{Y}=i)\cdot\mathbb{P}(Y=j\ |\ \hat{Y}=i)$, the post-stratified area estimate for the proportion of area of class $j$ is
 $\hat{\theta}_{j,\text{post}}=\sum_{i=1}^{K}A_{i}\frac{n_{ij}}{n_{i\cdot}}$
 with corresponding standard error
-$SE_{j,\text{post}}=\sqrt{\sum_{i=1}^{K}A_{i}^{2}\frac{\frac{n_{ij}}{n_{i\cdot}}(1-\frac{n_{ij}}{n_{i\cdot}})}{n_{i\cdot}}}.$
-The 95% confidence interval for the area of class $j$ is
-$C_{j,\text{post}}=(\hat{\theta}_{j,\text{post}}-1.96\cdot SE_{j,\text{post}},\ \hat{\theta}_{j,\text{post}}+1.96\cdot SE_{j,\text{post}}).$
+$\hat{\sigma}_{j,\text{post}}=\sqrt{\sum_{i=1}^{K}A_{i}^{2}\frac{\frac{n_{ij}}{n_{i\cdot}}(1-\frac{n_{ij}}{n_{i\cdot}})}{n_{i\cdot}}}.$
+The 95% confidence interval for the proportion of area of class $j$ is
+$C_{j,\text{post}}=(\hat{\theta}_{j,\text{post}}-1.96\cdot\hat{\sigma}_{j,\text{post}},\ \hat{\theta}_{j,\text{post}}+1.96\cdot\hat{\sigma}_{j,\text{post}}).$
 Area estimation is a special case of mean estimation in which we estimate the mean of a binary variable that is $1$ for all points in class $j$ and 00 for all points outside of class $j$.
 Thus, we can also perform area estimation using methods such as PPI and stratified estimation.
-If the $n$ ground truth points are obtained by stratified random sampling using the $K$ map land cover classes as the strata, the post-stratified confidence interval is identical to the stratified confidence interval.
-2.5 Evaluating performance using effective sample size
-We use effective sample size to compare the effectiveness of uncertainty quantification methods.
-Suppose a classical confidence interval of width $w$ is generated using $n$ ground truth points.
-If another uncertainty quantification method uses $n$ ground truth points (along with a map product) to generate a confidence interval of width $w^{\prime}$, then the effective sample size of the new method is
-$n_{\text{effective}}=n\cdot\left(\frac{w}{w^{\prime}}\right)^{2}.$
-This is the number of ground truth points that would be required to generate a classical confidence interval of width $w^{\prime}$.
-3 Area estimation
+If the $n$ ground truth points are obtained by stratified random sampling using the $K$ map land cover classes as the strata, the post-stratified confidence interval computed using this data is identical to the stratified confidence interval [olofsson2013making].
+However, unlike the stratified estimator, the post-stratified estimator can be applied even if the ground truth points are obtained using uniform random sampling.
+E.3 Area estimation examples
 In this section, we demonstrate uncertainty quantification for two area estimation use cases.
@@
 We then use real ground truth data and real map products developed by researchers to estimate change in tree cover in the Brazilian Amazon in 2000–2015.
-We compare the classical, map-only (pixel-counting), stratified, post-stratified, PPI, and stratified PPI estimators.
+We compare the GT-only, map-only (pixel-counting), stratified, post-stratified, PPI, and stratified PPI estimators.
 In each use case, we use binary data where each point has value 1 if it belongs to the land cover class of interest (e.g., cropland or deforested land), and value 0 otherwise.
 Then, estimating the area fraction of the land cover class is equivalent to estimating the mean of all points in the region.
-3.1 Simulation: Maize area in Iowa
+We use the ppi-python software package to implement the unweighted and weighted versions of PPI for our area estimation use cases.
+The software includes a tuning parameter $\lambda$ that we set using the map product points and ground-truth points.
+The confidence intervals are unbiased and guaranteed to have the appropriate (e.g., 95%) coverage of the true area fraction even with this data-driven tuning [angelopoulos2023ppi++].
+Data and code for these use cases are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
+E.3.1 Simulation: Maize area in Iowa
 Problem Formulation
@@
 We then experiment with degrading the map product quality as follows.
-We systematically add noise or bias to the map product points while keeping the ground truth points constant, and then compare our maize fraction estimates across the map-only (pixel-counting), classical, and PPI estimators.
+We systematically add noise or bias to the map product points while keeping the ground truth points constant, and then compare our maize fraction estimates across the map-only (pixel-counting), GT-only, and PPI estimators.
 We average results over 100 trials.
 Adding unbiased noise.
-We define adding unbiased noise at level $p$ to the map product as resampling each of the 100000 map product points with probability $p$ from the Bernoulli(0.35) distribution.
+We define adding “unbiased noise” at level $p$ to the map product as resampling each of the 100000 map product points with probability $p$ from the Bernoulli(0.35) distribution.
 In other words, about $p$ fraction of the points are randomly labeled, while the remaining $1-p$ points are unchanged.
@@
 This is what we expect, since our noise-adding process resamples from the Bernoulli(0.35) distribution, causing the overall fraction of maize to remain the same.
-The classical confidence interval stays constant because it only uses ground truth points.
+The GT-only confidence interval stays constant because it only uses ground truth points.
 The map-only confidence intervals are small because $N=100000$ is large, but they do not account for error in the map product.
 However, the map-only confidence intervals remain unbiased because the map still has true maize fraction close to 0.35 regardless of noise level.
-When the noise probability is low, the PPI confidence intervals are significantly smaller than the classical intervals.
+When the noise probability is low, the PPI confidence intervals are significantly smaller than the GT-only intervals.
 This is because the error of the map product ($\hat{Y}-Y$) is small at the ground truth locations, so the variance of this error is also small, resulting in a small PPI standard error.
-As noise increases, the variance of the error of the map product increases; the PPI confidence intervals grow wider, eventually converging to the classical confidence interval at noise probability $p=1$.
-(At $p=1$, the map is complete noise and therefore does not provide any information for estimating the area beyond the information contained in the ground truth samples.) Consequently, the PPI effective sample size decreases from $n_{\text{effective}}=578$ when $p=0.1$ to $n_{\text{effective}}=100$ when $p=1$ (which matches the number of actual ground truth points $n=100$).
+As noise increases, the variance of the error of the map product increases; the PPI confidence intervals grow wider, eventually converging to the GT-only confidence interval at noise probability $p=1$.
+(At $p=1$, the map is complete noise and therefore does not provide any information for estimating the area beyond the information contained in the ground truth sample.) Consequently, the PPI effective sample size decreases from $n_{\text{effective}}=578$ when $p=0.1$ to $n_{\text{effective}}=100$ when $p=1$ (which matches the number of actual ground truth points $n=100$).
 Thus, the effectiveness of PPI is dependent on the quality of the map product.
@@
 The estimated maize fraction confidence intervals after adding bias to the map product are shown in Figure 10.
-Again, the classical confidence interval stays constant and remains centered at the true maize fraction because it only uses the ground truth points.
+Again, the GT-only confidence interval stays constant and remains centered at the true maize fraction because it only uses the ground truth points.
 As bias increases, the fraction of map product points labeled as “maize” increases.
@@
 In contrast, the PPI estimates remain centered at the true maize fraction because PPI corrects for bias in the map product.
-The PPI confidence intervals are smaller than the classical intervals because the variance of the map product error $\hat{Y}-Y$ is small (since only about 40% of the map product points are resampled, regardless of bias level).
+The PPI confidence intervals are smaller than the GT-only intervals because the variance of the map product error $\hat{Y}-Y$ is small (since only about 40% of the map product points are resampled, regardless of bias level).
 Thus, PPI provides unbiased estimates with lower uncertainty than using only the ground truth points.
 Again, the post-stratified confidence intervals and effective sample sizes are nearly identical to those of PPI, as expected.
-We also plot the maize fraction confidence intervals against model accuracy for both the noise and bias experiments (Figure 7).
-As noise increases, model accuracy decreases, so low model accuracy is associated with wide PPI and post-stratified confidence intervals.
+We also plot the maize fraction confidence intervals against overall map accuracy for both the noise and bias experiments (Figure 11).
+As noise increases, map accuracy decreases, so low map accuracy is associated with wide PPI and post-stratified confidence intervals.
 Thus, the left panel of Figure 11 resembles an inverted version of the left panel of Figure 9.
-Figure 7: Simulation estimating maize fraction 95% confidence intervals at different map product accuracy levels obtained by varying noise (left) and bias (right), using $N=100000$ map product points and $n=100$ “ground truth” points.
+Figure 11: Simulation estimating maize fraction 95% confidence intervals at different map product overall accuracy levels obtained by varying noise (left) and bias (right), using $N=100000$ map product points and $n=100$ “ground truth” points.
 Post-stratified and PPI confidence intervals are overlapping.
-As bias increases and the fraction of map product points labeled as “maize” increases, model accuracy decreases linearly.
-(To see this, note that if each resampled point is labeled as maize with probability $p$, then an expected $0.35(p)+0.65(1-p)=0.65-0.3p$ fraction of the resampled points will match the ground truth.) Thus, the map-only estimates of maize fraction decrease linearly as model accuracy increases.
-In the right panel of Figure 7, a model accuracy of 0.86 corresponds to a map-only maize fraction estimate of 0.21, significantly lower than the true fraction of 0.35.
-This shows that even a map with high model accuracy can yield biased estimates under the map-only (pixel-counting) estimator.
-3.2 Deforestation area in Brazilian Amazon
+As bias increases and the fraction of map product points labeled as “maize” increases, map accuracy decreases linearly.
+(To see this, note that if each resampled point is labeled as maize with probability $p$, then an expected $0.35(p)+0.65(1-p)=0.65-0.3p$ fraction of the resampled points will match the ground truth.) Thus, the map-only estimates of maize fraction decrease linearly as map accuracy increases.
+In the right panel of Figure 11, a map accuracy of 0.86 corresponds to a map-only maize fraction estimate of 0.21, significantly lower than the true fraction of 0.35.
+This shows that even a map with high map accuracy can yield biased estimates under the map-only (pixel-counting) estimator.
+E.3.2 Deforestation area in Brazilian Amazon
 Problem Formulation
@@
 The GFCC product is derived from Landsat 5 and Landsat 7 surface reflectance imagery, and includes maps for tree cover percentage in each pixel in five-year increments from 2000 to 2015.
-We consider a map pixel deforested if its canopy cover percentage satisfies $\text{canopy}_{2015}-\text{canopy}_{2000}\leq-25\%$ (Appendix B).
+We consider a map pixel deforested if its canopy cover percentage satisfies $\text{canopy}_{2015}-\text{canopy}_{2000}\leq-25\%$ (Figure 6 in Appendix C).
 We use a ground truth Amazon deforestation and forest disturbance dataset created by [bullock2020satellite] using time-series analysis of Landsat imagery and high-resolution imagery from Google Earth between 1995 and 2017.
@@
 The original dataset is obtained by stratified random sampling, using a stratification map (also created by [bullock2020satellite]) that contains 7 strata related to deforestation status.
-To generate a classical estimate baseline, we construct a “uniformly random sampled” ground truth dataset by downsampling the points in each of the strata to match its area proportion.
+To generate a GT-only estimate baseline, we construct a “uniformly random sampled” ground truth dataset by downsampling the points in each of the strata to match its area proportion.
 For a fair comparison, we also downsample the original stratified ground truth dataset to match the size of the “uniformly random” dataset.
@@
 Figure 13: Brazilian Amazon deforestation 95% confidence intervals using $N=4.4\times 10^{9}$ map product points and $n=1386$ ground truth points.
-The stratified, post-stratified, and PPI estimators all result in smaller confidence intervals for estimated deforestation than classical estimation, with Stratified PPI having the best performance.
+The stratified, post-stratified, and PPI estimators all result in smaller confidence intervals for estimated deforestation than GT-only estimation, with the stratified estimator and stratified PPI having the best performances.
 Each method results in a point estimate that between 5% to 6% of the Brazilian Amazon was deforested between 2000 and 2015.
-The stratified, post-stratified, and PPI estimators all result in smaller 95% confidence intervals for estimated deforestation than classical estimation (Figure 9).
-The post-stratified and PPI confidence intervals are fairly similar, with respective effective sample sizes that are $1.9$ and $1.8$ times the actual number of ground truth samples.
-The stratified and stratified PPI confidence intervals have the best performance, with effective sample sizes that outperform the classical method by a factor of about $2.4$.
-(Put differently, stratified PPI has effective sample size $3382$, so using the remotely sensed deforestation map added the equivalent of $3382-1386=1996$ ground truth samples.) Thus, the methods that use both map product and ground truth reduce uncertainty in our estimates, compared to using only ground truth.
+The stratified, post-stratified, and PPI estimators all result in smaller 95% confidence intervals for estimated deforestation than GT-only estimation (Figure 13).
+The post-stratified and PPI confidence intervals are fairly similar, with respective effective sample sizes that are $1.9$ and $1.8$ times the actual number of ground truth sample points.
+The stratified and stratified PPI confidence intervals have the best performance, with effective sample sizes that outperform the GT-only method by a factor of about $2.4$.
+(Put differently, stratified PPI has effective sample size $3382$, so using the remotely sensed deforestation map added the equivalent of $3382-1386=1996$ ground truth sample points.) Thus, the methods that use both map product and ground truth reduce uncertainty in our estimates, compared to using only ground truth.
 In particular, the stratified estimator is effective because the strata map is designed so that points within each stratum have low variance in deforestation status.
@@
 However, we emphasize that this would not have been clear without the ground truth data, and ground truth data is still necessary to check the reliability of the map product for deforestation analysis.
-4 Estimating regression coefficients
-In this section, we apply our uncertainty quantification methods to estimating regression coefficients for associative relationships between a remotely sensed outcome and multiple covariates.
-Regressions are commonly used in the environmental sciences and economics to identify relationships — associative or causal — between covariates and a binary outcome (logistic regression) or between covariates and a continuous outcome (linear regression).
-Logistic regression coefficients give the change in the log odds ratio of a binary outcome corresponding to a one-unit increase in a covariate.
-Linear regression coefficients give the change in the value of a continuous outcome corresponding to a one-unit increase in a covariate.
-Thus, we can interpret these coefficients as effect sizes of the covariates.
-In particular, we estimate (1) logistic regression coefficients for Brazilian Amazon deforestation in 2000–2015, and (2) linear regression coefficients for US tree cover in 2021.
-We compare the classical method, map-only estimator, PPI, stratified estimator, and stratified PPI.
-Our methods only account for uncertainty in the remotely sensed target variable, while ignoring uncertainty in the covariates; we leave addressing the latter to future work.
-4.1 Deforestation logistic regression
-Problem Formulation
-We estimate the logistic regression coefficients for the association between Brazilian Amazon deforestation in 2000–2015 and several covariates: distance from major roads, distance from major rivers, elevation, and slope (Figure 10).
-We use logistic regression because the outcome variable (deforested or non-deforested) is binary.
-Given $n$ data points, the logistic regression model is
-$\log\frac{p}{1-p}=\theta_{\text{road}}\textit{x}_{\text{road}}+\theta_{\text{river}}\textit{x}_{\text{river}}+\theta_{\text{elevation}}\textit{x}_{\text{elevation}}+\theta_{\text{slope}}\textit{x}_{\text{slope}}$
-where $p$ is a $n\times 1$ vector of probabilities, $\textit{x}_{j}$ is a $n\times 1$ vector of covariate $j$, and $\theta_{j}$ is the scalar logistic regression coefficient for covariate $j$.
-We wish to estimate the $\theta_{j}$.
-A one-unit increase in covariate $j$ corresponds to a $\theta_{j}$ change in the log-odds of the outcome variable.
-We use the same Brazilian Amazon deforestation datasets as the previous section, with $n=1386$ ground truth points and $N=963548$ uniformly randomly sampled map product points.
-For the covariates, we use a map of Brazilian federal roads in 2000 [de2021transport] and the WWF HydroSHEDS Free Flowing Rivers Network map [grill2019mapping] (we include all rivers of order at most 3) to compute the distance from each data point to the nearest major road and major river.
-We obtain elevation and slope from the NASA Digital Elevation Model [jpl2020] at 30 m resolution.
-Prior work has shown that deforestation in the Brazilian Amazon is more likely to occur near major roads and navigable rivers due to human settlement and economic activities along these transportation networks [laurance2002predictors, kirby2006future, rosa2013predictive].
-Elevation and slope have also been associated with forest cover change; for example, deforestation is generally less likely on sloped terrain [sandel2013human].
-We do not include a comprehensive list of covariates that could influence deforestation, because the primary purpose of this regression is to illustrate uncertainty quantification rather than explain deforestation.
-Figure 10: Brazilian Amazon deforestation logistic regression datasets.
-We regress the deforestation binary target variable (left) against four continuous covariates (right): distance from roads, distance from rivers, elevation, and slope.
-Results
-The 95% confidence intervals for the four logistic regression coefficients are shown in Figure 11.
-The general degree of uncertainty in estimating the coefficients from our non-map-only methods are as follows.
-The log odds of deforestation change by about .$-.016$ to .$-.010$ for each one-kilometer increase in distance from roads, .$-.008$ to .$-.001$ for each one-kilometer increase in distance from rivers, $-3$ to $0.8$ for each one-kilometer increase in elevation, and $-0.28$ to $-0.11$ for each one-degree increase in slope.
-Thus, deforestation is negatively associated with all covariates except for elevation.
-The elevation coefficient confidence intervals intersect with zero for all non-map-only methods except for stratified PPI, suggesting that there might be no association between elevation and deforestation after accounting for the other three variables.
-The PPI methods result in the smallest confidence intervals, with effective sample size outperforming the classical method by a factor of 1.1 to 2.5 depending on the covariate.
-The most significant improvement is for the slope covariate.
-Unlike in our previous area estimation examples, the stratified estimator (i.e., weighted logistic regression with weights determined by the strata) is not effective and even reduces effective sample size relative to the classical method when estimating the coefficients for elevation and distance from rivers.
-Similarly, stratified PPI has slightly worse performance than PPI for all covariates except distance from roads.
-The strata were designed to minimize variance in deforestation status, which does not necessarily correspond to better estimates of regression coefficients for the covariates.
-For slope and distance from rivers, the map-only estimated coefficient computed from the map product falls outside the confidence intervals for all methods that use ground truth points.
-The map-only coefficient for slope is about 0, and the map-only coefficient for distance from rivers is about $-0.010$.
-In particular, for the slope coefficient, the methods that use ground truth points find an association that is statistically significantly different from zero.
-However, if we had used only the remote sensing map, we would not have discovered the negative relationship between slope and deforestation.
-Thus, using only the map product could result in misleading conclusions due to biases in the deforestation map product that are correlated with covariates (e.g., the map product overestimates deforestation at locations with high slope, or underestimates deforestation at locations with low slope).
-Figure 11: Brazilian Amazon deforestation logistic regression coefficient 95% confidence intervals using $N=963548$ map product points and $n=1386$ ground truth points.
-The PPI methods result in the smallest confidence intervals, while stratification is not generally effective.
-For slope and distance from rivers, the map-only coefficient falls outside the confidence intervals for all methods that use ground truth points.
-4.2 Tree cover linear regression
-Problem Formulation
-Lastly, we estimate linear regression coefficients for the association between 2021 tree cover percentage in the contiguous United States and several covariates: aridity index, elevation, and slope.
-Given $n$ data points, the linear regression model is
-$\mu=\theta_{\text{aridity}}\textit{x}_{\text{aridity}}+\theta_{\text{elevation}}\textit{x}_{\text{elevation}}+\theta_{\text{slope}}\textit{x}_{\text{slope}}$
-where $\mu$ is a $n\times 1$ vector of predictions, $\textit{x}_{j}$ is a $n\times 1$ vector of covariate $j$, and $\theta_{j}$ is the scalar linear regression coefficient for covariate $j$.
-We wish to estimate the $\theta_{j}$.
-A one-unit increase in covariate $j$ corresponds to a $\theta_{j}$ change in the outcome variable.
-We use $N=983238$ map product points sampled uniformly at random from the 2021 USFS Tree Canopy Cover (TCC) product [forest2023usfs] at 30 m resolution.
-The TCC product is derived from Landsat and Sentinel-2 top-of-atmosphere reflectance data, USDA NASS CDL data, and terrain data.
-Its tree cover percentage predictions are obtained from random forest models trained on reference data from human-labeled high-resolution imagery.
-For ground truth, we manually label $n=983$ high-resolution satellite images from Google Maps for tree cover percentage, at locations sampled uniformly at random from the study area.
-Because the source of ground truth is from a uniform random sample, we do not use any stratified methods.
-For the covariates, we use a map of the average Global Aridity Index from 1970-2000 (at 30 arc-seconds resolution) [trabucco2019global] and the NASA Digital Elevation Model [jpl2020] (at 30 m resolution).
-The aridity index is defined as a ratio between precipitation and evapo-transpiration, and is lower for more arid areas; its value ranges from 0.02 to 6.55 in our dataset.
-Increased aridity is associated with lower tree cover (and decreased vegetation in general) [bastin2017extent, williams2010forest, berdugo2020global].
-Elevation and slope have also been associated with tree cover globally; there is generally more tree cover at lower elevations (e.g., due to higher temperatures) [mayor2017elevation] and on sloped terrain [sandel2013human].
-We do not include a comprehensive list of covariates that could influence tree cover, because the primary purpose of this regression is to illustrate uncertainty quantification rather than explain tree cover.
-Figure 12: US tree cover linear regression datasets.
-We regress the tree cover percentage target variable (left) against three covariates (right): aridity index, elevation, and slope.
-We use the USFS Tree Canopy Cover map product and $n=983$ uniformly randomly sampled ground truth points.
-Results
-Figure 13: US tree cover linear regression coefficient 95% confidence intervals using $N=983238$ map product points and $n=983$ ground truth points.
-PPI outperforms the classical method.
-For elevation, the map-only coefficient falls outside the PPI confidence interval.
-The 95% confidence intervals for the four linear regression coefficients are shown in Figure 13.
-The PPI confidence intervals for coefficients are as follows.
-Tree cover percentage increases by 35 to 40 percentage points for each one-unit increase in aridity index, decreases by 2.4 to 4.8 percentage points for each one-kilometer increase in elevation, and increases by 0.9 to 1.3 percentage points for each one-degree increase in slope.
-The PPI effective sample size outperforms the classical method by a factor of 3.3 for aridity index, 2.4 for elevation, and 3.6 for slope.
-As in the other use cases, PPI uses the map product along with the ground truth to reduce uncertainty in the coefficient estimates compared to only using ground truth.
-For elevation, the map-only coefficient computed from the map product ($-5.3$) falls outside the PPI confidence interval.
-This implies that there is bias in the tree cover map product’s association with elevation (e.g., the map product overestimates tree cover at low elevation locations or underestimates tree cover at high elevation locations).
-Similar to the deforestation logistic regression use case, using only the map product could result in misleading conclusions.
-If we had used only the remotely sensed tree cover predictions, we would have overestimated the magnitude of the relationship between elevation and tree cover due to map errors.
-5 Discussion
-Using remote sensing-based maps to draw scientific inferences requires uncertainty quantification and propagation of uncertainty to downstream analyses — otherwise, we risk making, in the words of [mcroberts2011satellite], just pretty pictures.
-In this work, we applied several methods that use remote sensing map products, along with a small amount of ground truth data, to estimate area and regression coefficients while correctly taking into account the errors in the map products.
-We find that stratified, post-stratified, and prediction-powered inference estimates result in lower uncertainty than classical estimates using only ground truth data sampled.
-Unlike map-only estimates using the map product alone, these estimates are guaranteed to be unbiased.
+E.4 Discussion
 Bias is the main issue for map-only estimates.
-Our experiments illustrated that a map-only estimate of area or regression coefficients from a map without bias correction can yield incorrect inferences, and we differentiated between the influence of noise versus bias.
+Our experiments illustrated that a map-only area estimate from a map without bias correction can yield incorrect inferences, and we differentiated between the influence of noise versus bias.
 In our simulation estimating maize fraction, map-only estimates deviated from the true maize fraction of 0.35 as the bias of the crop type map increased.
-Furthermore, accuracy is not necessarily a good proxy for bias; even at seemingly high map accuracies ($\text{accuracy}>0.85$), the map-only approach estimated maize fraction at 0.21 (a 40% underestimate).
-In our experiments on real data, the map-only estimate for deforestation in the Brazilian Amazon fell within stratified and PPI confidence intervals, suggesting an overall low-bias map product.
-However, the map-only estimates of the logistic regression coefficients on distance from rivers and slope fell outside the intervals computed through the classical, stratified, PPI, and stratified PPI methods.
-The map-only estimator would have concluded that distance from rivers is significantly more predictive of deforestation than it actually is, and that slope has no relationship to deforestation when it actually does.
-Thus, even when a map product is unbiased for the average outcome, it can be biased for the purpose of estimating regression coefficients.
-Similarly, the map-only estimate of the elevation coefficient in our linear regression for tree cover also overestimates the effect of elevation compared to PPI.
-Effect of map accuracy.
-How much stratified, post-stratified, and PPI estimators reduce uncertainty relative to the classical estimator depends on the quality of the map product.
+Furthermore, map accuracy is not necessarily a good proxy for bias; even at seemingly high map accuracies ($\text{accuracy}>0.85$), the map-only approach estimated maize fraction at 0.21 (a 40% underestimate).
+Stratification vs. Post-stratification vs. PPI.
+How much stratified, post-stratified, and PPI estimators reduce uncertainty relative to the GT-only estimator depends on the quality of the map product.
 In stratification, the strata map should be designed such that the variance of the quantity of interest is low within each stratum.
@@
 In our maize fraction experiments, a highly accurate map ($\text{accuracy}=0.96$) was equivalent to having nearly 6×6$\times$6 × as much ground truth data, whereas a low-accuracy map ($\text{accuracy}=0.73$) only increased the effective sample size by 1.2×1.2$\times$1.2 ×.
-Stratification vs. Post-stratification vs. PPI.
-Comparing stratified, post-stratified, and PPI estimators, which estimator yields the smallest confidence intervals depends on how the strata were designed and the quality of the map product.
-Recall Table 2, which summarizes the conditions under which each method is most appropriate.
 Empirically, when estimating deforestation fraction in the Brazilian Amazon, we observed the smallest confidence intervals when stratification was used alone or in conjunction with PPI.
@@
 PPI and the post-stratified estimator yielded comparable interval widths to each other.
-This is expected, as the post-stratified estimator yields similar confidence intervals as PPI when $N$ is large (Appendix A).
-The picture changes, however, when we reuse the same stratified samples to estimate regression coefficients in regressions that the strata were not designed for.
-When we associated deforestation with covariates in a logistic regression, PPI and stratified PPI resulted in narrower confidence intervals than the stratified estimator across all 4 coefficients (distance from road, distance from river, elevation, and slope).
-The worse performance of the stratified estimator is because the strata were designed to minimize the variance of deforestation within each strata, which does not necessarily correspond to better regression estimation.
-Indeed, the stratified estimator was actually worse than the classical estimator for certain coefficients.
-An advantage of PPI compared to the stratified estimator, then, is its flexibility.
+This is expected, as the post-stratified estimator yields similar confidence intervals as PPI when $N$ is large (Appendix F).
+An advantage of PPI compared to the stratified estimator is its flexibility.
 While the stratified estimator requires us to choose a map product before sampling, PPI can be applied to any map product and any randomly sampled ground truth dataset.
@@
 Or, if stratified samples are already available from area estimation, they can be combined with PPI to estimate regression coefficients.
-Recommendations for map producers and users.
-Ultimately, the best methods combine map products with ground truth to yield the narrowest intervals (highest certainty).
-We recommend that producers of machine learning-generated remote sensing maps use a (simple or stratified) randomly sampled holdout ground truth dataset for uncertainty quantification.
-The criterion for this holdout set is that the model, including during hyperparameter tuning, was not trained on it; a standard test set would be adequate, provided that it is randomly sampled.
-Map producers should release this ground truth data alongside their map product to make their map suitable for downstream scientific inferences.
-In the absence of a provided holdout set (currently, releasing ground truth data is not the norm), we recommend that map users generate their own ground truth data for the variable of interest if possible.
-Limitations.
-All of the estimation methods described in this paper (except for map-only estimators) are limited to areas in which ground truth data is available.
-However, this would be true for any assumption-free, data-driven method that quantifies uncertainty and corrects for map bias.
-Furthermore, we can only ensure that our estimates are unbiased when the ground truth points are randomly sampled.
-Together, these attributes mean that such methods may be especially straightforward to use when ground truth data can be generated de novo, such as from human inspection of available satellite imagery, and more challenging to use when ground truth data requires bespoke surveys.
-Another limitation is that our estimates only quantify uncertainty in the target variable map product, and do not account for noise in either the ground truth data or (in the regression use cases) the covariates.
-We leave these investigations to future work.
-6 Acknowledgments
-This material is based upon work supported by the U.S.
-Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number(s) DE-SC0023112.
-This report was prepared as an account of work sponsored by an agency of the United States Government.
-Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights.
-Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof.
-The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
-[…]
-Appendix A Mathematical comparison of PPI and post-stratified area estimator
+Appendix F Mathematical comparison of PPI and post-stratified area estimator
 We show that when the number of map product points $N$ is large, the PPI and post-stratified area estimator confidence intervals are similar.
-In particular, we show (1) that they have approximately the same length when the number of samples is large and (2) they are centered on the same point.
+In particular, we show (1) that they have approximately the same length when the number of sample units is large and (2) they are centered on the same point.
 These two facts together provide a mathematical explanation for why PPI and the post-stratified estimator approximately agree in all of our experiments.
@@
 Standard errors.
-We now turn to the PPI standard error. We have
-$SE_{\text{PPI}}$ $\rightarrow\sqrt{\frac{1}{n}\mathrm{Var}(Y-\lambda\hat{Y})}$ since $N\gg n$
+We now turn to the PPI standard error. We have the standard deviation of the tuned PPI estimator from [angelopoulos2023ppi++]
+$\sigma_{\text{PPI}}$ $\rightarrow\sqrt{\frac{1}{n}\mathrm{Var}(Y-\lambda\hat{Y})}$ since $N\gg n$
 $=\sqrt{\frac{1}{n}\left(\mathrm{Var}(Y)+\lambda^{2}\mathrm{Var}(\hat{Y})-2\lambda\mathrm{Cov}(Y,\hat{Y})\right)}$ expanding out $\mathrm{Var}(Y-\lambda\hat{Y})$
@@
 $\lambda\rightarrow\frac{\mathrm{Cov}_{n}(Y,\hat{Y})}{\mathrm{Var}_{N}(\hat{Y})}\approx\frac{n_{00}n_{11}-n_{01}n_{10}}{n_{0\cdot}n_{1\cdot}}.$
+A standard error estimate $\hat{\sigma}_{\text{PPI}}$ of $\sigma_{\text{PPI}}$ also satisfies the above equations.
 We compare this with the post-stratified estimator standard error
-$SE_{\text{post}}$ $=\sqrt{A_{0}^{2}\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+A_{1}^{2}\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$
+$\hat{\sigma}_{\text{post}}$ $=\sqrt{A_{0}^{2}\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+A_{1}^{2}\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$
 $\approx\sqrt{\left(\frac{n_{0\cdot}}{n}\right)^{2}\times\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+\left(\frac{n_{1\cdot}}{n}\right)^{2}\times\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$ plugging in $A_{0}$ and $A_{1}$
@@
 We see that the final expressions are equal, which means the point estimates are approximately equal.
-Appendix B Binarizing the NASA Global Forest Cover Change map product for deforestation
-In the Brazilian Amazon deforestation use cases, the ground truth points from [bullock2020satellite] have binary deforestation values ($Y=1$ if deforested and $Y=0$ if not deforested between 2000-2015), while the NASA GFCC map product gives canopy cover percentages over time at each pixel.
-In order to binarize the continuous map product data, we classify a map pixel as “deforested” if its canopy cover percentage satisfies $\text{canopy}_{2015}-\text{canopy}_{2000}\leq-25\%$.
-We choose the $-25\%$ threshold by visually comparing the map-derived canopy cover change distributions for deforested and non-deforested ground truth points (Figure 14).
-Most of the deforested ground truth points have a map canopy cover decrease of more than $25\%$, while most of the non-deforested ground truth points have either a map canopy cover increase, no map canopy cover change, or a map canopy cover decrease of less than $25\%$.
-Figure 14: NASA GFCC map canopy cover change (2000-2015) distributions for non-deforested and deforested ground truth over $n=1386$ uniformly random sampled points.
-The distributions are visualized using kernel density estimation.
-We choose a -25% deforestation threshold (shown in red) to binarize the map product.
+Appendix G Simulating mean reversion bias in maps
+In Section 4, we simulate adding noise and nonlinear bias to the map products from Example 4 (forest cover linear regression).
+In this appendix, we simulate another model of map bias: linear mean reversion.
+Under this type of bias, the map underestimates high ground truth values and overestimates low ground truth values.
+Using the same MOSAIKS dataset as Section 4, we simulate different levels of mean reversion bias in the population map (error-in-$X$), the forest cover map (error-in-$Y$), or both (error-in-both).
+The elevation dataset is not altered.
+We compare the resulting 95% coefficient confidence intervals at different levels of simulated map bias using map-only, GT-only, and PPI methods.
+Bias
+To simulate bias for the ground truth population covariate $X_{\text{pop}}$, we linearly interpolate between $X_{\text{pop}}$ and the mean $\mu_{X_{\text{pop}}}$ of all $X_{\text{pop}}$ values in the full ground truth dataset (67968 points).
+The resulting simulated map product value at “bias level $c$” is
+$X_{\text{pop}}^{\text{map}}=c\mu_{X_{\text{pop}}}+(1-c)X_{\text{pop}}.$
+Similarly, for the forest cover response variable $Y$ with mean $\mu_{Y}$, the simulated map product value at each point is
+$Y^{\text{map}}=c\mu_{Y}+(1-c)Y.$
+We experiment with bias levels $c$ from 0 to 0.8 at increments of 0.1. When $c=0$, the simulated map is the same as the full ground truth map.
+Results
+The coefficient point estimates and 95% confidence intervals at each bias level are shown in Figure 14.
+We compare results for the map-only, GT-only, and PPI estimators.
+(For all experiments in this section, we use $B=200$ bootstrap iterations to construct PPI confidence intervals.)
+Similar to Section 4, the GT-only confidence interval stays constant regardless of the bias level, and it is wider than the PPI and map-only intervals due to the small number of ground truth points ($n=500$).
+At all bias levels across all three experiments (error in $X_{\text{pop}}$, $Y$, or both), the PPI confidence intervals contain the true coefficients for both covariates.
+This is expected, as PPI uses the small set of ground truth points to correct for bias in the map product, and the PPI confidence intervals are guaranteed to contain the true coefficient with 95% probability.
+At all bias levels, the PPI confidence intervals are significantly smaller than the GT-only intervals.
+This is because the biased map is a simple linear transformation of the true variable values and thus provides useful information for estimating the regression coefficients.
+As discussed in Section 4, PPI allows us to conclude that there is a positive association between forest cover and population, while the GT-only interval for $\beta_{\text{pop}}$ contains zero and does not detect this significant association.
+In all three experiments, as map bias increases, the map-only estimates become biased for at least one of the coefficients.
+The map-only confidence intervals are very narrow because the number of map product points is large ($N=67968$), so they fail to contain the true coefficients as the estimates become biased.
+As mean reversion bias in $X_{\text{pop}}$ increases, the map underestimates high population values and overestimates low population values.
+As a result, the map-only estimates for $\beta_{\text{pop}}$ increase in magnitude and overestimate the true coefficient, while map-only estimates for $\beta_{\text{elevation}}$ remain unbiased (since we do not add bias to the elevation map).
+As mean reversion bias in $Y$ (forest cover) increases, the map underestimates high forest cover values and overestimates low forest cover values.
+As a result, the map-only coefficient estimates for both covariates decrease in magnitude and attenuate toward zero.
+When there is mean reversion at level $c$ in both forest cover and population, the effects of the two biases cancel out to produce unbiased map-only estimates of $\beta_{\text{pop}}$.
+However, the map-only estimates of $\beta_{\text{elevation}}$ still decrease in magnitude and attenuate toward zero.
+Figure 14: Simulation estimating forest cover linear regression coefficient 95% confidence intervals under different levels of map mean reversion bias.
+PPI gives unbiased estimates with much narrower confidence intervals than the GT-only method.
+In each experiment, the map-only coefficient estimate for at least one covariate becomes increasingly biased as map bias increases.